### v2

# Labelling Data with Unknown References

Abstract
An evaluator is trustworthy when there exists some agreed-upon way to measure its performance as a labeller. The two ways to establish trustworthiness are either by testing it, or by assuming the evaluator ‘knows’ somehow the way to label the corpus. However, if labelled references (e.g., a development set) are unavailable, neither of these approaches work: the former requires the data, and the latter is an assumption, not evidence. To address this, we introduce an algorithm (the ‘No-Data Algorithm’) by which to establish trust in an evaluator without any existing references. Our algorithm works by successively posing challenges to said evaluator. We show that this is sufficient to establish trustworthiness w.h.p., in such a way that when the evaluator actually knows the way to label the corpus, the No-Data Algorithm accepts its output; and, conversely, flags untrustworthy evaluators when these are unable to prove it. We present formal proofs of correctness and limited experiments.
1 Introduction
One of the most fundamental problems in machine learning, and AI broadly, is measurement. It has been known for a while that some metrics have a certain level of unreliability; for example, by not corresponding to human judgements appropriately [34, 11, 3, 13]. However, the rise of LLMs as evaluators, typically known as LLMs-as-judges, has led to a substantial divide in the field regarding their trustworthiness as a measurement tool. Their use has also led to broader questions around measurement in AI; specifically, about the validity of some results [11, 33, 25].
A key component in the ‘measurement in AI question’, and the object of study of our work, is whether the tool providing the evaluation (the evaluator) may be trusted. This has become more critical given the fast adoption of foundation models in industry and scientific research, and the continuous need to scale. In these cases, trust is usually established based on some sort of ground truth: an existing set of labelled references by which to perform statistical analyses (e.g., a $t$-test). However, it is often the case that said references do not necessarily exist: perhaps the problem is too niche, or the annotation too costly. It follows then that the only way one can establish trust on an unknown evaluator in a no-reference scenario is via formal methods (i.e., formal proofs).
In this paper we introduce an algorithm (which we call the ‘No-Data Algorithm’) by which to establish trust in an evaluator without previously-labelled data. The algorithm is designed to work in-line with contemporary machine learning techniques, where an evaluator claims that it knows the map between datapoints and labels. It is up to a verifier to decide, based on the outcome of certain challenges, whether the evaluator may be trusted. The No-Data Algorithm requires a linear number of calls and does not require labelled datasets (e.g., development or training sets) or knowledge of the labels in the test set to establish trust.
We show that our algorithm, after $r$ calls, bounds confidence on the labeller to $(1/4)^{r}$. This success rate reflects the ‘true’ accuracy of the evaluator: if the evaluator knows how to label the set, the accuracy and success rates will be high; if the evaluator lies about knowing the labels, the accuracy could be high, but the success rate won’t. We present small experiments in synthetic data for traditional machine-learning (i.e., a decision tree) and contemporary (r. an LLM) methods which adjust to the theory’s predictions, although we note our contribution is theoretical in nature. [^1] Code and data may be found in https://github.com/adewynter/no_data_algorithm. Amongst our experimental findings we note that the behaviour of a reliable (read: trained) evaluator adjusts to our theoretical predictions, while an LLM requires more work due to prompting and spurious reasoning. Nonetheless, when properly tuned, it adjusts to the predictions. Our results demonstrate that trust in an evaluator may be established via formal methods without the need for references. A side effect of our experimentation opens a wide range of questions around faithful generation over arbitrary strings for a generative model.
2 Related Work
There has been considerable work on attempting to create reliable evaluators. Most of contemporary work relies on LLMs and various prompting strategies or call stacks, although the problem of (un)reliability/trustworthiness of an evaluator is older than generative models. For example, it has been known that the use of BLEU for tasks other than English-based machine translation is not accurate [22, 28]; and even then, that it does not correspond well to human judgements [34, 11].
More generally, it is known that metric choice, and hence the choice of evaluator, is system-dependent [4, 25, 10, 39, 32]. Poorly-developed rubrics also lead to unreliable human judgements [7, 20], and thus the evaluator trustworthiness problem is also a rubric problem. There are algorithms [27, 47] and paradigms [26] for learning with noisy teachers. Likewise, provable methods have been used with success in other areas, such as representation learning [15], and to improve robustness via synthetic data generation [8, 9]. However, they all assume that a (perhaps noisily) labelled dataset exists.
Relevant to our work, the use of LLMs as evaluators has promised extraordinary scaling capabilities, such as fast data generation and evaluation without humans [29, 6, 5, 23, 41, 48]. Some of these criticisms are that they do not correlate with human judgements well [3, 13, 4, 45, 37]; are sensitive to their prompt [24, 14, 46]; their performance could be the product of memorisation [19, 44, 35]; and even that they do not understand the task at all [43, 40, 46]. Their reasoning capabilities have been put into question, from the results being strongly dependent on the choice of metrics [36, 4], to findings that their output reasoning steps contain spurious reasoning [38, 18, 42]. Most of the work on improving trust, however, relies on prompting strategies or call stacks to other LLMs [21].
The main subroutine from the No-Data Algorithm, which we call the Evaluator-Verifier (EV) protocol is based off the well-known Arthur-Merlin (AM) protocol from Babai [2]. It is a type of zero-knowledge proof; see [12] and [1] for primers on the subject. That said, the EV protocol is different in many regards to the AM protocol, since it is designed to fit arbitrary inputs and the larger No-Data Algorithm. In particular, the EV protocol always reveals its secrets, which sets it apart from applications in theoretical computer science. Likewise, our use of the rubric and the aggregator functions are generalisations to partial functions that are not present (or needed) in previous works.
3 Background
4 Notation
Our work relates to datapoints $x\in X$ and labels $y\in Y$. Whenever the label set $Y$ is unspecified, it is the binary label set $Y=\{0,1\}$. We also let $X\subset\{0,1\}^{n}$; that is, every datapoint is a $n$-bit binary string modelling a phenomenon. The assumptions on $X$ and $Y$ are to simplify our proofs; see Section 9 for a discussion on their relevancy.
Likewise, for a binary string $x\in\{0,1\}^{n}$, we denote its relevant subset set as $S_{x}\subset 2^{x}$. The construction of this specific set will depend on the setup, and will be discussed in detail in Section 6.1.
For two functions $f\colon B\rightarrow C,g\colon A\rightarrow B$ we write the composition $f\circ g\colon A\rightarrow C$ as $fg$.
For two sets $A$, $B$, we denote equality up to isomorphism as $A\cong B$, where $A\cong B$ if $f(A)=f(B)$ but $A\neq B$ for some $f\colon U\rightarrow V$. We also consider the special case of equality up to permutation over subsets as $A\equiv B$, where $A\equiv B$ if $\forall a\in S_{A},\exists b\in S_{B}\,.\,g(a)=g(b)$, for some fixed function $g\colon U\rightarrow V$. We assume all functions to be deterministic throughout, unless stated otherwise.
5 Definitions and Problem Setup
In line with standard learning theory, we assume that there exists an unknown map $f\colon X\rightarrow Y$ that takes datapoints to labels faithfully–that is, $f$ provides the true labelling. Our task is to determine whether a function $E\colon X\rightarrow Y$ (a labeller, or evaluator) is equivalent to $f$, without any knowledge of the map itself.
For that, we introduce four definitions:
A criterion $c\colon X\rightarrow\{0,1\}$ is a map that takes in a datapoint $X$ and maps it to a single bit.
The rubric $C\colon\times_{i,\dots,n}\{0,1\}\rightarrow\{0,1\}^{n}$ takes in the binary outputs from an ordered set of criteria $\mathcal{C}=\{c_{1},\dots,c_{n}\}$ and returns a bitstring that is the concatenation of these outputs: $C=c_{1}|c_{2}|\dots|c_{n}$. We say a rubric is total if it explicitly decomposes nonlinear criteria with arity $>1$ (e.g. the xor $c_{i}=c_{a}\oplus c_{b}$), and denote it as $\bar{\mathcal{C}}$. When the evaluation of said criteria can be done separately (e.g., by testing first $c_{a}$, then $c_{b}$, and then their xor $c_{a}\oplus c_{b}$), we say it is a total evaluation.
The aggregator $\sigma\colon\{0,1\}^{n}\rightarrow Y$ is a function that takes in a binary string of length $n$ (the output from the rubric) and maps it to the label set.
Finally, for two $x,x^{\prime}\in X$, we say that $x$ is similar to $x^{\prime}$ if both $x\cong x^{\prime}$ and $x\equiv x^{\prime}$.
Intuitively, the rubric is the set of reasons why a datapoint $x$ would take a given label (for example, an annotation rubric passed in to human labellers, the last layer in a model, etc.); while the aggregator decides the final label based on that rubric (e.g., via an activation function, a majority vote, etc.).
We are now ready to introduce the core assumption of our work:
Assumption 1.
The map $f\colon X\rightarrow Y$ is a composition of the rubric and the aggregator. That is,
$f=\sigma C.$ (1)
While not a particular strong assumption (after all, datapoints must have a reason to acquire a label and, when there is more than one reason, there must be a way to decide how to aggregate them) this decomposition is crucial to our work.
6 The No-Data Algorithm
The No-Data Algorithm is an algorithm with two components (players): the evaluator and the verifier. [^2] In theoretical computer science, the evaluator is typically known as the prover. We use this more-intuitive naming to highlight the distinction between our work (data-labelling) and others (e.g. theorem-proving). Throughout the algorithm, for every datapoint $x\in X$, the evaluator must convince the verifier that it is able to label $x$ correctly (i.e., that it knows $f$).
For this, the players run a multi-round sub-game which we refer to as the Evaluator-Verifier (EV) protocol. The EV protocol takes in $x$, and either succeeds or ends in failure, w.h.p., based on challenges posed by the verifier on a $x^{\prime}$ similar to $x$ generated by the evaluator. It then returns the proposed label and the status (success or failure).
The algorithm then, in case of failure, decides to flip the label with probability $\phi$. At the end of the loop, it aggregates the results of the EV protocol and returns the final labelling and success count. The final labelling depends on the choice of $\phi$, and the evaluator’s trustworthiness (and hence that of said labelling) is be determined by measuring the success count.
The full algorithm may be found in Algorithm 1. Given that the EV protocol is a fundamental part of the No-Data Algorithm, we focus the remainder of this section to describe it, and provide a characterisation of the No-Data Algorithm (bounds, correctness, etc.) in Section 7.1.
1:  Input: Unlabelled data $X$, evaluator $E$, verifier $V$, criterion $C$, flip probability $\phi$, rounds $r$
2:  $predictions\leftarrow\{\}$
3:  $successes\leftarrow\{\}$
4:  for $x\in X$ do
5:     $success,y\leftarrow\textsc{EV}(x,E,V,C,r)$// Run for $r$ rounds
6:     if $success$ then
7:        $predictions\leftarrow predictions\cup\{y\}$
8:        $successes\leftarrow successes\cup\{1\}$
9:     else
10:        $\tilde{y}\leftarrow\lnot y\text{ with probability }\phi$ // The opposite label
11:        $predictions\leftarrow predictions\cup\{\tilde{y}\}$
12:        $successes\leftarrow successes\cup\{0\}$
13:     end if
14:  end for
15:  return  $predictions$, $successes$
Algorithm 1 The No-Data Algorithm. It works by calling a subroutine (the EV Protocol) to provably determine the trustworthiness of the labels produced by the evaluator. Based on its output (success or failure), the algorithm then determines which label to assign to the corpus, with a stochastic label flip $\phi$ added to it in case of failures.
6.1 The Evaluator-Verifier (EV) Protocol
In the EV protocol both players are given a datapoint $x\in X$, and the rubric $C$, but nothing else. In particular, $\sigma$ is not explicitly given to either player. However, the evaluator claims to know the map $f\colon X\rightarrow Y$. Since $C$ is given, then the claim is that it has prior knowledge of $\sigma$. The implication–were the evaluator to convince the verifier–then is that its choice of label $y$ for $x$ may be trusted.
Remark that the verifier does not know if $y$ is correct. In order to be convinced, the verifier chooses one out of two challenges, neither of which rely on $\sigma$ or $y$. Based on whether the evaluator passes or fails them, the verifier builds its own belief on the evaluator’s trustworthiness.
The entire point of the EV protocol (and thus of the No-Data Algorithm) is for the evaluator to convince the verifier that its choice of label for $x$ is correct. If the evaluator fails, it follows that the correct label for $x$ could be the opposite label. Hence, regardless of success, at the end of the EV protocol the evaluator reveals its label. See Figure 1 for a diagram of the EV protocol. We now describe each of the components and flow.
6.1.1 Setup
At every round of the EV protocol, the evaluator takes in $x$ and generates a similar datapoint $x^{\prime}$ and a ‘partial’ label $\tilde{y}^{\prime}$ that–ideally–is equal to $y$. It is partial because there is no access to $f$, and hence the best the evaluator can do is to generate a guess $\tilde{y}^{\prime}$ based on the encoding $C(x^{\prime})$. If the evaluator indeed knows $f$, an application of $\sigma$ to $C(x^{\prime})$ would lead to the ‘true’ label $y$.
That said, once it has generated the tuple $\langle x^{\prime},\tilde{y}^{\prime}\rangle$, the evaluator may not change them in this round.
Remark that isomorphism is not a sufficient condition for $\tilde{y}^{\prime}=y$ since $C$ might not be total. Hence, both structural and valuation checks are needed. Therefore, after the tuple has been generated, the verifier then chooses uniformly at random one of the two following challenges:
Figure 1: Diagram of the EV protocol. At every round, the evaluator (blue) generates a datapoint $x^{\prime}$ similar to $x$, and a partial label $\tilde{y}^{\prime}$. It then responds to one of two (chosen uniformly at random) challenges by the verifier (orange). The game is repeated by $r$ rounds and returns a succeeded state, unless there is a failure in the challenges, and thus returns a failed state. In either case, it also returns the proposed partial label, $\tilde{y}^{\prime}$.
Challenge 1 (Equality up to Permutation)
Setup: The verifier tests the evaluator by asking it to show a similar point that leads to the same total evaluation of $C$. This challenge is surrogated by testing whether $x\equiv x^{\prime}$ over $\bar{\mathcal{C}}$.
Check: The verifier asserts that
$\forall s\in S_{x}\exists t\in S_{x^{\prime}}\,.\,\forall c\in\bar{\mathcal{C}}\,c(s)=c(t);$ (2)
namely, whether there exists at least one substring in $x^{\prime}$ matching the (total) evaluation of a criterion from $x$.
Challenge 2 (Equality up to Isomorphism)
Setup: The verifier asks for the rationale as to why $f(x^{\prime})=y$. Since $y$ is not available, it checks whether $x^{\prime}\cong x$.
Check: The verifier asserts that $C(x^{\prime})=C(x)$.
6.1.2 Verification
If the evaluator failed the tests, the EV Protocol returns a failure state. Otherwise, it restarts from Section 6.1.1 over and over until the maximum number of rounds $r$ has been reached, or the evaluator fails a test. In either case, at the end of the call the EV Protocol also returns $\tilde{y}^{\prime}$.
Remark 1.
Since the challenges are public, the evaluator could cheat and generate an $x^{\prime}$ to pass one of the challenges, but not both. Otherwise, the implication is that it knows $f$, and it will pass the challenges every time. This is why the challenges assert both structure ($x\equiv x^{\prime}$) and final encoding ($x\cong x^{\prime}$). This observation will be elaborated further in our correctness proofs ( Section 7.1).
Remark 2.
The check for Challenge 1 (demanding equality up to criterion) is very strict. We discuss this in Section 9 and present a relaxed extension in Appendix C.
7 Algorithm Characterisation
In this section we provide proofs of correctness (Section 7.1). Runtime bounds may be found in Appendix B.
7.1 Correctness
Lemma 1 (EV Protocol Correctness Bound).
The probability that the verifier fails to detect a lie by the evaluator in the the EV protocol sub-game after $r$ rounds is $(1/4)^{r}$.
Proof.
(Sketch; a complete proof is in Appendix A.1). We separate the proof in two parts: whether the challenges are necessary and sufficient to determine whether the evaluator is lying; and the probability bound to which it happens.
On the first part, we note that Challenge 1 determines the internal structure of the string, but does not account for the valuation of $C(x^{\prime})$. This challenge is necessary because two $x,x^{\prime}$ could lead to the same evaluation of some $c\in\mathcal{C}$ (e.g., disjunctive clauses), and that would not mean that they are the same up to permutation.
However, Challenge 2 requires strict equality in $C(x^{\prime})=C(x)$, and it is indeed order-dependent, but does not account for said internal structure. Hence both challenges provide some, but not all, information about $x,x^{\prime}$ and the claimed $f(x),f(x^{\prime})$. They are then both sufficient and necessary to determine if an evaluator is lying. More importantly, they are mutually exclusive in the information they provide. This is needed because a lying evaluator could know ahead of time the challenges.
For the second part we handle two cases where the evaluator could lie (or cheat): either (1) it can generate an appropriate $x^{\prime}$, but not a correct $C(x^{\prime})$; or (2) it provides a correct $C(x^{\prime})$ but an incorrect $x^{\prime}$. In either case, the probability that it gets lucky and/or passes the challenge by cheating is $1/4$; after $r$ rounds, it is $(1/4)^{r}$. ∎
Remark 3.
The exponential bound from Lemma 1 means that, for example, after running the EV protocol for 3 rounds, the probability that the verifier fails to detect a lie from the evaluator is 1.6%.
Theorem 1 (No-Data Algorithm Correctness Bound).
Suppose an evaluator has accuracy $\alpha$ on a balanced dataset $X$ of size $n$ with binary labels $Y=\{0,1\}$. Assume the evaluator always lies on any $x\in X$ iff it incorrectly labels the datapoint. Then if the EV protocol is ran for $r$ rounds, the expected accuracy of the No-Data Algorithm is given by
$\mathrm{E}[\text{correct}]\leq 1-(1-\alpha)\left(1-\phi+\phi\left(\frac{1}{4}\right)^{r}\right),$ (3)
for a chosen value of $\phi$ and over the same dataset $X$.
Proof.
(Sketch; a complete proof is in Appendix A.4). The accuracy of any evaluator is directly tied to its success in the algorithm: if it knows the label, then the checks by the EV protocol will pass. Even under the constraint that the mislabelled datapoints are tied to the evaluator’s lies, errors could still get past undetected. From Lemma 1, we know the error bound for every label flip, and therefore the probability of mislabelling the entire dataset. We may compute the latter via an application of the union bound, that is, $\Pr[\text{error}]\leq(1-\alpha)(1-\phi+\phi(1/4)^{r})$. ∎
The implication of Theorem 1 is that a good evaluator with low probability to flip the labels will obtain a good label accuracy. A bad evaluator with high probability to flip the labels will also obtain a good accuracy. Another interpretation of this, however, is that there is no way to recover unknowable labels when the evaluator is unable to convince the verifier–we discuss this further in Section 9.
It is worth mentioning that the assumption for Theorem 1 around always lying when mislabelling (and hence always being wrong) are not necessarily applicable to all evaluators, since a particularly tricky evaluator could still guess the right label and yet not pass the challenges. We consider this out of scope, since the ability to guess a label without explaining how is not a viable way to establish trust in scenarios without any references. Consequentially, for the No-Data Algorithm it is more helpful to look at the success count.
8 Experiments
We describe the two experiments and results in this section and refer the reader to Appendix D for a more detailed treatment, including prompts and specifics on the models used. Throughout this section, we run the No-Data Algorithm with $r=3$.
8.0.1 Data
We created two mostly disjoint synthetic datasets so that we could control the labels, data, and their rubrics. We refer to these datasets as in-phenomenon, or IP; and out-of-phenomenon, or OOP, as opposed to in-distribution (r. out-of-distribution) to highlight the fact that the entire measurement is altered by the change in data–and, that, semantically, we are modelling different phenomena. The IP dataset is our baseline: the evaluator will not be lying on knowing this phenomenon. The OOP dataset is our reference for a corpus that the evaluator has not seen before (and thus it will lie about knowing it). For both problems, $\sigma$ is the majority vote.
We chose to generate our own data to ablate out any potential memorisation concerns, and to control the rubrics. These may be found in Table 1. The test sets for both IP and OOP included 498 entries and are balanced.
8.0.2 Evaluator
We used a decision tree (DT) for our evaluation, and an LLM (o3-mini; [31]). Both models observed a balanced training split of the IP dataset: the DT was trained on it, while the LLM observed five exemplars in the prompt. The rationale to train the DT was to fully ensure that the evaluator learnt both $\sigma$ and $C$. For generation it only had access to the rubric, but not $\sigma$. The LLM observed it via the prompt, and was made aware of the aggregation function only for the labelling step. For generation it was asked to pick the best candidate from the dataset, as opposed to outputting a new point, as we observed marked performance decreases otherwise. During OOP, the LLM was given the rubric for IP and an OOP datapoint; while the DT was trained in IP (and, ideally, learnt the rubric), and classified IP and OOP datapoints.
Rubric Definition
In-Phenomenon
$c_{0}$ $y=1$ if $x$ has an even number of ones; else 00
$c_{1}$ $y=1$ if $x$ starts with a zero or contains $10101$, but not both; else 00
$c_{2}$ $y=1$ if $x$ has strictly more than five ones; else 00
Out-of-Phenomenon
$c_{0}$ $y=1$ if $x$ contains the substring $111$; else 00
$c_{1}$ $y=1$ if $x$ ends with a one; else 00
$c_{2}$ $y=1$ if $x$ contains the substring $110001$; else 00
Table 1: Rubrics and criteria for our experiments. For both problems, $\sigma$ was the majority vote. Remark that $c_{1}$ for the IP rubric is an exclusive-or operation. This means that, while this rubric contains three criteria, the total rubric contains five: $\{c_{0},c_{2},c_{1,a},c_{1,b},c_{1}\}$; where $c_{1,a},c_{1,b}$ are the clauses for $c_{1}$, and $c_{1}$ is the exclusive-or itself.
8.0.3 Verifier
The verifier is a rule-matching algorithm that has access to the relevant rubrics (e.g., when testing IP, it observes the rubric for IP), and works as described in Section 6.1.
8.1 Results
The results are in Table 2. Both the DT and the LLM adjusted to the predictions correctly, successfully convincing the verifier when it was an oracular generator, reaching approximately their knowable accuracy, and outputting low success rates when lying. Indeed, the accuracy of the models in the OOP dataset hovers around 54%–close to random guessing. A similar accuracy can be obtained by the output of the No-Data Algorithm, but the success rates between IP (100%) and OOP (5%) provide better information as to in which situations the evaluator is trustworthy. The tuned $\phi$ value, based on the known labels, also yielded accuracies close to the expected score from the known dataset. Of note, which we will cover in Section 9, is the number of flips performed by the algorithm in the DT OOP dataset: 46%.
As part of our experimentation, we performed ablation studies on the relationship of lying with evaluators and generators (Section 8.1.1); and the effectiveness of generator strategies in LLMs (Section 8.1.2).
8.1.1 Ablation: Lying Evaluators and Lying Generators
The results are in Table 3. Throughout that study DT randomly (w.p. $1/10$) flipped the label when evaluating. The lies tested were (in order in the table) when the model claims to
  1. 1.
    know $\sigma$: that is, it can output an encoding $C(x)$, but does not necessarily know $\bar{C}(x)$;
  2. 2.
    know $f$: that is, it can retrieve a datapoint $x^{\prime}\in X$ such that $y^{\prime}=y$, but nothing else; and
  3. 3.
    approximately know $f$ up to a probability $p=1/10$ (r. it could pass both challenges with probability $1-p$).
The last ‘lie’ is more akin to how a realistic generator would behave, sometimes understanding the challenges well, but failing them with some probability.
As before, the results indicate that lying evaluators tend to score low in the success rate, even when their accuracies are close to the known case. Remark that (a) the evaluator that lies on knowing $\sigma$ had a success rate much higher than that of lying of knowing $f$; and (b) the evaluator with approximately knowledge of $f$ had a comparatively low (44%) success rate. We discuss both findings in Section 9.
DT (known) DT (unknown) LLM (known) LLM (unknown)
IP
Successes/Flips — / — 100.0 / 0.0 — / — 81.3 / 1.8
Accuracy/F1 62.2 / 58.8 62.2 / 59.8 99.8 / 99.8 97.6 / 97.6
OOP
Successes/Flips — / — 4.8 / 46.4 — / — 28.0 / 6.0
Accuracy/F1 54.2 / 54.2 52.8 / 52.1 60.6 / 66.2 59.0 / 64.0
Table 2: Results for the known (test score if the labels were known) and unknown (output of the No-Data Algorithm) cases. The known case is a reference that in practice does not exist. For the unknown case, if we used the base generator, the performance is exactly the same as in the known case. The flip $\phi$ is set near the accuracy in the known case ($0.6$ for DT; $0.9$ for LLM). This is because the generator acts as an oracle, convincing the verifier every time, and hence the No-Data algorithm reduces to simple evaluation (the known case). However, the success rate is extremely low in DT OOP (unknown): even though there is no way for us to know whether the evaluator’s score is trustworthy, the success rate indicates deception.
IP (Flips) IP (No Flips) OOP (Flips) OOP (No Flips)
No Lie: DT
Successes / Flips 100.0 / 0.0 100.0 / —- 6.0 / 35.3 6.4 / —-
Accuracy/ F1 60.6 / 58.8 59.0 / 57.4 48.6 / 47.3 45.6 / 44.8
Lie 1: $\sigma$ unknown
Successes / Flips 17.0 / 35.1 15.1 / —- 2.8 / 35.1 2.2 / —-
Accuracy/ F1 58.6 / 58.8 42.6 / 45.0 52.4 / 53.8 49.0 / 49.4
Lie 2: $f$ unknown
Successes / Flips 0.6 / 37.4 0.2 / —- 3.0 / 35.7 2.8 / —-
Accuracy/ F1 50.6 / 48.1 39.4 / 42.4 51.4 / 52.2 51.2 / 52.8
Lie 3: generator w.p. $p=1/10$
Successes / Flips 43.6 / 24.9 51.3 / —- 4.4 / 37.9 4.6 / —-
Accuracy/ F1 55.8 / 54.2 49.0 / 49.4 47.6 / 49.5 46.6 / 48.6
Table 3: Performance for a lying DT evaluator on the No-Data Algorithm with $\phi=0.6$. From top to bottom: Lie 1: $\sigma$ unknown. Lie 2: $f$ unknown. Lie 3: $f$ approximately known (lying with probability $p=1/10$) in the generation step. In all these calls, the evaluator itself also outputs the incorrect label with probability $p=1/10$. Note the comparatively high success rates in the IP case of Lies 1 and 2. We discuss these findings in Section 9. Also note the consistently low success rate for OOP across all experiments, thus indicating the robustness of this algorithm to any evaluator strategy in scenarios where the data is completely unknown to the evaluator.
8.1.2 Ablation: LLM Generation Strategies
Given the LLMs’ ability to follow instructions, it is likely that their performance could be sensitive to how the datapoints are generated during the EV protocol. Hence in this study we compared the performance of both the LLM whose results we presented in the main results (o3-mini) with another LLM, GPT-4o [30]. We tested two different generation strategies: picking a matching datapoint from the data, as in Section 8.1, and actually generating a datapoint from the rubric. The results are in Table 4. We found that, again, the experimental results adjusted to the predictions, with lying and low-performing models being easily spotted by the No-Data Algorithm. However, we also found that actual generation led to bad performance in both LLMs. We note that this is expected: in general, classifying an element in a language is much easier than generating an element [17, 16].
o3-mini (IP) o3-mini (OOP) GPT-4o (IP) GPT-4o (OOP)
Picking
Accuracy/F1 97.6 / 97.6 59.0 / 70.0 58.0 / 67.1 56.4 / 65.4
Successes/Flips 81.3 / 1.8 27.9 / 6.0 28.7 / 71.3 10.8 / 8.8
Generating
Accuracy/F1 94.2 / 94.2 60.6 / 65.5 56.6 / 63.8 53.4 / 63.6
Successes/Flips 55.4 / 5.0 28.7 / 7.8 5.8 / 10.8 34.5 / 4.6
Table 4: Results of our ablation study on LLM generation strategies, comparing a high-performing model (o3-mini) with a lower-performing model (GPT-4o) as a LLM-based baseline. We report the highest-performing run at $\phi=0.9$. While both models were able to broadly adjust to their original accuracy, their success rates are markedly different when the models are generating a datapoint versus picking one.
9 Discussion
9.1 Theory
Theorem 1 states that when the labeller has high accuracy, a low $\phi$ will be able to reproduce the label set. It is then possible to measure the successes to establish a confidence bound on the aggregate labels provided by the evaluator. However, this result has a different interpretation: when the accuracy of the evaluator is zero, there is no way to determine the labelling, regardless of $\phi$–it is completely unknowable. We argue this interpretation also makes sense: otherwise, it would be possible to create information out of thin air (that is, without any prior knowledge of $f$). Moreover, the primary purpose of the No-Data Algorithm is to establish trustworthiness on the evaluator, not label the data.
One important limitation of the theory is that we assumed that $Y$ is a binary set, and that $X$ is the set of $k$-bit binary strings. The first assumption can be addressed by ‘binarising’ a set $Y$ with arity $>2$ by separating it into two classes. The second, however, must be understood more carefully: if this were a natural-language problem, it would be inaccurate to assume that the $k$-bit encoding for $x\in X$ will retain correctness by solely working with the binary encoding of the strings. Extra information (e.g., semantics captured in a latent space) must be added to $x$, before proceeding as usual.
9.2 Experiments
The results adjusted to the theory’s predictions. In particular, in both the known/unknown comparison and the lying ablation study, the LLM and DT evaluators consistently had low success rates for the OOP case, thus indicating that the No-Data Algorithm is robust to evaluator strategies attempting deception. Three things we noted throughout our experimentation were the high number of flips in the DT OOP case; the (relatively) high success rate when the evaluator lied about knowing $\sigma$; and the low success rate of the evaluator with approximate knowledge of $f$.
The first illustrates our interpretation of Theorem 1: since the model’s accuracy is near-random in the OOP dataset–as indicated by the known case–the verifier will almost always catch a lie. Since the flip probability is set to $\phi=0.5$, half the labels got flipped and no information was gained from the labelling. However, we learnt that the evaluator cannot be trusted–the main point of the No-Data Algorithm.
In the case of the evaluator lying about $\sigma$, the explanation is simple. This success rate depends on the cardinality of $\bar{C}$: when it is equal to $\mathcal{C}$, the model will always pass Challenge 1 (and fail Challenge 2). However, when $\lvert\bar{C}\rvert>\lvert C\rvert$, there will be criteria that the evaluator cannot possibly solve without knowing how $C(x)$ is constructed based on its internal structure, and hence it will fail them with some probability. In our experiments, the internal structure depends on both clauses in $c_{2}$.
Finally, the (relatively) low success rate of the evaluator with approximate knowledge of $f$ illustrates one critical characteristic of the No-Data Algorithm: the success rate is a very sensitive metric, highly dependent on both the evaluator’s labelling and generation reliabilities.
10 Conclusion
We introduced an algorithm to determine trustworthiness of an evaluator in the case where labelled data does not exist. It works by performing checks that we showed to be sufficient and necessary to determine (w.h.p.) whether the evaluator knew the labels. Experimentally, the results adjusted to the theoretical predictions, with the No-Data Algorithm being able to determine lying evaluators under various scenarios. Of note was when the evaluator was randomly guessing the labels: while its accuracy was deceptively close to that of its known case, the success rate of the No-Data Algorithm was low, thus flagging that said labelling could not be trusted.
One natural implication of the No-Data Algorithm’s correctness proof is that, when the evaluator cannot be fully trusted, the data cannot be annotated. We argued that this scenario is unsolvable: there is no prior information on the datapoint, and hence it will be effectively creating information out of thin air. This is why the algorithm’s focus is less on the labelling returned, and more on the successes: if an evaluator has low accuracy, but high success rate, it is very likely it is not lying; just wrong (e.g., the IP case in the experiments). Alternatively, high accuracy but low success rate indicates deception (r. the OOP case). The addition of the flip probability $\phi$ and the success count allows for some–limited–heuristics by which the No-Data Algorithm can produce a labelling when the evaluator’s output is rejected. Nonetheless, since the primary purpose of this algorithm is to establish trustworthiness, and not to label the set, it is up to the user to decide (to its own margin of trust) whether to accept or reject the labelling.
As mentioned, some of our work’s limitations are solvable (e.g., label sets of arity $>$ 2), and others are natural (the interpretation of Theorem 1). However, other limitations are more practical in nature and should be the subject of further work. Namely, it is well-known that LLMs are sensitive to prompting. Establishing trust in an LLM evaluator in our setup was equivalent to establishing trust in the prompt, not the model. This nuance could require extensions to the No-Data Algorithm to determine (w.h.p.) whether it is the model or the prompt what is failing to convince the verifier.
11 Limitations
Most of the limitations in the theory are covered in Section 9, and relevant extensions to more common setups are provided in Appendix C. Even when our focus was not experimentation, it could be argued that these were limited in size and scope. We agree–time constraints made us prioritise rigour over experimental results. We have open-sourced the code and data for our work, but also plan to extend it ourselves with other datasets and open models in the near future.
12 Ethics
We argue that the nature of our work precludes any misuse.
Appendix A Detailed Proofs
In this section we provide more careful treatments of the proofs of Lemma 1 (Appendix A.1) and Theorem 1 (Appendix A.4).
A.1 Correctness of the EV protocol
The key in this proof is to note that accepting a label does not mean it will be the correct label, but that the challenges in the EV protocol are sufficient to ascertain said correctness with some probability.
We separate our proof in two: the first part concerns itself with correctness, or whether the challenges will provide sufficient information to determine if the evaluator can be lying. The second part deals with robustness under lying, cheating, or lucky scenarios by the evaluator.
A.1.1 Part 1 (Correctness)
We prove the following lemma:
Lemma 2.
Let $x,x^{\prime}\in X$ be two binary strings such that $x\neq x^{\prime}$. In order to test whether an evaluator knows $f(x)=f(x^{\prime})$ without calling $f$, it is necessary to challenge the evaluator in both $x\equiv x^{\prime}$ and $x\cong x^{\prime}$. Moreover, these conditions are sufficient.
Proof.
(Necessary) It follows from the definition of $C$. Since $f(x)=f(x^{\prime})$, testing $x\equiv x^{\prime}$ is a necessary condition, but not sufficient. This challenge yields information around the internal structure of $x,x^{\prime}$ up to permutation. It is necessary because it is possible to have two $x,x^{\prime}$ such that $x^{\prime}\not\in\{\pi_{i}(x)\}$ (i.e., two strings that aren’t permutations of one another), and yet $C(x^{\prime})=C(x)$. However, this does not prove that x is similar to x’, since there could be a subset $t\in S_{x^{\prime}}.t\not\in S_{x}$ and $c(t)=c(s)$ for some $c\in\mathcal{C}$ (for example, if $c$ contains a disjunctive clause). Hence running this challenge provides information as to whether the encoding of $C(x^{\prime})$ contains all information fed into $\sigma$ for a specific $x,x^{\prime}$. It accounts for variation in internal structure of the string so that $C(x)=C(x^{\prime})$, but does not account for the ordering of these bitstrings.
On the other hand, $x\equiv x^{\prime}$ will determine whether the encoding for the bitstrings lead to the same valuation, by checking strict equality. This is because, under our setup, the definition of similarity for $x,x^{\prime}$ implies $\sigma C(x^{\prime})=\sigma C(x)$. This makes it order-sensitive, but does not account for the internal structure of the string.
Since neither of them provide information about the other, it follows that both are needed to fully determine $f(x)=f(x^{\prime})$.
(Sufficient) It follows directly from the earlier statement. Suppose there exists a pair $x,x^{\prime}$ such that either (or both) $x\not\equiv x^{\prime}$, $x\not\cong x^{\prime}$ and also $f(x)=f(x^{\prime})$. This would imply that (1) either the string , a contradiction.
This concludes the proof. ∎
While a natural consequence of Lemma 2 is that only checking for isomorphism is required for testing equality, we note that the goal of the challenges is to determine if the evaluator is lying, not the label itself.
A.1.2 Part 2 (Robustness)
We prove the following lemma:
Lemma 3.
The probability of error (i.e., the verifier accepts a $y^{\prime}\neq y$) at the $r^{\text{th}}$ round is $\left(1/4\right)^{r}$.
Proof.
In here we have two scenarios: either the evaluator knows $f$, or it is lying or cheating.
Knowing $f$ implies that the evaluator will generate a ‘good’ $x^{\prime}$ every time, such that $\tilde{y^{\prime}}=y$; and hence it will pass the checks from the verifier w.p.1. For the proof we focus then on the second case: that the evaluator is lying or cheating.
There are then two (disjoint) possibilities for a generated $\langle x^{\prime},y^{\prime}\rangle$: either that the evaluator is lying, but gets lucky; or that it is cheating and can generate the answer for one of the challenges. It is worth noting that generating an answer to pass both challenges is equivalent to knowing $f$.
A.2 Lie 1: Generate a ‘good’ $x^{\prime}$
Suppose that the evaluator generates an $x^{\prime}$ such that $S_{x^{\prime}}=S_{x}$, and thus can pass Challenge 1. It cannot pass Challenge 2, since $S_{x^{\prime}}=S_{x}$ implies that either $C(x^{\prime})\neq C(x)$ (and hence it fails it); or $C(x^{\prime})=C(x)$ (and hence either it gets lucky, or it knows $f$). Since the challenges appear with probability $1/2$, the likelihood that the evaluator gets lucky in this round is $1/4$.
A.3 Lie 2: $f(x^{\prime})\neq y^{\prime}$
Now suppose that the evaluator anticipates Challenge 2, and picks an $x^{\prime}$ such that $C(x^{\prime})=C(x)$. Similar to Lie 1, then either $S_{x^{\prime}}\neq S_{x}$ (and thus it fails Challenge 1); or $S_{x}=S_{x^{\prime}}$ (and hence it gets lucky or it knows $f$). The likelihood of getting lucky this round is again $1/4$.
Therefore, the probability of the verifier to not catch the evaluator’s lie is $1/4$ in any round. At the $r^{\text{th}}$ round, $\left(1/4\right)^{r}$. This concludes the proof. ∎
Therefore, from Lemma 2 we know that both challenges together are sufficient to ascertain the truthfulness of a label, and from Lemma 3 that the ability of the verifier to do this is bounded. Putting both lemmas together we obtain the proof for Lemma 1.
A.4 Correctness Bounds of the No-Data Algorithm
By a simple application of the union bound on the probability that the evaluator was wrong, but not detected by the verifier, and the probability that the evaluator was wrong, detected by the verifier, but without having its label flipped.
For this, it is much easier to work with the errors $\epsilon=1-\alpha$, and remark that, from Lemma 1 the first term (undetected, no flip) is given by
$\Pr[\text{undetected}]$ $=\sum_{f(x)\neq y}\left(\frac{1}{4}\right)^{r}\text{, and hence}$ (4)
$\mathrm{E}[\text{undetected}]$ $=\epsilon\left(\frac{1}{4}\right)^{r}.$ (5)
Likewise, the second term accounts only for the situation where the mislabelling was detected and the label was not flipped:
$\mathrm{E}[\text{detected, no flip}]=\epsilon\left(1-\phi\right)\left(1-\left(\frac{1}{4}\right)^{r}\right).$ (6)
This yields:
[…]
$\leq\left(\frac{1}{4}\right)^{r}+\left(1-\phi\right)\left(1-\left(\frac{1}{4}\right)^{r}\right)$ (8)
$\mathrm{E}[\text{wrong}]$ $\leq\epsilon\left(1-\phi+\phi\left(\frac{1}{4}\right)^{r}\right)\text{, by linearity of expectation.}$. (9)
Substituting back $\Pr[\text{wrong}]=1-\Pr[\text{right}]$ and $\epsilon=1-\alpha$ yields the desired value. This concludes the proof.
Appendix B Characterisation Bounds
Lemma 4.
Suppose the evaluator and verifier run in polynomial time for any input $x\in X$. Then, if the No-Data Algorithm is ran over a dataset $D$; and with $r$ rounds of EV protocol, the runtime is then $\operatorname{O}\!\left(r\lvert D\rvert\right)$.
Proof.
Straightforward. ∎
Appendix C Extensions
C.1 Distance-based Constraints
The constraint that $C(x)=C(x^{\prime})$ is too strict for some applications. We sketch out in this section a variant of the No-Data Algorithm where, instead of requiring equality, we request closeness with respect to some arbitrary metric $\delta$ in Challenge 2. Suppose now that the failure is given whenever $\delta(C(x),C(x^{\prime}))\geq\epsilon$, for some $\epsilon\in[0,1/2]$. The bounds of correctness are roughly equivalent to these described in Lemma 1 and Theorem 1. However, the term $(1/4)^{r}$ now takes a dependency on the probability of generating a string that is far from the desired constraint.
C.2 $k$-ary Label Sets
To convert a dataset of arity $k>2$, it is necessary to rerun the No-Data algorithm $k-1$ times. On each run, the label set $\{l_{1},\dots,l_{k}\}$ is mapped as $\{l_{1}\},\{l_{2},\dots,l_{k}\}\mapsto\{0\},\{1\}$. In these cases, the runtime of the algorithm is increased by a factor of $k-1$.
Appendix D Detailed Experimental Methods
D.1 Detailed Setup
For the LLM we used o3-mini (version 2025-01-31), and called it via the Azure OpenAI API. We set the maximum tokens to 50,000 and left all other parameters as default. For the decision tree we used the implementation from sklearn [^3] https://scikit-learn.org/stable/index.html with the default parameters. All data analysis was performed on a consumer-grade laptop. Random seeds are set for the training in the code.
D.2 Prompts
The prompts we used for our LLM stack are in Prompt 1 (evaluation) and Prompt 2 (generation). All exemplars are omitted for brevity. Whenever a parsing error occurred, we retried the call up to five times. If it failed on the fifth, we returned a random string.
You are labelling binary strings based on a rubric (given below).
First return the parts of the criterion that match the string, and the values.
Then return the label based on the aggregate function.
# Rubric:
<NL RUBRIC GOES HERE>
# Aggregation Function:
Majority vote
Return your answer in the form:
|reasons|
(list of reasons)
|reasons|
|label|
(the label)
|label|
Table 1: System prompt for evaluation. Anchors (‘|reasons|’, ‘|label|’) are used for parsing. The text contained within ‘|reasons|’ is, as specified by the exemplars (not pictured) a chain-of-thought approach determining criterion-by-criterion matching. The rubric, explained in natural language, is inserted in the marker. Throughout the experiments we only used the IP rubric.
You are given five datapoints and a set of criteria.
Based on these, pick the datapoint that matches the criteria exactly.
# Rubric:
<NL RUBRIC GOES HERE>
# Datapoints:
<DATAPOINTS GO HERE>
Return your rationale, and then your chosen datapoint ONLY in the form:
|datapoint|
(the datapoint)
|datapoint|
Table 2: System prompt for generation. This prompt does not utilise exemplars. The rubric, explained in natural language, is inserted in the marker. Throughout the experiments we only used the IP rubric and selected datapoints from the IP training dataset, ensuring that at least one of the entries matched the criterion. Given that the model did not follow the output format exactly, some trial-and-error was needed to ensure a comparatively high parse success rate.
You are a datapoint generator over binary strings.
Given a rubric (given below), a datapoint, and a label, return a similar datapoint that has the same
label, and fulfils the same conditions as the rubric.
For convenience, always start with the same datapoint: 0100. It will be easier to work with.
# Rubric:
<NL RUBRIC GOES HERE>
Return your rationale, and then the final datapoint in the form:
|datapoint|
(the datapoint)
|datapoint|
Table 3: System prompt for generation used in the ablation study from Section 8.1.2. Unlike Prompt 2, this prompt does take exemplars (not pictured): the first five entries in the training dataset, along with chain-of-thought-style rationale.

### v3

@@
 We show that this is sufficient to establish trustworthiness w.h.p., in such a way that when the evaluator actually knows the way to label the corpus, the No-Data Algorithm accepts its output; and, conversely, flags untrustworthy evaluators when these are unable to prove it.
-We present formal proofs of correctness and limited experiments.
+We present formal proofs of correctness, empirical tests, and applications to LLMs-as-judges on low-resource languages.
 1 Introduction
-One of the most fundamental problems in machine learning, and AI broadly, is measurement.
-It has been known for a while that some metrics have a certain level of unreliability; for example, by not corresponding to human judgements appropriately [34, 11, 3, 13].
-However, the rise of LLMs as evaluators, typically known as LLMs-as-judges, has led to a substantial divide in the field regarding their trustworthiness as a measurement tool.
-Their use has also led to broader questions around measurement in AI; specifically, about the validity of some results [11, 33, 25].
-A key component in the ‘measurement in AI question’, and the object of study of our work, is whether the tool providing the evaluation (the evaluator) may be trusted.
-This has become more critical given the fast adoption of foundation models in industry and scientific research, and the continuous need to scale.
-In these cases, trust is usually established based on some sort of ground truth: an existing set of labelled references by which to perform statistical analyses (e.g., a $t$-test).
-However, it is often the case that said references do not necessarily exist: perhaps the problem is too niche, or the annotation too costly.
-It follows then that the only way one can establish trust on an unknown evaluator in a no-reference scenario is via formal methods (i.e., formal proofs).
-In this paper we introduce an algorithm (which we call the ‘No-Data Algorithm’) by which to establish trust in an evaluator without previously-labelled data.
-The algorithm is designed to work in-line with contemporary machine learning techniques, where an evaluator claims that it knows the map between datapoints and labels.
+One of the most fundamental problems in AI is the question of measurement.
+It has been known for a while that some metrics have a certain level of unreliability; for example, by not corresponding to human judgements appropriately (Reiter, 2018; Gehrmann et al., 2023; Bavaresco et al., 2024; Hada et al., 2024).
+With the rise of LLMs as evaluators, typically known as LLMs-as-judges, scaling measurement to multiple subproblems has become ubiquitous (de Wynter, 2025).
+However, their reliability and trustworthiness as a measurement tool has led to a substantial divide in the research community, and specifically around the validity of some results (Gehrmann et al., 2023; Rabinovich and Anaby Tavor, 2025; Moghe et al., 2023).
+A key component of the question of measurement, and the focus of our work, is whether the tool providing the evaluation (the evaluator) may be trusted.
+This trust is typically established with some sort of ground truth, such as an existing set of labelled references with which to perform statistical analyses (e.g., a $t$-test), or a benchmark with previously-published results.
+Due to uniqueness of the problem, cost of annotation, or other reasons, references may not readily exist, or be too scarce to be statistically significant.
+Likewise, contamination of public benchmarks has become a major concern (see, e.g., Sainz et al.
+2023).
+It follows then that the only way one can scientifically establish trust on an evaluator whose capabilities are unknown, in a no-reference scenario, is with a formal proof.
+In this paper we introduce an algorithm (the ‘No-Data Algorithm’) by which to establish trust in an evaluator without previously-labelled data.
+The algorithm is designed to work in-line with contemporary machine learning techniques–namely, LLMs-as-judges–where an evaluator claims it knows the map between datapoints and labels.
 It is up to a verifier to decide, based on the outcome of certain challenges, whether the evaluator may be trusted.
 The No-Data Algorithm requires a linear number of calls and does not require labelled datasets (e.g., development or training sets) or knowledge of the labels in the test set to establish trust.
-We show that our algorithm, after $r$ calls, bounds confidence on the labeller to $(1/4)^{r}$.
-This success rate reflects the ‘true’ accuracy of the evaluator: if the evaluator knows how to label the set, the accuracy and success rates will be high; if the evaluator lies about knowing the labels, the accuracy could be high, but the success rate won’t.
-We present small experiments in synthetic data for traditional machine-learning (i.e., a decision tree) and contemporary (r.
-an LLM) methods which adjust to the theory’s predictions, although we note our contribution is theoretical in nature.
+We show that our algorithm, after $r$ calls, bounds confidence on the evaluator to a $(1/4)^{r}$ probability of error.
+This success rate reflects the evaluator’s knowledge–or lack thereof: if it knows how to label the set, the accuracy and success rate, as reported by the No-Data Algorithm, will be high.
+Conversely, if the evaluator lies about knowing the labels, the accuracy could be high–perhaps by luck–but the success rate will not.
+We present empirical tests with traditional machine-learning (i.e., a decision tree) and contemporary (r.
+LLMs) evaluation methods.
+We also showcase an application of the No-Data Algorithm by judging an LLM-as-a-judge’s ability to label a dataset in West Frisian, a low-resource language.
+Both experiments adjust to the theory’s predictions.
 [^1] Code and data may be found in https://github.com/adewynter/no_data_algorithm.
-Amongst our experimental findings we note that the behaviour of a reliable (read: trained) evaluator adjusts to our theoretical predictions, while an LLM requires more work due to prompting and spurious reasoning.
-Nonetheless, when properly tuned, it adjusts to the predictions.
-Our results demonstrate that trust in an evaluator may be established via formal methods without the need for references.
-A side effect of our experimentation opens a wide range of questions around faithful generation over arbitrary strings for a generative model.
+Our results demonstrate that trust in an evaluator may be established via formal methods without the need for references, and enables their application in scenarios where labels may be extremely scarce, such as market research, low-resource languages (the subject of our experiments), and the medical domain.
 2 Related Work
@@
 Most of contemporary work relies on LLMs and various prompting strategies or call stacks, although the problem of (un)reliability/trustworthiness of an evaluator is older than generative models.
-For example, it has been known that the use of BLEU for tasks other than English-based machine translation is not accurate [22, 28]; and even then, that it does not correspond well to human judgements [34, 11].
-More generally, it is known that metric choice, and hence the choice of evaluator, is system-dependent [4, 25, 10, 39, 32].
-Poorly-developed rubrics also lead to unreliable human judgements [7, 20], and thus the evaluator trustworthiness problem is also a rubric problem.
-There are algorithms [27, 47] and paradigms [26] for learning with noisy teachers.
-Likewise, provable methods have been used with success in other areas, such as representation learning [15], and to improve robustness via synthetic data generation [8, 9].
+For example, it has been known that the use of BLEU for tasks other than English-based machine translation is not accurate (Liu et al., 2016; Novikova et al., 2017); and even then, that it does not correspond well to human judgements (Reiter, 2018; Gehrmann et al., 2023).
+More generally, it is known that metric choice, and hence the choice of evaluator, is system-dependent (Chen et al., 2024; Moghe et al., 2023; Flamich et al., 2025; von Däniken et al., 2024; Pombal et al., 2025).
+Poorly-developed rubrics also lead to unreliable human judgements (Clark et al., 2021; van der Lee et al., 2019), and thus the evaluator trustworthiness problem is also a rubric problem.
+There are algorithms (Northcutt et al., 2017; Zhang and Sabuncu, 2018) and paradigms (Northcutt et al., 2021) for learning with noisy teachers.
+Likewise, provable methods have been used with success in other areas, such as representation learning (Jovanović et al., 2023), and to improve robustness via synthetic data generation (Dimitrov et al., 2022; Fischer et al., 2022).
 However, they all assume that a (perhaps noisily) labelled dataset exists.
-Relevant to our work, the use of LLMs as evaluators has promised extraordinary scaling capabilities, such as fast data generation and evaluation without humans [29, 6, 5, 23, 41, 48].
-Some of these criticisms are that they do not correlate with human judgements well [3, 13, 4, 45, 37]; are sensitive to their prompt [24, 14, 46]; their performance could be the product of memorisation [19, 44, 35]; and even that they do not understand the task at all [43, 40, 46].
-Their reasoning capabilities have been put into question, from the results being strongly dependent on the choice of metrics [36, 4], to findings that their output reasoning steps contain spurious reasoning [38, 18, 42].
-Most of the work on improving trust, however, relies on prompting strategies or call stacks to other LLMs [21].
-The main subroutine from the No-Data Algorithm, which we call the Evaluator-Verifier (EV) protocol is based off the well-known Arthur-Merlin (AM) protocol from Babai [2].
-It is a type of zero-knowledge proof; see [12] and [1] for primers on the subject.
+Relevant to our work, the use of LLMs as evaluators has promised extraordinary scaling capabilities, such as fast data generation and evaluation without humans (Open AI, 2023; Chiang et al., 2024; Chiang and Lee, 2023; Liu et al., 2023; Wei et al., 2024; Zheng et al., 2023).
+Some of these criticisms are that they do not correlate with human judgements well (Bavaresco et al., 2024; Hada et al., 2024; Chen et al., 2024; de Wynter et al., 2025; Tjuatja et al., 2024); are sensitive to their prompt (Lu et al., 2022; Hida et al., 2024; Ye and Durrett, 2022); their performance could be the product of memorisation (Lee et al., 2023; de Wynter et al., 2023; Sainz et al., 2023); and even that they do not understand the task at all (de Wynter and Yuan, 2024; Webson and Pavlick, 2022; Ye and Durrett, 2022).
+Their reasoning capabilities have been put into question, from the results being strongly dependent on the choice of metrics (Schaeffer et al., 2023; Chen et al., 2024), to findings that their output reasoning steps contain spurious reasoning (Turpin et al., 2023; Lanham et al., 2023; Wu et al., 2024).
+Most of the work on improving trust, however, relies on prompting strategies or call stacks to other LLMs (Li et al., 2024).
+The main subroutine from the No-Data Algorithm, which we call the Evaluator-Verifier (EV) protocol is based off the well-known Arthur-Merlin (AM) protocol from Babai (1985).
+It is a type of zero-knowledge proof; see Goldreich (2009) and Arora and Barak (2009) for primers on the subject.
 That said, the EV protocol is different in many regards to the AM protocol, since it is designed to fit arbitrary inputs and the larger No-Data Algorithm.
@@
 Likewise, our use of the rubric and the aggregator functions are generalisations to partial functions that are not present (or needed) in previous works.
-3 Background
-4 Notation
+3 Definitions and Problem Setup
+3.1 Notation
 Our work relates to datapoints $x\in X$ and labels $y\in Y$.
@@
 The assumptions on $X$ and $Y$ are to simplify our proofs; see Section 7 for a discussion on their relevancy.
-Likewise, for a binary string $x\in\{0,1\}^{n}$, we denote its relevant subset set as $S_{x}\subset 2^{x}$.
+For a binary string $x\in\{0,1\}^{n}$, we denote its relevant subset set as $S_{x}\subset 2^{x}$.
 The construction of this specific set will depend on the setup, and will be discussed in detail in Section 4.1.
@@
 We assume all functions to be deterministic throughout, unless stated otherwise.
-5 Definitions and Problem Setup
-In line with standard learning theory, we assume that there exists an unknown map $f\colon X\rightarrow Y$ that takes datapoints to labels faithfully–that is, $f$ provides the true labelling.
+3.2 Definitions
+In line with standard learning theory (Valiant, 1984), we assume that there exists an unknown map $f\colon X\rightarrow Y$ that takes datapoints to labels faithfully–that is, $f$ provides the true labelling.
 Our task is to determine whether a function $E\colon X\rightarrow Y$ (a labeller, or evaluator) is equivalent to $f$, without any knowledge of the map itself.
-For that, we introduce four definitions:
-A criterion $c\colon X\rightarrow\{0,1\}$ is a map that takes in a datapoint $X$ and maps it to a single bit.
-The rubric $C\colon\times_{i,\dots,n}\{0,1\}\rightarrow\{0,1\}^{n}$ takes in the binary outputs from an ordered set of criteria $\mathcal{C}=\{c_{1},\dots,c_{n}\}$ and returns a bitstring that is the concatenation of these outputs: $C=c_{1}|c_{2}|\dots|c_{n}$.
-We say a rubric is total if it explicitly decomposes nonlinear criteria with arity $>1$ (e.g.
-the xor $c_{i}=c_{a}\oplus c_{b}$), and denote it as $\bar{\mathcal{C}}$.
+This is measured with its error rate, or, conversely, its accuracy.
+We also introduce four definitions which we will use in our work:
+A criterion $c\colon X\rightarrow\{0,1\}$ maps a datapoint $X$ to a single bit.
+A rubric $C\colon\times_{i,\dots,n}\{0,1\}\rightarrow\{0,1\}^{n}$ maps the binary outputs from an ordered set of criteria $\mathcal{C}=\{c_{1},\dots,c_{n}\}$ to a bitstring that is the concatenation of these outputs: $C=c_{1}|c_{2}|\dots|c_{n}$.
+We say a rubric is total if it explicitly decomposes nonlinear criteria with arity >1>1 (e.g.
+the xor $c_{i}=c_{a}\oplus c_{b}$), and write it as $\bar{\mathcal{C}}$.
 When the evaluation of said criteria can be done separately (e.g., by testing first $c_{a}$, then $c_{b}$, and then their xor $c_{a}\oplus c_{b}$), we say it is a total evaluation.
-The aggregator $\sigma\colon\{0,1\}^{n}\rightarrow Y$ is a function that takes in a binary string of length $n$ (the output from the rubric) and maps it to the label set.
+The aggregator $\sigma\colon\{0,1\}^{n}\rightarrow Y$ is a function that maps $n$-bit binary strings (rubric outputs) to $Y$.
 Finally, for two $x,x^{\prime}\in X$, we say that $x$ is similar to $x^{\prime}$ if both $x\cong x^{\prime}$ and $x\equiv x^{\prime}$.
 Intuitively, the rubric is the set of reasons why a datapoint $x$ would take a given label (for example, an annotation rubric passed in to human labellers, the last layer in a model, etc.); while the aggregator decides the final label based on that rubric (e.g., via an activation function, a majority vote, etc.).
+Natural-language examples are in Appendix E.
 We are now ready to introduce the core assumption of our work:
 Assumption 1.
-The map $f\colon X\rightarrow Y$ is a composition of the rubric and the aggregator. That is,
+The map $f\colon X\rightarrow Y$ is a composition of the rubric and the aggregator:
 $f=\sigma C.$ (1)
-While not a particular strong assumption (after all, datapoints must have a reason to acquire a label and, when there is more than one reason, there must be a way to decide how to aggregate them) this decomposition is crucial to our work.
+It is not a strong assumption: after all, datapoints must have a reason to acquire a label.
+When there is more than one reason, there must be a way to decide how to aggregate them.
+However, this decomposition is crucial to our work.
 4 The No-Data Algorithm
-The No-Data Algorithm is an algorithm with two components (players): the evaluator and the verifier.
-[^2] In theoretical computer science, the evaluator is typically known as the prover.
-We use this more-intuitive naming to highlight the distinction between our work (data-labelling) and others (e.g.
-theorem-proving).
-Throughout the algorithm, for every datapoint $x\in X$, the evaluator must convince the verifier that it is able to label $x$ correctly (i.e., that it knows $f$).
+The No-Data Algorithm is an algorithm with two components (players): the evaluator and the verifier.Throughout the algorithm, for every datapoint $x\in X$, the evaluator must convince the verifier that it is able to label $x$ correctly (i.e., that it knows $f$).
 For this, the players run a multi-round sub-game which we refer to as the Evaluator-Verifier (EV) protocol.
-The EV protocol takes in $x$, and either succeeds or ends in failure, w.h.p., based on challenges posed by the verifier on a $x^{\prime}$ similar to $x$ generated by the evaluator.
-It then returns the proposed label and the status (success or failure).
-The algorithm then, in case of failure, decides to flip the label with probability $\phi$.
-At the end of the loop, it aggregates the results of the EV protocol and returns the final labelling and success count.
-The final labelling depends on the choice of $\phi$, and the evaluator’s trustworthiness (and hence that of said labelling) is be determined by measuring the success count.
-The full algorithm may be found in Algorithm 1.
-Given that the EV protocol is a fundamental part of the No-Data Algorithm, we focus the remainder of this section to describe it, and provide a characterisation of the No-Data Algorithm (bounds, correctness, etc.) in Section 7.1.
+The EV protocol takes in $x$, and either succeeds or fails, w.h.p., based on challenges posed by the verifier on a $x^{\prime}$ similar to $x$ generated by the evaluator.
+The EV protocol returns the proposed label and the status (success or failure).
+The No-Data Algorithm then flips the label with probability $\phi$ in case of failure.
+After observing every $x\in X$, it returns the success count and final labelling.
+The evaluator’s trustworthiness (and hence the final labelling’s) is established with the success count.
+The full algorithm is in Algorithm 1. We and provide a characterisation of the No-Data Algorithm (bounds, correctness, etc.) in Section 5.
 1:Input: Unlabelled data $X$, evaluator $E$, verifier $V$, criterion $C$, flip probability $\phi$, rounds $r$
-2:  $predictions\leftarrow\{\}$
-3:  $successes\leftarrow\{\}$
+[…]
+[…]
 4:for $x\in X$ do
 5:  $success,y\leftarrow\textsc{EV}(x,E,V,C,r)$// Run for $r$ rounds
-6:     if $success$ then
+6:  if successsuccess then
 7:   $predictions\leftarrow predictions\cup\{y\}$
@@
 14:end for
-15:  return  $predictions$, $successes$
+15:return  predictionspredictions, successessuccesses
 Algorithm 1 The No-Data Algorithm.
-It works by calling a subroutine (the EV Protocol) to provably determine the trustworthiness of the labels produced by the evaluator.
-Based on its output (success or failure), the algorithm then determines which label to assign to the corpus, with a stochastic label flip $\phi$ added to it in case of failures.
+For every $x\in X$ it calls the EV protocol to determine the trustworthiness of the label returned by the evaluator, $y=E(x)$.
+Based on its output, the algorithm flips $y$ w.p.
+$\phi$ in case of failure.
+Finally, it returns the success count and the labels.
 4.1 The Evaluator-Verifier (EV) Protocol
-In the EV protocol both players are given a datapoint $x\in X$, and the rubric $C$, but nothing else.
+In the EV protocol both players are given a datapoint $x\in X$, rubric $C$, and nothing else.
 In particular, $\sigma$ is not explicitly given to either player.
-However, the evaluator claims to know the map $f\colon X\rightarrow Y$.
-Since $C$ is given, then the claim is that it has prior knowledge of $\sigma$.
-The implication–were the evaluator to convince the verifier–then is that its choice of label $y$ for $x$ may be trusted.
-Remark that the verifier does not know if $y$ is correct.
-In order to be convinced, the verifier chooses one out of two challenges, neither of which rely on $\sigma$ or $y$.
-Based on whether the evaluator passes or fails them, the verifier builds its own belief on the evaluator’s trustworthiness.
-The entire point of the EV protocol (and thus of the No-Data Algorithm) is for the evaluator to convince the verifier that its choice of label for $x$ is correct.
-If the evaluator fails, it follows that the correct label for $x$ could be the opposite label.
-Hence, regardless of success, at the end of the EV protocol the evaluator reveals its label.
-See Figure 1 for a diagram of the EV protocol.
-We now describe each of the components and flow.
+The evaluator claims to know the map $f\colon X\rightarrow Y$.
+Since $C$ is given, the claim is then that it has prior knowledge of $\sigma$.
+The verifier’s goal is to be sufficiently sure that the evaluator’s choice of label $x$ for a given $x\in X$ may be trusted.
 4.1.1 Setup
-At every round of the EV protocol, the evaluator takes in $x$ and generates a similar datapoint $x^{\prime}$ and a ‘partial’ label $\tilde{y}^{\prime}$ that–ideally–is equal to $y$.
-It is partial because there is no access to $f$, and hence the best the evaluator can do is to generate a guess $\tilde{y}^{\prime}$ based on the encoding $C(x^{\prime})$.
-If the evaluator indeed knows $f$, an application of $\sigma$ to $C(x^{\prime})$ would lead to the ‘true’ label $y$.
-That said, once it has generated the tuple $\langle x^{\prime},\tilde{y}^{\prime}\rangle$, the evaluator may not change them in this round.
-Remark that isomorphism is not a sufficient condition for $\tilde{y}^{\prime}=y$ since $C$ might not be total.
-Hence, both structural and valuation checks are needed.
-Therefore, after the tuple has been generated, the verifier then chooses uniformly at random one of the two following challenges:
-Figure 1: Diagram of the EV protocol.
-At every round, the evaluator (blue) generates a datapoint $x^{\prime}$ similar to $x$, and a partial label $\tilde{y}^{\prime}$.
-It then responds to one of two (chosen uniformly at random) challenges by the verifier (orange).
-The game is repeated by $r$ rounds and returns a succeeded state, unless there is a failure in the challenges, and thus returns a failed state.
-In either case, it also returns the proposed partial label, $\tilde{y}^{\prime}$.
+The EV protocol is played for $r$ rounds.
+At every round, the evaluator generates a similar datapoint $x^{\prime}$ to $x$, and a ‘partial’ label $\tilde{y}^{\prime}$ that–supposedly–is equal to $y$.
+It is partial because there is no access to $f$, and hence the best the evaluator can do is to guess $\tilde{y}^{\prime}$ based on $C(x^{\prime})$.
+Once it has generated the tuple $\langle x^{\prime},\tilde{y}^{\prime}\rangle$, the evaluator may not change it in this round.
+The verifier poses one of two challenges, selected uniformly at random.
+Either challenge measures equality between $x^{\prime}$ and $x$, either up to isomorphism or up to permutation.
+If the evaluator fails, the protocol terminates and returns failure.
+Otherwise, it starts again.
+Either in success or failure, the protocol also returns $\tilde{y}^{\prime}$.
+See Figure 1 for a diagram of the flow of the EV protocol.
+The verifier challenges are central to the work, and perform both structural and valuation checks:
+Figure 1: EV protocol flow.
+At every round, the evaluator (blue) generates an $x^{\prime}$ similar to $x$, and a partial label $\tilde{y}^{\prime}$.
+It then answers one of two (chosen uniformly at random) challenges by the verifier (orange).
+If the evaluator does not pass the challenge, the protocol returns failure.
+Otherwise, the game is repeated.
+If the rounds are over, it returns a succeeded state.
+In either case, it also returns $\tilde{y}^{\prime}$.
 Challenge 1 (Equality up to Permutation)
-Setup: The verifier tests the evaluator by asking it to show a similar point that leads to the same total evaluation of $C$.
-This challenge is surrogated by testing whether $x\equiv x^{\prime}$ over $\bar{\mathcal{C}}$.
+Setup: The verifier tests the evaluator by asking it to show a similar point that leads to the same total evaluation of $C$, namely, whether $x\equiv x^{\prime}$ over $\bar{\mathcal{C}}$.
 Check: The verifier asserts that
 $\forall s\in S_{x}\exists t\in S_{x^{\prime}}\,.\,\forall c\in\bar{\mathcal{C}}\,c(s)=c(t);$ (2)
-namely, whether there exists at least one substring in $x^{\prime}$ matching the (total) evaluation of a criterion from $x$.
+that is, if there is at least one substring in $x^{\prime}$ matching the (total) evaluation of a criterion from $x$.
 Challenge 2 (Equality up to Isomorphism)
-Setup: The verifier asks for the rationale as to why $f(x^{\prime})=y$. Since $y$ is not available, it checks whether $x^{\prime}\cong x$.
+Setup: The verifier asks why $f(x^{\prime})=y$. Since $y$ is not available, it checks whether $x^{\prime}\cong x$.
 Check: The verifier asserts that $C(x^{\prime})=C(x)$.
-6.1.2 Verification
-If the evaluator failed the tests, the EV Protocol returns a failure state.
-Otherwise, it restarts from Section 6.1.1 over and over until the maximum number of rounds $r$ has been reached, or the evaluator fails a test.
-In either case, at the end of the call the EV Protocol also returns $\tilde{y}^{\prime}$.
 Remark 1.
@@
 This is why the challenges assert both structure ($x\equiv x^{\prime}$) and final encoding ($x\cong x^{\prime}$).
-This observation will be elaborated further in our correctness proofs ( Section 7.1).
-Remark 2.
-The check for Challenge 1 (demanding equality up to criterion) is very strict. We discuss this in Section 9 and present a relaxed extension in Appendix C.
-7 Algorithm Characterisation
-In this section we provide proofs of correctness (Section 7.1). Runtime bounds may be found in Appendix B.
-7.1 Correctness
+This observation is central to our correctness proofs (Section 5).
+5 Characterisation
 Lemma 1 (EV Protocol Correctness Bound).
@@
 ∎
-Remark 3.
-The exponential bound from Lemma 1 means that, for example, after running the EV protocol for 3 rounds, the probability that the verifier fails to detect a lie from the evaluator is 1.6%.
 Theorem 1 (No-Data Algorithm Correctness Bound).
@@
 ∎
-The implication of Theorem 1 is that a good evaluator with low probability to flip the labels will obtain a good label accuracy.
-A bad evaluator with high probability to flip the labels will also obtain a good accuracy.
-Another interpretation of this, however, is that there is no way to recover unknowable labels when the evaluator is unable to convince the verifier–we discuss this further in Section 9.
+The implication of Theorem 1 is that a good evaluator with relatively high accuracy and success rate, and a well-calibrated $\phi$, will have a true accuracy higher than what the No-Data Algorithm shows.
+This means that reliable evaluators, as per the algorithm, may be used with the expectation that their ‘true’ performance will be better.
+Another interpretation of this result is that there is no way to recover unknowable labels when the evaluator is unable to convince the verifier.
 It is worth mentioning that the assumption for Theorem 1 around always lying when mislabelling (and hence always being wrong) are not necessarily applicable to all evaluators, since a particularly tricky evaluator could still guess the right label and yet not pass the challenges.
-We consider this out of scope, since the ability to guess a label without explaining how is not a viable way to establish trust in scenarios without any references.
-Consequentially, for the No-Data Algorithm it is more helpful to look at the success count.
+We consider this out of scope, since the ability to guess a label without explaining how is not a viable way to establish trust, especially in scenarios without any references.
 6 Experiments
-We describe the two experiments and results in this section and refer the reader to Appendix D for a more detailed treatment, including prompts and specifics on the models used.
-Throughout this section, we run the No-Data Algorithm with $r=3$.
-8.0.1 Data
-We created two mostly disjoint synthetic datasets so that we could control the labels, data, and their rubrics.
-We refer to these datasets as in-phenomenon, or IP; and out-of-phenomenon, or OOP, as opposed to in-distribution (r.
-out-of-distribution) to highlight the fact that the entire measurement is altered by the change in data–and, that, semantically, we are modelling different phenomena.
-The IP dataset is our baseline: the evaluator will not be lying on knowing this phenomenon.
-The OOP dataset is our reference for a corpus that the evaluator has not seen before (and thus it will lie about knowing it).
-For both problems, $\sigma$ is the majority vote.
-We chose to generate our own data to ablate out any potential memorisation concerns, and to control the rubrics.
-These may be found in Table 1.
+We present two experiments: an empirical test of the theory (Section 6.1), and an application to a realistic scenario (Section 6.2).
+Ablation experiments are in Appendix D.
+We used two corpora: the data it knows how to label (in-phenomenon, or IP), and one it doesn’t (out-of-phenomenon; OOP).
+We use this naming to highlight the fact that the entire measurement is altered by the change in data–and, that, semantically, we are modelling different phenomena (i.e., distinct $f\colon X\rightarrow Y$).
+We refer to the hypothetical setup where we know the labels of a dataset as knowable (r.
+unknowable).
+Knowable scenarios are only to illustrate how the algorithm’s labelling adheres to the ‘true’ predictions, and in practice they would be unavailable.
+We run the No-Data Algorithm with $r=3$.
+See Appendix F for prompts and specifics on the models and methodology used.
+6.1 Experiment 1: Synthetic Data without Natural Language
+In this experiment we aimed to understand the behaviour of the No-Data Algorithm with a corpus close to the theoretical setup.
+We created two disjoint synthetic datasets over binary strings.
+This had the advantage of ablating out any potential memorisation concerns, and allowed us to control the rubrics, as shown in Rubric 1.
+For both OOP and IP, $\sigma$ is the majority vote.
 The test sets for both IP and OOP included 498 entries and are balanced.
-8.0.2 Evaluator
-We used a decision tree (DT) for our evaluation, and an LLM (o3-mini; [31]).
-Both models observed a balanced training split of the IP dataset: the DT was trained on it, while the LLM observed five exemplars in the prompt.
-The rationale to train the DT was to fully ensure that the evaluator learnt both $\sigma$ and $C$.
-For generation it only had access to the rubric, but not $\sigma$.
-The LLM observed it via the prompt, and was made aware of the aggregation function only for the labelling step.
-For generation it was asked to pick the best candidate from the dataset, as opposed to outputting a new point, as we observed marked performance decreases otherwise.
-During OOP, the LLM was given the rubric for IP and an OOP datapoint; while the DT was trained in IP (and, ideally, learnt the rubric), and classified IP and OOP datapoints.
+6.1.1 Setup: Evaluators and Verifiers
+Evaluators:
+We used a decision tree (DT), and the highest-performing LLM we tested (o3-mini; OpenAI 2025a).
+[^2] The full results, in Appendix D, further include another closed- and two open-source models.
+Both observed a balanced training split of the IP dataset (DT was trained in it; the LLM observed it with five exemplars).
+We trained the DT to ensure that the evaluator learnt both IP’s $\sigma$ and $C$, even though during generation it only had access to $C$, and not $\sigma$.
+As ‘training’, the LLM observed the rubric in the prompt, and was made aware of the aggregation function only for the labelling step.
+For generation it was asked to pick the best candidate from the (training) dataset instead of outputting a new point, as we observed marked performance decreases otherwise (Appendix D).
+During OOP, the LLM was given the rubric for IP and an OOP datapoint.
+We tuned $\phi$ based on the knowable scenarios, to $\phi=0.4,0.1$ for the DT and the LLM, respectively.
+Verifier:
+We used a rule-matching algorithm that has access to the relevant rubrics (e.g., when testing IP, it observes the rubric for IP), implemented as described in Section 4.1.
 Rubric Definition
 In-Phenomenon
-$c_{0}$ $y=1$ if $x$ has an even number of ones; else 00
-$c_{1}$ $y=1$ if $x$ starts with a zero or contains $10101$, but not both; else 00
-$c_{2}$ $y=1$ if $x$ has strictly more than five ones; else 00
+$c_{0}$ $y=1$ if $x$ has an even number of ones; else 0
+$c_{1}$ $y=1$ if $x$ starts with a zero or contains $10101$, but not both; else 0
+$c_{2}$ $y=1$ if $x$ has strictly more than five ones; else 0
 Out-of-Phenomenon
-$c_{0}$ $y=1$ if $x$ contains the substring $111$; else 00
-$c_{1}$ $y=1$ if $x$ ends with a one; else 00
-$c_{2}$ $y=1$ if $x$ contains the substring $110001$; else 00
-Table 1: Rubrics and criteria for our experiments.
-For both problems, $\sigma$ was the majority vote.
-Remark that $c_{1}$ for the IP rubric is an exclusive-or operation.
-This means that, while this rubric contains three criteria, the total rubric contains five: $\{c_{0},c_{2},c_{1,a},c_{1,b},c_{1}\}$; where $c_{1,a},c_{1,b}$ are the clauses for $c_{1}$, and $c_{1}$ is the exclusive-or itself.
-8.0.3 Verifier
-The verifier is a rule-matching algorithm that has access to the relevant rubrics (e.g., when testing IP, it observes the rubric for IP), and works as described in Section 6.1.
-8.1 Results
+$c_{0}$ $y=1$ if $x$ contains the substring $111$; else 0
+$c_{1}$ $y=1$ if $x$ ends with a one; else 0
+$c_{2}$ $y=1$ if $x$ contains the substring $110001$; else 0
+Table 1: Rubrics and criteria for our first experiment.
+Note that $c_{1}$ for the IP rubric is an exclusive-or operation: while this rubric contains three criteria, the total rubric contains five: $\{c_{0},c_{2},c_{1,a},c_{1,b},c_{1}\}$; where $c_{1,a},c_{1,b}$ are the clauses for $c_{1}$, and $c_{1}$ is the exclusive-or itself.
+6.1.2 Results
 The results are in Table 1.
-Both the DT and the LLM adjusted to the predictions correctly, successfully convincing the verifier when it was an oracular generator, reaching approximately their knowable accuracy, and outputting low success rates when lying.
-Indeed, the accuracy of the models in the OOP dataset hovers around 54%–close to random guessing.
-A similar accuracy can be obtained by the output of the No-Data Algorithm, but the success rates between IP (100%) and OOP (5%) provide better information as to in which situations the evaluator is trustworthy.
-The tuned $\phi$ value, based on the known labels, also yielded accuracies close to the expected score from the known dataset.
+The DT and the LLM adjusted to the predictions: accuracies with the No-Data Algorithm were within 2-2 and 0% of the knowable case in both IP and OOP.
+The success rate differences between IP (100% for DT; 81% for the LLM) and OOP (r.
+5% and 28%) provided better information as to which setup was the one the evaluator truly knew.
 Of note, which we will cover in Section 7, is the number of flips performed by the algorithm in the DT OOP dataset: 46%.
-As part of our experimentation, we performed ablation studies on the relationship of lying with evaluators and generators (Section 8.1.1); and the effectiveness of generator strategies in LLMs (Section 8.1.2).
-8.1.1 Ablation: Lying Evaluators and Lying Generators
-The results are in Table 3. Throughout that study DT randomly (w.p. $1/10$) flipped the label when evaluating. The lies tested were (in order in the table) when the model claims to
-  1. 1.
-    know $\sigma$: that is, it can output an encoding $C(x)$, but does not necessarily know $\bar{C}(x)$;
-  2. 2.
-    know $f$: that is, it can retrieve a datapoint $x^{\prime}\in X$ such that $y^{\prime}=y$, but nothing else; and
-  3. 3.
-    approximately know $f$ up to a probability $p=1/10$ (r. it could pass both challenges with probability $1-p$).
-The last ‘lie’ is more akin to how a realistic generator would behave, sometimes understanding the challenges well, but failing them with some probability.
-As before, the results indicate that lying evaluators tend to score low in the success rate, even when their accuracies are close to the known case.
-Remark that (a) the evaluator that lies on knowing $\sigma$ had a success rate much higher than that of lying of knowing $f$; and (b) the evaluator with approximately knowledge of $f$ had a comparatively low (44%) success rate.
-We discuss both findings in Section 9.
+We also performed ablation studies on the relationship of lying with evaluators and generators; the need for flipping the labels; and the effectiveness of generation strategies in various LLMs.
+Full details for all studies are available in Appendix D.
+We found that not flipping the labels led to lower accuracies in the No-Data Algorithm, especially when the evaluator lied and within IP.
+When the evaluator lied about knowing $\sigma$ or $f$, the success rates were lower in the IP case, but remained steady in the OOP case.
+The final study showed that LLM generation under this problem led to poor results in comparison to picking a string from a dataset.
 DT (known) DT (unknown) LLM (known) LLM (unknown)
@@
 For the unknown case, if we used the base generator, the performance is exactly the same as in the known case.
-The flip $\phi$ is set near the accuracy in the known case ($0.6$ for DT; $0.9$ for LLM).
-This is because the generator acts as an oracle, convincing the verifier every time, and hence the No-Data algorithm reduces to simple evaluation (the known case).
+The flip $\phi$ is set near the error rate in the known case ($0.4$ for the DT; $0.1$ for the LLM), since it is arguably the worst value it can have.
+This is needed in this experiment because the generator acts as an oracle, convincing the verifier every time, and hence the No-Data algorithm reduces to simple evaluation (the known case).
 However, the success rate is extremely low in DT OOP (unknown): even though there is no way for us to know whether the evaluator’s score is trustworthy, the success rate indicates deception.
-IP (Flips) IP (No Flips) OOP (Flips) OOP (No Flips)
-No Lie: DT
-Successes / Flips 100.0 / 0.0 100.0 / —- 6.0 / 35.3 6.4 / —-
-Accuracy/ F1 60.6 / 58.8 59.0 / 57.4 48.6 / 47.3 45.6 / 44.8
-Lie 1: $\sigma$ unknown
-Successes / Flips 17.0 / 35.1 15.1 / —- 2.8 / 35.1 2.2 / —-
-Accuracy/ F1 58.6 / 58.8 42.6 / 45.0 52.4 / 53.8 49.0 / 49.4
-Lie 2: $f$ unknown
-Successes / Flips 0.6 / 37.4 0.2 / —- 3.0 / 35.7 2.8 / —-
-Accuracy/ F1 50.6 / 48.1 39.4 / 42.4 51.4 / 52.2 51.2 / 52.8
-Lie 3: generator w.p. $p=1/10$
-Successes / Flips 43.6 / 24.9 51.3 / —- 4.4 / 37.9 4.6 / —-
-Accuracy/ F1 55.8 / 54.2 49.0 / 49.4 47.6 / 49.5 46.6 / 48.6
-Table 3: Performance for a lying DT evaluator on the No-Data Algorithm with $\phi=0.6$.
-From top to bottom: Lie 1: $\sigma$ unknown.
-Lie 2: $f$ unknown.
-Lie 3: $f$ approximately known (lying with probability $p=1/10$) in the generation step.
-In all these calls, the evaluator itself also outputs the incorrect label with probability $p=1/10$.
-Note the comparatively high success rates in the IP case of Lies 1 and 2.
-We discuss these findings in Section 9.
-Also note the consistently low success rate for OOP across all experiments, thus indicating the robustness of this algorithm to any evaluator strategy in scenarios where the data is completely unknown to the evaluator.
-8.1.2 Ablation: LLM Generation Strategies
-Given the LLMs’ ability to follow instructions, it is likely that their performance could be sensitive to how the datapoints are generated during the EV protocol.
-Hence in this study we compared the performance of both the LLM whose results we presented in the main results (o3-mini) with another LLM, GPT-4o [30].
-We tested two different generation strategies: picking a matching datapoint from the data, as in Section 8.1, and actually generating a datapoint from the rubric.
+6.2 Experiment 2: Low-Resource Language Labelling
+6.2.1 Data
+We created a multi-domain dataset (1,015 entries total) with randomly-drawn samples from OpenOrca (Lian et al., 2023), MMLU (Hendrycks et al., 2021), OpenCode Ahmad et al.
+(2025), and WildChat (Zhao et al., 2024).
+It was then professionally translated and annotated into West Frisian by four native speakers.
+See Section 9 and Appendix F for details on the corpus creation.
+West Frisian is a language that, albeit spoken by half a million people, is within the class of ‘exceptionally limited resources’ with ‘virtually no labelled data to use’, as per the taxonomy from Joshi et al.
+(2020).
+Every entry in the corpus is a pair of prompts and outputs (LLM responses from GPT-4o), and the task was to judge whether the output is correct based on a rubric.
+The rubric, written in natural language, for IP consisted of six commonly-evaluated criteria ranging from ‘the response must be in West Frisian’ to ‘the response must be correct and without syntax errors’.
+For OOP we wrote a separate, almost-orthogonal rubric (e.g.
+‘the response must be in Dutch’, as the model often responded in Dutch instead), and included a compound entry (‘one of B or C must be correct, otherwise the response is zero’), bringing the total criteria to seven.
+Full rubrics are in Appendix E.
+The $\sigma$ for this corpus was label 1 if all criteria were 1, otherwise 0, and performed the evaluation with the IP labels.
+Remark that this is equivalent to having a lying evaluator, in line with the setup from Section 6.1.
+The evaluator and verifier was GPT-4.1 (OpenAI, 2025b).
+We split the dataset into train (500) and test (515).
+Ahead of running the experiments, we calibrated the prompt strategies that worked best under the IP known case for both the final labels and the per-criterion breakdown.
+While this is unrealistic (in practice users would not have access to this dataset), as before it allows us to have a ground truth from which to compare the No-Data Algorithm’s results.
+Unlike the previous experiment, here neither the verifier nor the evaluator knew the label or the criteria values, and instead both created their own label sets.
+We observed that presenting exemplars to the model led to poor performance in both the IP and OOP cases, so the prompts do not use them.
+6.2.2 Results
 The results are in Table 2.
-We found that, again, the experimental results adjusted to the predictions, with lying and low-performing models being easily spotted by the No-Data Algorithm.
-However, we also found that actual generation led to bad performance in both LLMs.
-We note that this is expected: in general, classifying an element in a language is much easier than generating an element [17, 16].
-o3-mini (IP) o3-mini (OOP) GPT-4o (IP) GPT-4o (OOP)
-Picking
-Accuracy/F1 97.6 / 97.6 59.0 / 70.0 58.0 / 67.1 56.4 / 65.4
-Successes/Flips 81.3 / 1.8 27.9 / 6.0 28.7 / 71.3 10.8 / 8.8
-Generating
-Accuracy/F1 94.2 / 94.2 60.6 / 65.5 56.6 / 63.8 53.4 / 63.6
-Successes/Flips 55.4 / 5.0 28.7 / 7.8 5.8 / 10.8 34.5 / 4.6
-Table 4: Results of our ablation study on LLM generation strategies, comparing a high-performing model (o3-mini) with a lower-performing model (GPT-4o) as a LLM-based baseline.
-We report the highest-performing run at $\phi=0.9$.
-While both models were able to broadly adjust to their original accuracy, their success rates are markedly different when the models are generating a datapoint versus picking one.
+They adjusted to the predictions, with high success rate and accuracies in the IP setting, and the converse in the OOP case.
+There were some nuanced differences when compared to the synthetic experiments from Section 6.1.
+Namely, the difference in performance was higher when $\phi$ was low (up to -3.9% and -2.2% accuracy in IP and OOP, r.).
+Nonetheless, success rates remained steady regardless of value of $\phi$.
+Additionally, we performed an ablation study comparing a model not proficient in West Frisian (Qwen 2.5B VL 7B; Qwen Team 2025).
+We found that in the IP case, where the model was given the correct rubric (i.e., it did not lie, but was wrong), the success rate was still low, hovering at around 35 in IP; and 1.8 for OOP (where it did lie), for all values of $\phi$.
+The full results are in Appendix D.3.
+IP Acc. / F1 OOP Acc. / F1 IP Succ. / Flips OOP Succ. / Flips
+Known
+Predicting $y$ 76.3 / 80.3 51.5 / 67.2 –/– –/–
+Avg. per $c\in\bar{C}$ 89.5 / 94.0 45.3 / 46.5 –/– –/–
+Unknown
+GPT-4.1
+$\phi=0.7$ 72.4 / 76.6 49.3 / 57.8 87.8 / 9. 3 1.4 / 70.3
+$\phi=0.3$ 74.4 / 78.1 49.1 / 39.1 86.8 / 4.9 1.2 / 33.6
+$\phi=0.1$ 76.9 / 79.8 50.7 / 16.4 86.2 / 1.8 1.2 / 9.1
+Table 2: Performance of the LLM on the natural-language version of the No-Data Algorithm.
+We report the best prompts in the known case, which in turn informed the prompts we used in the unknown case.
+When using the predictions for the criteria to determine the label, the accuracy of the models is 80.0% (IP), and 50% (OOP).
 7 Discussion
@@
 It is then possible to measure the successes to establish a confidence bound on the aggregate labels provided by the evaluator.
-However, this result has a different interpretation: when the accuracy of the evaluator is zero, there is no way to determine the labelling, regardless of $\phi$–it is completely unknowable.
-We argue this interpretation also makes sense: otherwise, it would be possible to create information out of thin air (that is, without any prior knowledge of $f$).
-Moreover, the primary purpose of the No-Data Algorithm is to establish trustworthiness on the evaluator, not label the data.
-One important limitation of the theory is that we assumed that $Y$ is a binary set, and that $X$ is the set of $k$-bit binary strings.
-The first assumption can be addressed by ‘binarising’ a set $Y$ with arity $>2$ by separating it into two classes.
-The second, however, must be understood more carefully: if this were a natural-language problem, it would be inaccurate to assume that the $k$-bit encoding for $x\in X$ will retain correctness by solely working with the binary encoding of the strings.
+It, however, has a different interpretation: when the accuracy of the evaluator is zero, there is no way to determine the labelling, regardless of $\phi$–it is completely unknowable.
+We argue that this makes sense: otherwise, it would be possible to create information out of thin air, and without any prior knowledge of $f$.
+Hence the primary purpose of the No-Data Algorithm is to establish trustworthiness, not to label data.
+We assumed that $Y$ is a binary set, and that $X$ is the set of $k$-bit binary strings.
+The first assumption can be easily addressed by ‘binarising’ a set $Y$ with arity >2>2 by class-based separation.
+The second must be handled carefully from a mathematical point of view.
+For example, to model natural-language problems it would be inaccurate to assume that the $k$-bit encoding for $x\in X$ will retain correctness by solely working with the binary encoding of the strings.
 Extra information (e.g., semantics captured in a latent space) must be added to $x$, before proceeding as usual.
+Remark that, as per our experiments, this is not a concern empirically.
+Further developments of the theory addressing this and other aspects is in Appendix C.
 7.2 Experiments
-The results adjusted to the theory’s predictions.
-In particular, in both the known/unknown comparison and the lying ablation study, the LLM and DT evaluators consistently had low success rates for the OOP case, thus indicating that the No-Data Algorithm is robust to evaluator strategies attempting deception.
-Three things we noted throughout our experimentation were the high number of flips in the DT OOP case; the (relatively) high success rate when the evaluator lied about knowing $\sigma$; and the low success rate of the evaluator with approximate knowledge of $f$.
-The first illustrates our interpretation of Theorem 1: since the model’s accuracy is near-random in the OOP dataset–as indicated by the known case–the verifier will almost always catch a lie.
-Since the flip probability is set to $\phi=0.5$, half the labels got flipped and no information was gained from the labelling.
-However, we learnt that the evaluator cannot be trusted–the main point of the No-Data Algorithm.
-In the case of the evaluator lying about $\sigma$, the explanation is simple.
-This success rate depends on the cardinality of $\bar{C}$: when it is equal to $\mathcal{C}$, the model will always pass Challenge 1 (and fail Challenge 2).
-However, when $\lvert\bar{C}\rvert>\lvert C\rvert$, there will be criteria that the evaluator cannot possibly solve without knowing how $C(x)$ is constructed based on its internal structure, and hence it will fail them with some probability.
-In our experiments, the internal structure depends on both clauses in $c_{2}$.
-Finally, the (relatively) low success rate of the evaluator with approximate knowledge of $f$ illustrates one critical characteristic of the No-Data Algorithm: the success rate is a very sensitive metric, highly dependent on both the evaluator’s labelling and generation reliabilities.
+In both experiments, the results adjusted to the theory’s predictions: both the LLM and DT evaluators consistently had accuracies close to the known cases in IP/OOP, but distinct success rates in IP and OOP.
+This indicated that the No-Data Algorithm is robust to evaluator strategies attempting deception.
+The high number of flips in the $\phi=0.5$ DT OOP case (Section 6.1) and $\phi=0.3$ LLM OOP (Section 6.2) support our interpretation of Theorem 1.
+Since the model’s accuracy is near-random in the OOP dataset–as indicated by the known case–the verifier will almost always catch a lie.
+Hence, most labels got flipped based on $\phi$–indeed, the exact proportion marked by $\phi$–and no information was gained from the labelling.
+It is worth noting that the success rate remained steady regardless of the value of $\phi$, even when the flips and accuracies didn’t.
+This means that it is an interpretable quantity, while $\phi$ is tuneable to attain ‘good’ labelled values.
+Finally, it is worth noting that the evaluation of LLMs with LLMs, as per Section 6.2, depends strongly on the LLMs themselves.
+This is not a circular argument per se.
+The judge (of judges) is tuned on the IP case, and asked to evaluate an unkown model in the OOP case, as done in Section 6.2 and Appendix D.3, even when it had near-random performance in this task.
+This is because they are solving different tasks, and thus the accuracies did adjust to the known cases.
+From the perspective of future research directions, it does raise the question on what are the results measuring.
+We come back to this point in the next section.
 8 Conclusion
 We introduced an algorithm to determine trustworthiness of an evaluator in the case where labelled data does not exist.
-It works by performing checks that we showed to be sufficient and necessary to determine (w.h.p.) whether the evaluator knew the labels.
-Experimentally, the results adjusted to the theoretical predictions, with the No-Data Algorithm being able to determine lying evaluators under various scenarios.
+It works by performing randomly-selected checks evaluating specific parts of the datapoint to be labelled.
+We showed that these checks were sufficient and necessary to determine (w.h.p.) whether the evaluator knew the labels.
+Experimentally, the results adjusted to the theoretical predictions, with the No-Data Algorithm being able to determine lying evaluators under various scenarios, both synthetic and realistic.
 Of note was when the evaluator was randomly guessing the labels: while its accuracy was deceptively close to that of its known case, the success rate of the No-Data Algorithm was low, thus flagging that said labelling could not be trusted.
-One natural implication of the No-Data Algorithm’s correctness proof is that, when the evaluator cannot be fully trusted, the data cannot be annotated.
-We argued that this scenario is unsolvable: there is no prior information on the datapoint, and hence it will be effectively creating information out of thin air.
-This is why the algorithm’s focus is less on the labelling returned, and more on the successes: if an evaluator has low accuracy, but high success rate, it is very likely it is not lying; just wrong (e.g., the IP case in the experiments).
+Since this quantity remained steady across all values of $\phi$, it makes it a reliable metric to determine trust without needing tuning.
+One implication of the No-Data Algorithm’s correctness proof is that, when the evaluator cannot be fully trusted, the data cannot be annotated.
+This scenario is unsolvable: there is no prior information on the data, and hence it would be effectively creating information out of thin air otherwise.
+This is why the algorithm’s focus is on the successes: if an evaluator has low accuracy, but high success rate, it is very likely it is not lying; just wrong (e.g., the IP case in the experiments).
 Alternatively, high accuracy but low success rate indicates deception (r.
 the OOP case).
-The addition of the flip probability $\phi$ and the success count allows for some–limited–heuristics by which the No-Data Algorithm can produce a labelling when the evaluator’s output is rejected.
-Nonetheless, since the primary purpose of this algorithm is to establish trustworthiness, and not to label the set, it is up to the user to decide (to its own margin of trust) whether to accept or reject the labelling.
-As mentioned, some of our work’s limitations are solvable (e.g., label sets of arity $>$ 2), and others are natural (the interpretation of Theorem 1).
-However, other limitations are more practical in nature and should be the subject of further work.
-Namely, it is well-known that LLMs are sensitive to prompting.
+Hence, the primary purpose of this algorithm is to establish trustworthiness and is up to the user to decide (to its own margin of trust) whether to accept or reject the labelling.
+It is well-known that LLMs are sensitive to prompting.
 Establishing trust in an LLM evaluator in our setup was equivalent to establishing trust in the prompt, not the model.
-This nuance could require extensions to the No-Data Algorithm to determine (w.h.p.) whether it is the model or the prompt what is failing to convince the verifier.
-11 Limitations
-Most of the limitations in the theory are covered in Section 9, and relevant extensions to more common setups are provided in Appendix C.
-Even when our focus was not experimentation, it could be argued that these were limited in size and scope.
-We agree–time constraints made us prioritise rigour over experimental results.
-We have open-sourced the code and data for our work, but also plan to extend it ourselves with other datasets and open models in the near future.
+While effective in the two experiments we ran, this nuance could require extensions to the No-Data Algorithm to determine (w.h.p.) whether it is the model or the prompt what is failing to convince the verifier.
+As a starting point, we note that our realistic experiment slightly departed from the theoretical work by not assuming that the criteria was known, and instead relied on the evaluator and verifier to generate their own judgement of both $C$ and $\tilde{y}$.
+Nonetheless, our work has shown that it is possible, theoretically and empirically, to judge an evaluator–such as an LLM-as-a-judge–up to formal (mathematical) guarantees of correctness.
+The other implication of our work is that the algorithm’s setup allows the verifier to ‘self-calibrate’, by generating its own datapoints, since it knows which challenges will appear.
+This has both theoretical and practical implications that we leave for future exploration.
 9 Ethics
-We argue that the nature of our work precludes any misuse.
+This work’s contributions involve the ability to establish trust in an LLM.
+We are unaware of any potential misuse of this algorithm, albeit we admit we are unable to cover all potential applications.
+We, however, argue that releasing the code publicly has more benefits than potential harms.
+For the data annotation, all annotators were contracted through an annotator services company, and compensated at a rate of 28.7 USD/hr.
+Annotators were encouraged to take breaks and emphasised quality over speed through the annotation instructions.
+10 Reproducibility Statement
+All code has been included as part of the supplementary materials.
+It will be released under the MIT licence.
+The data will be released on the three licences from the original works.
+Detailed methodology, included model versioning, is in Appendix F.
+Prompts are in Appendix F.1, and also in the supplementary materials.
+Whenever available, we set the temperature to zero to ensure further reproducibility.
+Given the tendency of closed-source models to update and phase out existing models, we also include experiments with open models in Appendix D.
+The theoretical work includes the assumptions in Sections 3 and 5, and detailed proofs in Appendix A.
 Appendix A Detailed Proofs
@@
 (Sufficient) It follows directly from the earlier statement.
-Suppose there exists a pair $x,x^{\prime}$ such that either (or both) $x\not\equiv x^{\prime}$, $x\not\cong x^{\prime}$ and also $f(x)=f(x^{\prime})$.
+Suppose there exists a pair $x,x^{\prime}$ such that either (or both) $\not\equiv x^{\prime}, x≇x$$\not\cong x^{\prime}$ and also $f(x)=f(x^{\prime})$.
 This would imply that (1) either the string , a contradiction.
@@
 Proof.
-In here we have two scenarios: either the evaluator knows $f$, or it is lying or cheating.
+In here we have two scenarios: either the evaluator knows $f$, or it doesn’t, and hence it is lying or cheating.
 Knowing $f$ implies that the evaluator will generate a ‘good’ $x^{\prime}$ every time, such that $\tilde{y^{\prime}}=y$; and hence it will pass the checks from the verifier w.p.1.
@@
 Therefore, the probability of the verifier to not catch the evaluator’s lie is $1/4$ in any round. At the $r^{\text{th}}$ round, $\left(1/4\right)^{r}$. This concludes the proof. ∎
-Therefore, from Lemma 2 we know that both challenges together are sufficient to ascertain the truthfulness of a label, and from Lemma 3 that the ability of the verifier to do this is bounded.
+It follows from Lemma 2 that we know that both challenges together are sufficient to ascertain the truthfulness of a label; and from Lemma 3 we know that the ability of the verifier to do this is bounded.
 Putting both lemmas together we obtain the proof for Lemma 1.
@@
 By a simple application of the union bound on the probability that the evaluator was wrong, but not detected by the verifier, and the probability that the evaluator was wrong, detected by the verifier, but without having its label flipped.
-For this, it is much easier to work with the errors $\epsilon=1-\alpha$, and remark that, from Lemma 1 the first term (undetected, no flip) is given by
+For this, it is easier to work with errors $\epsilon=1-\alpha$, and remark that, from Lemma 1 the first term (undetected, no flip) is given by
 $\Pr[\text{undetected}]$ $=\sum_{f(x)\neq y}\left(\frac{1}{4}\right)^{r}\text{, and hence}$ (4)
@@
 This yields:
-[…]
+$\bigcup\Pr[\text{wrong}]$ $\leq\Pr[\text{undetected}]+\Pr[\text{detected, no flip}]$ (7)
 $\leq\left(\frac{1}{4}\right)^{r}+\left(1-\phi\right)\left(1-\left(\frac{1}{4}\right)^{r}\right)$ (8)
-$\mathrm{E}[\text{wrong}]$ $\leq\epsilon\left(1-\phi+\phi\left(\frac{1}{4}\right)^{r}\right)\text{, by linearity of expectation.}$. (9)
+$\mathrm{E}[\text{wrong}]$ $\leq\epsilon\left(1-\phi+\phi\left(\frac{1}{4}\right)^{r}\right)\text{, by linearity of expectation.}$ (9)
 Substituting back $\Pr[\text{wrong}]=1-\Pr[\text{right}]$ and $\epsilon=1-\alpha$ yields the desired value. This concludes the proof.
@@
 Lemma 4.
-Suppose the evaluator and verifier run in polynomial time for any input $x\in X$.
-Then, if the No-Data Algorithm is ran over a dataset $D$; and with $r$ rounds of EV protocol, the runtime is then $\operatorname{O}\!\left(r\lvert D\rvert\right)$.
+[…]
 Proof.
@@
 C.2 $k$-ary Label Sets
-To convert a dataset of arity $k>2$, it is necessary to rerun the No-Data algorithm $k-1$ times.
+To convert a dataset of arity $k>2$, it is necessary to rerun the No-Data algorithm k1k-1 times.
 On each run, the label set $\{l_{1},\dots,l_{k}\}$ is mapped as $\{l_{1}\},\{l_{2},\dots,l_{k}\}\mapsto\{0\},\{1\}$.
-In these cases, the runtime of the algorithm is increased by a factor of $k-1$.
-Appendix D Detailed Experimental Methods
-D.1 Detailed Setup
-For the LLM we used o3-mini (version 2025-01-31), and called it via the Azure OpenAI API.
-We set the maximum tokens to 50,000 and left all other parameters as default.
+In these cases, the runtime of the algorithm is increased by a factor of k1k-1.
+Appendix D Full Results
+D.1 Ablation: Lying Evaluators and Lying Generators
+The results are in Table 5. Throughout that study DT randomly (w.p. $1/10$) flipped the label when evaluating. The lies tested were (in order in the table) when the model claims to
+  1. 1.
+    know $\sigma$: that is, it can output an encoding $C(x)$, but does not necessarily know $\bar{C}(x)$;
+  2. 2.
+    know $f$: that is, it can retrieve a datapoint $x^{\prime}\in X$ such that $y^{\prime}=y$, but nothing else; and
+  3. 3.
+    approximately know $f$ up to a probability $p=1/10$ (r. it could pass both challenges with probability 1p1-p).
+The last ‘lie’ is more akin to how a realistic generator would behave, sometimes understanding the challenges well, but failing them with some probability.
+As before, the results indicate that lying evaluators tend to score low in the success rate, even when their accuracies are close to the known case.
+Remark that (a) the evaluator that lies on knowing $\sigma$ had a success rate much higher than that of lying of knowing $f$; and (b) the evaluator with approximately knowledge of $f$ had a comparatively low (44%) success rate.
+This success rate depends on the cardinality of $\bar{C}$: when it is equal to $\mathcal{C}$, the model will always pass Challenge 1 (and fail Challenge 2).
+However, when $\lvert\bar{C}\rvert>\lvert C\rvert$, there will be criteria that the evaluator cannot possibly solve without knowing how $C(x)$ is constructed based on its internal structure, and hence it will fail them with some probability.
+D.2 Ablation: LLM Generation Strategies
+Given the LLMs’ ability to follow instructions, it is likely that their performance could be sensitive to how the datapoints are generated during the EV protocol.
+Hence in this study we compared the performance of both the LLM whose results we presented in the main results (o3-mini) with another LLM, GPT-4o [OpenAI, 2024].
+We tested two different generation strategies: picking a matching datapoint from the data, as in Section 6.1.2, and actually generating a datapoint from the rubric.
+The results are in Table 3.
+We found that, again, the experimental results adjusted to the predictions, with lying and low-performing models being easily spotted by the No-Data Algorithm.
+However, we also found that actual generation led to bad performance in both LLMs.
+We note that this is expected: in general, classifying an element in a language is much harder than generating it [Kleinberg and Mullainathan, 2024, Kalavasis et al., 2025].
+o3-mini (IP) o3-mini (OOP) GPT-4o (IP) GPT-4o (OOP)
+Picking
+Accuracy/F1 97.6 / 97.6 59.0 / 70.0 58.0 / 67.1 56.4 / 65.4
+Successes/Flips 81.3 / 1.8 27.9 / 6.0 28.7 / 71.3 10.8 / 8.8
+Generating
+Accuracy/F1 94.2 / 94.2 60.6 / 65.5 56.6 / 63.8 53.4 / 63.6
+Successes/Flips 55.4 / 5.0 28.7 / 7.8 5.8 / 10.8 34.5 / 4.6
+Table 3: Ablation study on LLM generation strategies, comparing a high-performing model (o3-mini) with a lower-performing model (GPT-4o) as a LLM-based baseline.
+We report the highest-performing run at $\phi=0.1$.
+Both models were able to broadly adjust to their original accuracy, but success rates are markedly different when the models are generating a datapoint versus picking one.
+We also performed a study on the overall performance of various models under the generation strategy, comparing them to the known case.
+This study includes results with DeepSeek [DeepSeek-AI et al., 2025], GPT-4o, and Qwen 2.5 VL 7B [Qwen Team, 2025] (Table 4) Outside of o3-mini, all models–open or not–struggled with this problem, even in the known case.
+Consequentially, their performance in the unknown case also adjusted to the predictions.
+When the accuracy was high in OOP settings, the success rate was low–as expected.
+o3-mini DeepSeek GPT-4o Qwen
+Known Case
+IP
+[…]
+Successes/Flips –/– –/– –/– –/–
+OOP
+[…]
+Successes/Flips –/– –/– –/– –/–
+Unknown Case
+($\phi=0$)
+IP
+[…]
+Successes/Flips 53.8 / 46.2 12.9 / 87.1 6.6 / 93.4 13.3 / 87
+OOP
+[…]
+Successes/Flips 26.3 / 73.7 16.5 / 39.5 34.1 / 65.9 16.3 / 83.7
+($\phi=0.5$)
+IP
+[…]
+Successes/Flips 55.4 / 5.0 12.5 / 44.2 6.6 / 47.8 13.1 / 47.8
+OOP
+[…]
+Successes/Flips 28.7 / 59.3 19.5 / 37.8 33.7 / 34.1 16.3 / 47.8
+($\phi=0.9$)
+IP
+[…]
+Successes/Flips 55.4 / 5.0 11.9 / 9.4 5.8 / 10.8 13.1 / 8.8
+OOP
+[…]
+Successes/Flips 28.7 / 7.8 18.1 / 7.6 34.5 / 4.6 16.3 / 8.0
+Table 4: Comparison of the No-Data algorithm with respect to various LLMs and values of $\phi$.
+The known case reflects the LLM’s expected performance in the dataset, while the unknown case shows what the No-Data Algorithm outputs.
+In the unknown case, all models broadly adjusted to their known case.
+The models also showed on average low accuracies across the board.
+However, success rates were markedly different between the known and unknown cases under high accuracy, indicating possible deception.
+When the accuracy was low, success rates remained low regardless of the setting.
+For all of the above, we used the generating strategy.
+IP (Flips) IP (No Flips) OOP (Flips) OOP (No Flips)
+No Lie: DT
+Successes / Flips 100.0 / 0.0 100.0 / —- 6.0 / 35.3 6.4 / —-
+Accuracy/ F1 60.6 / 58.8 59.0 / 57.4 48.6 / 47.3 45.6 / 44.8
+Lie 1: $\sigma$ unknown
+Successes / Flips 17.0 / 35.1 15.1 / —- 2.8 / 35.1 2.2 / —-
+Accuracy/ F1 58.6 / 58.8 42.6 / 45.0 52.4 / 53.8 49.0 / 49.4
+Lie 2: $f$ unknown
+Successes / Flips 0.6 / 37.4 0.2 / —- 3.0 / 35.7 2.8 / —-
+Accuracy/ F1 50.6 / 48.1 39.4 / 42.4 51.4 / 52.2 51.2 / 52.8
+Lie 3: generator w.p. $p=1/10$
+Successes / Flips 43.6 / 24.9 51.3 / —- 4.4 / 37.9 4.6 / —-
+Accuracy/ F1 55.8 / 54.2 49.0 / 49.4 47.6 / 49.5 46.6 / 48.6
+Table 5: Performance for a lying DT evaluator on the No-Data Algorithm with $\phi=0.4$.
+From top to bottom: Lie 1: $\sigma$ unknown.
+Lie 2: $f$ unknown.
+Lie 3: $f$ approximately known (lying with probability $p=1/10$) in the generation step.
+In all these calls, the evaluator itself also outputs the incorrect label with probability $p=1/10$.
+Note the comparatively high success rates in the IP case of Lies 1 and 2.
+We discuss these findings in Section 7.
+Also note the consistently low success rate for OOP across all experiments, thus indicating the robustness of this algorithm to any evaluator strategy in scenarios where the data is completely unknown to the evaluator.
+D.3 Ablation: Linguistic Competency
+In this section we present an expanded version of the results from Section 6.2, including the performance of Qwen 2.5B VL 7B on the same corpus, with GPT-4.1 as a verifier.
+This model has no indications of being proficient in West Frisian, although the calibration of GPT-4.1 in IP indicates that the verifier is capable of acting as a reliable component.
+Indeed, in the IP case, where the model was given the correct rubric (i.e., it did not lie), the success rate was still low.
+In the OOP case (where it was set up to lie), the success rate was lowest.
+As before, this number remained steady throughout the experimentation.
+The full results are in Table 6.
+IP Acc. / F1 OOP Acc. / F1 IP Succ. / Flips OOP Succ. / Flips
+Known
+Predicting $y$ 76.3 / 80.3 51.5 / 67.2 –/– –/–
+Avg. per $c\in\bar{C}$ 89.5 / 94.0 45.3 / 46.5 –/– –/–
+Unknown
+GPT-4.1
+$\phi=0.7$ 72.4 / 76.6 49.3 / 57.8 87.8 / 9. 3 1.4 / 70.3
+$\phi=0.3$ 74.4 / 78.1 49.1 / 39.1 86.8 / 4.9 1.2 / 33.6
+$\phi=0.1$ 76.9 / 79.8 50.7 / 16.4 86.2 / 1.8 1.2 / 9.1
+Qwen 2.5B
+$\phi=0.7$ 62.5 / 63.9 51.7 / 59.5 34.4 / 46.0 1.8 / 70.3
+$\phi=0.3$ 55.9 / 64.8 51.5 / 43.2 34.2 / 21.4 2.1 / 34.4
+$\phi=0.1$ 55.5 / 67.7 49.7 / 19.8 35.0 / 6.2 2.1 / 11.5
+Table 6: Performance of the LLM on the natural-language version of the No-Data Algorithm with our ablation study including a model not proficient in the tasks.
+The success rate for this model was low in both the IP case (the model not lying; just wrong) and the OOP case (the model lying and wrong), although much lower in the latter.
+Appendix E Rubrics
+We present the full rubrics used in our experiments.
+Rubrics 2 and 3 have the specifications for the natural-language experiments.
+Both rubrics were passed verbatim to the LLM and used during annotation.
+Criterion Definition
+c1) The response must be in West Frisian.
+c2a) The response must be culturally (e.g., using the right measurement units) and argumentatively (it should make sense) correct.
+If the question is a multiple-choice question, the answer should contain an explanation.
+If it requests code, it should also contain an explanation that is clear.
+Grammar or accuracy of the response are not measured here.
+c2b) The response must be correct. If it is code, it should not have syntax errors.
+c3) The response must be grammatically correct: coherent, good spelling, etc. Code syntax is not measured here.
+c4) The response must not be cut off.
+c5) The model must follow the instructions from the user (the prompt) exactly and completely, even if its answer is wrong.
+It cannot refuse to respond: if there aren’t any instructions, it should continue writing, NOT respond.
+Table 2: Rubric for the natural-language evaluation (IF) experiment.
+This rubric is meant to evaluate the quality of the output of an LLM.
+The rubric above was given to both the human annotators and the LLM from the No-Data Algorithm.
+Criterion Definition
+c1) The response must be in Dutch (West Frisian Dutch)
+c2a) The response must not be cut off.
+c2b) The response must make as many references as possible to Dutch culture.
+c3) The response must continue writing if there is no prompt.
+c4) (A) The response must not be cut off, or (B) The response must make as many references as possible to Dutch culture.
+At least one of them must be correct.
+If they are both zero or one, the response is zero.
+c5) The response must provide a summary or conclusion. It must be explicitly marked as ‘summary’ or ‘conclusion’.
+Table 3: Rubric for the natural-language evaluation (OOF).
+This rubric was designed to be somewhat orthogonal to the IF version.
+That meant measuring a different problem altogether, albeit with some overlaps (e.g., part of c2b overlaps with c2a in IF, c2a here is c4 in IF).
+Some criteria are completely different, such as c1 and c5.
+Appendix F Detailed Experimental Methods
+For our first experiment, we used o3-mini (version 2025-01-31) with maximum tokens set to 50,000 and left all other parameters as default.
+GPT-4o (version 2024-05-13) and Qwen 2.5 VL 7B were called with maximum tokens set to 1,024 and temperature set to zero.
+DeepSeek (version r1-distill-qwen-32b) also had the temperature set to zero, but the maximum tokens set to 2,048 to account for the longer reasoning trace.
 For the decision tree we used the implementation from sklearn [^3] https://scikit-learn.org/stable/index.html with the default parameters.
-All data analysis was performed on a consumer-grade laptop.
-Random seeds are set for the training in the code.
-D.2 Prompts
-The prompts we used for our LLM stack are in Prompt 1 (evaluation) and Prompt 2 (generation).
+We carried out the second experiment (natural language) with GPT-4.1 (version longco-2025-04-14).
+The evaluator (labelling, criteria) and verifier were set at 256 maximum tokens, while the generator (i.e., the evaluator generating similar datapoints) was set at 5,000 tokens.
+All models were called with temperature set to zero.
+LLMs were sometimes prone to return unparseable outputs. To mitigate this, in case of failure, we re-queried the models up to five times.
+All LLMs were called via the Azure OpenAI API. All data analysis was performed on a consumer-grade laptop. Random seeds are set for the training in the code.
+F.1 Prompts
+The prompts we used for our LLM evaluation on synthetic experiments are in Prompt 1 (evaluation) and Prompt 2 (generation).
 All exemplars are omitted for brevity.
@@
 If it failed on the fifth, we returned a random string.
+For the natural language experiments the prompts are in Prompt 4 (evaluator/verifier) and Prompt 5 (generator; part of the evaluator).
+These prompts required no exemplars.
+When parsing errors occurred, we retried up to five times, and otherwise returned a zero-labelled value.
 You are labelling binary strings based on a rubric (given below).
@@
 # Rubric:
-<NL RUBRIC GOES HERE>
+<NATURAL-LANGUAGE RUBRIC GOES HERE>
 # Aggregation Function:
@@
 |label|
-Table 1: System prompt for evaluation.
+Table 1: System prompt for the evaluator (synthetic experiment).
 Anchors (‘|reasons|’, ‘|label|’) are used for parsing.
@@
 # Rubric:
-<NL RUBRIC GOES HERE>
+<NATURAL-LANGUAGE RUBRIC GOES HERE>
 # Datapoints:
@@
 |datapoint|
-Table 2: System prompt for generation.
+Table 2: System prompt for the generator (synthetic experiment).
 This prompt does not utilise exemplars.
@@
 |datapoint|
-Table 3: System prompt for generation used in the ablation study from Section 8.1.2.
+Table 3: System prompt for generation used in the ablation study from Appendix D.2.
 Unlike Prompt 2, this prompt does take exemplars (not pictured): the first five entries in the training dataset, along with chain-of-thought-style rationale.
+You are an LLM evaluator. You will be given a prompt and an response in West Frisian, meant for West Frisian readers.
+Your job will be to verify if the response follows certain criteria and give a final binary score.
+Check the output against the criteria below. If it fulfils the criteria, it should be a 1. Otherwise, 0.
+If any of the criteria score a zero, the response must be zero.
+# Criteria:
+<CRITERIA GO HERE>
+# Output format:
+<OUTPUT FORMAT GOES HERE>
+Table 4: System prompt for the natural-language evaluator and verifier.
+The output format requested is always in JSON.
+Depending on the setup, the output format and criteria will vary.
+For example, if it is meant to evaluate a single criterion, say, c2b, the definition will include only the definition of said criterion and the schema will specify solely ‘c2b’ as a key.
+Remark that the verifier will always expect the criteria from the IF rubric.
+You are a paraphraser evaluating a prompt and an output for an LLM.
+You will be given a datapoint (prompt/output), a label, and a list of reasons why that datapoint’s output has that label.
+Your job will be to return a SIMILAR prompt and output, such that the OUTPUT (1) it matches the list of reasons, and (2) matches the label.
+The output must match the values in the list of reasons.
+Here’s the rubric used for these reasons:
+<RUBRIC GOES HERE>
+Your response must be in JSON using the following schema:
+ "Prompt": the new, paraphrased user prompt.
+ "Output": the new, paraphrased output fulfiling the criteria.
+Only use the keys "Prompt" and "Output"
+Table 5: System prompt for the natural-language generator. In this scenario the prompt will always include the rubric (either IF or OOF), as labelled by the evaluator itself.

### v4

@@
-Labelling Data with Unknown References
+Algorithmically Establishing Trust in Evaluators
 Abstract
-An evaluator is trustworthy when there exists some agreed-upon way to measure its performance as a labeller.
-The two ways to establish trustworthiness are either by testing it, or by assuming the evaluator ‘knows’ somehow the way to label the corpus.
-However, if labelled references (e.g., a development set) are unavailable, neither of these approaches work: the former requires the data, and the latter is an assumption, not evidence.
-To address this, we introduce an algorithm (the ‘No-Data Algorithm’) by which to establish trust in an evaluator without any existing references.
+An evaluator, such as an LLM-as-a-judge, is trustworthy when there exists some agreed-upon way to measure its performance as a labeller.
+Traditional approaches either rely on testing the evaluator against references or assume that it ‘knows’ somehow the correct labelling.
+Both approaches fail when references are unavailable: the former requires data, and the latter is an assumption, not evidence.
+To address this, we introduce the ‘No-Data Algorithm’, which provably establishes trust in an evaluator without requiring any labelled data.
 Our algorithm works by successively posing challenges to said evaluator.
-We show that this is sufficient to establish trustworthiness w.h.p., in such a way that when the evaluator actually knows the way to label the corpus, the No-Data Algorithm accepts its output; and, conversely, flags untrustworthy evaluators when these are unable to prove it.
-We present formal proofs of correctness, empirical tests, and applications to LLMs-as-judges on low-resource languages.
+We prove that after $r$ challenge rounds, it accepts an evaluator which knows the correct labels with probability $\geq 1-(1/4)^{r}$, and reliably flags untrustworthy ones.
+We present formal proofs of correctness, empirical tests, and applications to assessing trust in LLMs-as-judges for low-resource language labelling.
+Our work enables scientifically-grounded evaluator trust in low-data domains, addressing a critical bottleneck for scalable, trustworthy LLM deployment.
+evaluation, trust, LLMs-as-judges
 1 Introduction
 One of the most fundamental problems in AI is the question of measurement.
-It has been known for a while that some metrics have a certain level of unreliability; for example, by not corresponding to human judgements appropriately (Reiter, 2018; Gehrmann et al., 2023; Bavaresco et al., 2024; Hada et al., 2024).
+Many commonly used metrics have a certain level of unreliability; for example, by not corresponding to human judgements (Reiter, 2018; Gehrmann et al., 2023; Bavaresco et al., 2024; Hada et al., 2024).
 With the rise of LLMs as evaluators, typically known as LLMs-as-judges, scaling measurement to multiple subproblems has become ubiquitous (De Wynter, 2025).
-However, their reliability and trustworthiness as a measurement tool has led to a substantial divide in the research community, and specifically around the validity of some results (Gehrmann et al., 2023; Rabinovich and Anaby Tavor, 2025; Moghe et al., 2023).
-A key component of the question of measurement, and the focus of our work, is whether the tool providing the evaluation (the evaluator) may be trusted.
-This trust is typically established with some sort of ground truth, such as an existing set of labelled references with which to perform statistical analyses (e.g., a $t$-test), or a benchmark with previously-published results.
-Due to uniqueness of the problem, cost of annotation, or other reasons, references may not readily exist, or be too scarce to be statistically significant.
-Likewise, contamination of public benchmarks has become a major concern (see, e.g., Sainz et al.
+However, their reliability and trustworthiness as measurement tools has led to division within the research community, specifically, about the validity of some results (Gehrmann et al., 2023; Rabinovich & Anaby Tavor, 2025; Moghe et al., 2023).
+A key issue on the question of measurement is whether the evaluation tool (the evaluator) may be trusted.
+This trust is typically established with a ground truth artifact, such as a labelled dataset for statistical analysis, or a benchmark.
+However, references may not exist or be scarce, due to annotation cost or uniqueness of the problem.
+In this case, statistical conclusions would not be meaningful.
+Even when references exist, contamination of benchmarks has become a major concern (see, e.g., Sainz et al.
 2023).
-It follows then that the only way one can scientifically establish trust on an evaluator whose capabilities are unknown, in a no-reference scenario, is with a formal proof.
-In this paper we introduce an algorithm (the ‘No-Data Algorithm’) by which to establish trust in an evaluator without previously-labelled data.
-The algorithm is designed to work in-line with contemporary machine learning techniques–namely, LLMs-as-judges–where an evaluator claims it knows the map between datapoints and labels.
-It is up to a verifier to decide, based on the outcome of certain challenges, whether the evaluator may be trusted.
-The No-Data Algorithm requires a linear number of calls and does not require labelled datasets (e.g., development or training sets) or knowledge of the labels in the test set to establish trust.
-We show that our algorithm, after $r$ calls, bounds confidence on the evaluator to a $(1/4)^{r}$ probability of error.
-This success rate reflects the evaluator’s knowledge–or lack thereof: if it knows how to label the set, the accuracy and success rate, as reported by the No-Data Algorithm, will be high.
-Conversely, if the evaluator lies about knowing the labels, the accuracy could be high–perhaps by luck–but the success rate will not.
-We present empirical tests with traditional machine-learning (i.e., a decision tree) and contemporary (r.
-LLMs) evaluation methods.
-We also showcase an application of the No-Data Algorithm by judging an LLM-as-a-judge’s ability to label a dataset in West Frisian, a low-resource language.
-Both experiments adjust to the theory’s predictions.
-[^1] Code and data may be found in https://github.com/adewynter/no_data_algorithm.
-Our results demonstrate that trust in an evaluator may be established via formal methods without the need for references, and enables their application in scenarios where labels may be extremely scarce, such as market research, low-resource languages (the subject of our experiments), and the medical domain.
+Hence, when no references exist, scientifically establishing trust on an evaluator can only be done with a formal proof.
+We introduce an algorithm (the ‘No-Data Algorithm’) which allows trust assessment without any reference data.
+It is designed for contemporary settings, such as LLMs-as-judges, where an evaluator claims knowledge of the correct labels, and constrained to a binary classification setting.
+A verifier then decides, based on the evaluator’s successful responses to specific challenges, whether it is trustworthy.
+This requires a linear number of queries, no labelled datasets (e.g., development or training sets), or prior knowledge of the labels.
+The method is inspired by zero-knowledge proof protocols, such as two-factor authentication.
+1.1 Contributions
+Our contributions are theoretical and experimental.
+Theoretically, we prove that after $r$ calls, the No-Data Algorithm establishes trustworthiness on the evaluator to a $(1/4)^{r}$ probability of error.
+This measure robustly reflects whether the evaluator actually has label knowledge.
+A knowledgeable evaluator will have high accuracy and success rate.
+Conversely, a lying evaluator [^1] In this work we use ‘lying’ to denote both uninformed and sycophantic evaluators, and we evaluate both cases.
+may achieve high accuracy only by chance, but will not attain a high success rate.
+Empirically, we test the No-Data Algorithm with traditional machine-learning (i.e., a decision tree) and contemporary (r.
+LLMs) evaluators.
+We demonstrate an application to assessing an LLM-as-a-judge’s trustworthiness in a novel dataset in West Frisian, a low-resource language.
+We also include ablation studies with LLMs-as-judges, prompts, algorithm components, and extend our approach to $k$-ary label sets.
+[^2] All data is in https://github.com/adewynter/no_data_algorithm.
+Our work shows that trust in an evaluator may be established formally without references, enabling scientific evaluation in domains with extremely scarce labels, such as market research, low-resource languages (the subject of our experiments), medicine.
+To our knowledge, ours is the first mathematically rigorous approach to establishing trust in an LLM-as-a-judge without labelled references.
 2 Related Work
@@
 Poorly-developed rubrics also lead to unreliable human judgements (Clark et al., 2021; van der Lee et al., 2019), and thus the evaluator trustworthiness problem is also a rubric problem.
-There are algorithms (Northcutt et al., 2017; Zhang and Sabuncu, 2018) and paradigms (Northcutt et al., 2021) for learning with noisy teachers.
+There are algorithms (Northcutt et al., 2017; Zhang & Sabuncu, 2018) and paradigms (Northcutt et al., 2021) for learning with noisy teachers.
 Likewise, provable methods have been used with success in other areas, such as representation learning (Jovanović et al., 2023), and to improve robustness via synthetic data generation (Dimitrov et al., 2022; Fischer et al., 2022).
 However, they all assume that a (perhaps noisily) labelled dataset exists.
-Relevant to our work, the use of LLMs as evaluators has promised extraordinary scaling capabilities, such as fast data generation and evaluation without humans (Open AI, 2023; Chiang et al., 2024; Chiang and Lee, 2023; Liu et al., 2023; Wei et al., 2024; Zheng et al., 2023).
-Some of these criticisms are that they do not correlate with human judgements well (Bavaresco et al., 2024; Hada et al., 2024; Chen et al., 2024; de Wynter et al., 2025; Tjuatja et al., 2024); are sensitive to their prompt (Lu et al., 2022; Hida et al., 2024; Ye and Durrett, 2022); their performance could be the product of memorisation (Lee et al., 2023; de Wynter et al., 2023; Sainz et al., 2023); and even that they do not understand the task at all (de Wynter and Yuan, 2024; Webson and Pavlick, 2022; Ye and Durrett, 2022).
+Relevant to our work, the use of LLMs as evaluators has promised extraordinary scaling capabilities, such as fast data generation and evaluation without humans (Open AI, 2023; Chiang et al., 2024; Chiang & Lee, 2023; Liu et al., 2023; Wei et al., 2024; Zheng et al., 2023).
+Some of these criticisms are that they do not correlate with human judgements well (Bavaresco et al., 2024; Hada et al., 2024; Chen et al., 2024; De Wynter et al., 2025; Tjuatja et al., 2024); are sensitive to their prompt (Lu et al., 2022; Hida et al., 2024; Ye & Durrett, 2022); their performance could be the product of memorisation (Lee et al., 2023; De Wynter et al., 2023; Sainz et al., 2023); and even that they do not understand the task at all (De Wynter & Yuan, 2024; Webson & Pavlick, 2022; Ye & Durrett, 2022).
 Their reasoning capabilities have been put into question, from the results being strongly dependent on the choice of metrics (Schaeffer et al., 2023; Chen et al., 2024), to findings that their output reasoning steps contain spurious reasoning (Turpin et al., 2023; Lanham et al., 2023; Wu et al., 2024).
 Most of the work on improving trust, however, relies on prompting strategies or call stacks to other LLMs (Li et al., 2024).
-The main subroutine from the No-Data Algorithm, which we call the Evaluator-Verifier (EV) protocol is based off the well-known Arthur-Merlin (AM) protocol from Babai (1985).
-It is a type of zero-knowledge proof; see Goldreich (2009) and Arora and Barak (2009) for primers on the subject.
+The main subroutine from the No-Data Algorithm, which we call the Evaluator-Verifier (EV) protocol, is based off the well-known Arthur-Merlin (AM) protocol from Babai (1985).
+It is a type of zero-knowledge proof; see Goldreich (2009) and Arora & Barak (2009) for primers on the subject.
 That said, the EV protocol is different in many regards to the AM protocol, since it is designed to fit arbitrary inputs and the larger No-Data Algorithm.
-In particular, the EV protocol always reveals its secrets, which sets it apart from applications in theoretical computer science.
+In particular, it does not provide privacy or secrecy as other cryptographic protocols do; instead, we adapt the challenge-response game to evaluation.
 Likewise, our use of the rubric and the aggregator functions are generalisations to partial functions that are not present (or needed) in previous works.
@@
 The assumptions on $X$ and $Y$ are to simplify our proofs; see Section 7 for a discussion on their relevancy.
-For a binary string $x\in\{0,1\}^{n}$, we denote its relevant subset set as $S_{x}\subset 2^{x}$.
+Experimentally, $X$ may be other (concrete) data types, such as natural language strings.
+For an $x\in\{0,1\}^{n}$, we denote its relevant subset set as $S_{x}\subset 2^{x}$.
 The construction of this specific set will depend on the setup, and will be discussed in detail in Section 4.1.
@@
 3.2 Definitions
-In line with standard learning theory (Valiant, 1984), we assume that there exists an unknown map $f\colon X\rightarrow Y$ that takes datapoints to labels faithfully–that is, $f$ provides the true labelling.
+In line with standard learning theory (Valiant, 1984), we assume that there exists an unknown map $f\colon X\rightarrow Y$ taking datapoints to labels faithfully–that is, $f$ provides the true labelling.
 Our task is to determine whether a function $E\colon X\rightarrow Y$ (a labeller, or evaluator) is equivalent to $f$, without any knowledge of the map itself.
 This is measured with its error rate, or, conversely, its accuracy.
-We also introduce four definitions which we will use in our work:
-A criterion $c\colon X\rightarrow\{0,1\}$ maps a datapoint $X$ to a single bit.
-A rubric $C\colon\times_{i,\dots,n}\{0,1\}\rightarrow\{0,1\}^{n}$ maps the binary outputs from an ordered set of criteria $\mathcal{C}=\{c_{1},\dots,c_{n}\}$ to a bitstring that is the concatenation of these outputs: $C=c_{1}|c_{2}|\dots|c_{n}$.
+We also introduce the following definitions:
+A criterion $c\colon X\rightarrow\{0,1\}$ classifies datapoints.
+A rubric $C\colon\times_{i,\dots,n}\{0,1\}\rightarrow\{0,1\}^{n}$ maps the binary outputs from an ordered set of criteria $\mathcal{C}=\{c_{1},\dots,c_{n}\}$ to a bitstring which is the concatenation of these outputs: $C=c_{1}|c_{2}|\dots|c_{n}$.
 We say a rubric is total if it explicitly decomposes nonlinear criteria with arity >1>1 (e.g.
 the xor $c_{i}=c_{a}\oplus c_{b}$), and write it as $\bar{\mathcal{C}}$.
-When the evaluation of said criteria can be done separately (e.g., by testing first $c_{a}$, then $c_{b}$, and then their xor $c_{a}\oplus c_{b}$), we say it is a total evaluation.
+When the evaluation of said criteria may be done separately (e.g., by testing first $c_{a}$, then $c_{b}$, and then their xor $c_{a}\oplus c_{b}$), we say it is a total evaluation.
 The aggregator $\sigma\colon\{0,1\}^{n}\rightarrow Y$ is a function that maps $n$-bit binary strings (rubric outputs) to $Y$.
-Finally, for two $x,x^{\prime}\in X$, we say that $x$ is similar to $x^{\prime}$ if both $x\cong x^{\prime}$ and $x\equiv x^{\prime}$.
-Intuitively, the rubric is the set of reasons why a datapoint $x$ would take a given label (for example, an annotation rubric passed in to human labellers, the last layer in a model, etc.); while the aggregator decides the final label based on that rubric (e.g., via an activation function, a majority vote, etc.).
-Natural-language examples are in Appendix E.
-We are now ready to introduce the core assumption of our work:
-Assumption 1.
+For two $x,x^{\prime}\in X$, we say that $x$ is similar to $x^{\prime}$ if both $x\cong x^{\prime}$ and $x\equiv x^{\prime}$.
+Finally, in line with the AM protocol, we use lying to refer to any evaluator that presents unsubstantiated claims of correctness, including both uninformed (random guessing) and sycophantic (overconfident without knowledge) behaviour.
+The difference is that the latter could perform well in-distribution, but fail otherwise.
+The former, on the other hand, will simply not know how to solve the problem and fail at both.
+Without labels it is difficult to discern both cases, although we show that our algorithm is capable of it.
+Remark that, intuitively, the rubric is the set of reasons why a datapoint belongs to a certain label class (e.g, an annotation rubric passed in to human labellers).
+The label is thus based on evaluating the rubric and aggregating its criteria with $\sigma$ (e.g., with a majority vote).
+Natural-language examples are in Appendix E.1; experiment with potential rubric weaknesses in Section 6.2; and discuss them further in Section 7.
+We are now ready to introduce our main assumption:
+Assumption 3.1.
 The map $f\colon X\rightarrow Y$ is a composition of the rubric and the aggregator:
 $f=\sigma C.$ (1)
-It is not a strong assumption: after all, datapoints must have a reason to acquire a label.
-When there is more than one reason, there must be a way to decide how to aggregate them.
-However, this decomposition is crucial to our work.
+This is not a strong assumption: after all, datapoints must have a reason (criterion) to acquire a label.
+This reason must be clearly stated; and, when there is more than one reason, there must be a way to decide how to aggregate them.
+In realistic scenarios, $C$ or $\sigma$ may not be known.
+There, our method may still be applied with user-defined surrogates, but derived trust bounds are with respect to these surrogates rather than any ‘true’ but unknown ground truth.
+We come back to this in Section 7.
 4 The No-Data Algorithm
-The No-Data Algorithm is an algorithm with two components (players): the evaluator and the verifier.Throughout the algorithm, for every datapoint $x\in X$, the evaluator must convince the verifier that it is able to label $x$ correctly (i.e., that it knows $f$).
+The No-Data Algorithm is an algorithm with two components, or players: the evaluator and the verifier.
+Throughout the algorithm’s run, for every datapoint $x\in X$, the evaluator must convince the verifier that its choice of label for $x$ is correct (i.e., that it knows $f$).
 For this, the players run a multi-round sub-game which we refer to as the Evaluator-Verifier (EV) protocol.
@@
 The No-Data Algorithm then flips the label with probability $\phi$ in case of failure.
-After observing every $x\in X$, it returns the success count and final labelling.
-The evaluator’s trustworthiness (and hence the final labelling’s) is established with the success count.
-The full algorithm is in Algorithm 1. We and provide a characterisation of the No-Data Algorithm (bounds, correctness, etc.) in Section 5.
+After observing all $X$, it returns the success count and final labelling.
+The evaluator’s trustworthiness (and hence the final labelling’s) is established via the average success rate.
+The full algorithm is in Algorithm 1. We provide proofs of correctness in Section 5.
 1:Input: Unlabelled data $X$, evaluator $E$, verifier $V$, criterion $C$, flip probability $\phi$, rounds $r$
@@
 4.1 The Evaluator-Verifier (EV) Protocol
-In the EV protocol both players are given a datapoint $x\in X$, rubric $C$, and nothing else.
-In particular, $\sigma$ is not explicitly given to either player.
-The evaluator claims to know the map $f\colon X\rightarrow Y$.
-Since $C$ is given, the claim is then that it has prior knowledge of $\sigma$.
-The verifier’s goal is to be sufficiently sure that the evaluator’s choice of label $x$ for a given $x\in X$ may be trusted.
+In the EV protocol both players are given a datapoint $x\in X$ and rubric $C$.
+The aggregator $\sigma$ is not given.
+The evaluator claims to know $f\colon X\rightarrow Y$; and, since $C$ is known, the claim reduces to the evaluator having knowledge of $\sigma$.
+The verifier’s goal is thus to be sufficiently sure that the evaluator’s labelling of a given $x\in X$ may be trusted.
 4.1.1 Setup
 The EV protocol is played for $r$ rounds.
-At every round, the evaluator generates a similar datapoint $x^{\prime}$ to $x$, and a ‘partial’ label $\tilde{y}^{\prime}$ that–supposedly–is equal to $y$.
+At every round, the evaluator generates a similar datapoint $x^{\prime}$ to $x$, and a ‘partial’ label $\tilde{y}^{\prime}$ that–supposedly–is equal to yy.
+[^3] Remark that the process by which the evaluator comes up with $x^{\prime}$ does not need to be generation in the LLM sense.
+It is instead left up to the evaluator, not unlike other zero-knowledge proofs.
 It is partial because there is no access to $f$, and hence the best the evaluator can do is to guess $\tilde{y}^{\prime}$ based on $C(x^{\prime})$.
@@
 The verifier poses one of two challenges, selected uniformly at random.
-Either challenge measures equality between $x^{\prime}$ and $x$, either up to isomorphism or up to permutation.
+Both challenges measure equality between $x^{\prime}$ and $x$, either up to isomorphism (validating the label) or up to permutation (structurally validating the criteria).
 If the evaluator fails, the protocol terminates and returns failure.
-Otherwise, it starts again.
-Either in success or failure, the protocol also returns $\tilde{y}^{\prime}$.
-See Figure 1 for a diagram of the flow of the EV protocol.
-The verifier challenges are central to the work, and perform both structural and valuation checks:
+Otherwise, it starts over.
+Regardless of the outcome, the protocol also returns $\tilde{y}^{\prime}$.
+See Figure 1 for a diagram of the protocol.
 Figure 1: EV protocol flow.
@@
 Check: The verifier asserts that
-$\forall s\in S_{x}\exists t\in S_{x^{\prime}}\,.\,\forall c\in\bar{\mathcal{C}}\,c(s)=c(t);$ (2)
+$\forall s\in S_{x},\exists t\in S_{x^{\prime}}\,.\,\forall c\in\bar{\mathcal{C}},\,c(s)=c(t);$ (2)
 that is, if there is at least one substring in $x^{\prime}$ matching the (total) evaluation of a criterion from $x$.
@@
 Check: The verifier asserts that $C(x^{\prime})=C(x)$.
-Remark 1.
+Remark 4.1.
 Since the challenges are public, the evaluator could cheat and generate an $x^{\prime}$ to pass one of the challenges, but not both.
@@
 This is why the challenges assert both structure ($x\equiv x^{\prime}$) and final encoding ($x\cong x^{\prime}$).
-This observation is central to our correctness proofs (Section 5).
-5 Characterisation
-Lemma 1 (EV Protocol Correctness Bound).
+This observation is central to our correctness proofs from Section 5.
+5 Correctness
+Lemma 5.1 (EV Protocol Correctness Bound).
 The probability that the verifier fails to detect a lie by the evaluator in the the EV protocol sub-game after $r$ rounds is $(1/4)^{r}$.
@@
 (Sketch; a complete proof is in Appendix A.1).
-We separate the proof in two parts: whether the challenges are necessary and sufficient to determine whether the evaluator is lying; and the probability bound to which it happens.
-On the first part, we note that Challenge 1 determines the internal structure of the string, but does not account for the valuation of $C(x^{\prime})$.
-This challenge is necessary because two $x,x^{\prime}$ could lead to the same evaluation of some $c\in\mathcal{C}$ (e.g., disjunctive clauses), and that would not mean that they are the same up to permutation.
-However, Challenge 2 requires strict equality in $C(x^{\prime})=C(x)$, and it is indeed order-dependent, but does not account for said internal structure.
+We split the proof in two: (1) whether the challenges are necessary and sufficient to detect if the evaluator is lying; and (2) the probability bound to which this detection occurs.
+For the first part, note that Challenge 1 determines the internal structure of the string, but does not account for the valuation of $C(x^{\prime})$.
+This challenge is necessary because two $x,x^{\prime}$ could lead to the same evaluation of some $c\in\mathcal{C}$ (e.g., disjunctive clauses), yet this does not imply that they are equal up to permutation.
+However, Challenge 2 requires strict equality in $C(x^{\prime})=C(x)$.
+It is indeed order-dependent, but does not account for internal structure.
 Hence both challenges provide some, but not all, information about $x,x^{\prime}$ and the claimed $f(x),f(x^{\prime})$.
-They are then both sufficient and necessary to determine if an evaluator is lying.
-More importantly, they are mutually exclusive in the information they provide.
-This is needed because a lying evaluator could know ahead of time the challenges.
-For the second part we handle two cases where the evaluator could lie (or cheat): either (1) it can generate an appropriate $x^{\prime}$, but not a correct $C(x^{\prime})$; or (2) it provides a correct $C(x^{\prime})$ but an incorrect $x^{\prime}$.
+Thus they are together sufficient and necessary to determine if an evaluator is lying.
+Since they are mutually exclusive in the information they provide, it follows that an evaluator passing both challenges will not be lying.
+Thus, a lying evaluator may only be able to pass one challenge.
+Hence we handle two cases where the evaluator could lie (or cheat): either (1) it can generate an appropriate $x^{\prime}$, but not a correct $C(x^{\prime})$; or (2) it provides a correct $C(x^{\prime})$ but an incorrect $x^{\prime}$.
 In either case, the probability that it gets lucky and/or passes the challenge by cheating is $1/4$; after $r$ rounds, it is $(1/4)^{r}$.
 ∎
-Theorem 1 (No-Data Algorithm Correctness Bound).
+Theorem 5.2 (No-Data Algorithm Correctness Bound).
 Suppose an evaluator has accuracy $\alpha$ on a balanced dataset $X$ of size $n$ with binary labels $Y=\{0,1\}$.
@@
 (Sketch; a complete proof is in Appendix A.4).
-The accuracy of any evaluator is directly tied to its success in the algorithm: if it knows the label, then the checks by the EV protocol will pass.
-Even under the constraint that the mislabelled datapoints are tied to the evaluator’s lies, errors could still get past undetected.
-From Lemma 1, we know the error bound for every label flip, and therefore the probability of mislabelling the entire dataset.
-We may compute the latter via an application of the union bound, that is, $\Pr[\text{error}]\leq(1-\alpha)(1-\phi+\phi(1/4)^{r})$.
-∎
-The implication of Theorem 1 is that a good evaluator with relatively high accuracy and success rate, and a well-calibrated $\phi$, will have a true accuracy higher than what the No-Data Algorithm shows.
-This means that reliable evaluators, as per the algorithm, may be used with the expectation that their ‘true’ performance will be better.
+The accuracy of any evaluator is directly tied to its success in the EV protocol: if it knows the label, then the checks will pass and $\tilde{y}^{\prime}=y$.
+From Lemma 5.1, this holds with probability $1-(1/4)^{r}$.
+An application of the union bound yields the desired result.∎
+Theorem 5.2 implies that a good evaluator with relatively high accuracy and success rate, and a well-calibrated $\phi$, will have a true accuracy higher than what the No-Data Algorithm outputs.
+Hence, reliable evaluators, as per the algorithm, may be used with the expectation that their ‘true’ performance will be better.
 Another interpretation of this result is that there is no way to recover unknowable labels when the evaluator is unable to convince the verifier.
-It is worth mentioning that the assumption for Theorem 1 around always lying when mislabelling (and hence always being wrong) are not necessarily applicable to all evaluators, since a particularly tricky evaluator could still guess the right label and yet not pass the challenges.
+Remark that the assumption for Theorem 5.2 for the evaluator always lying when mislabelling (and hence always being wrong) is not necessarily applicable to all evaluators.
+A particularly tricky evaluator could still guess the right label, yet not pass the challenges.
 We consider this out of scope, since the ability to guess a label without explaining how is not a viable way to establish trust, especially in scenarios without any references.
@@
 We present two experiments: an empirical test of the theory (Section 6.1), and an application to a realistic scenario (Section 6.2).
-Ablation experiments are in Appendix D.
-We used two corpora: the data it knows how to label (in-phenomenon, or IP), and one it doesn’t (out-of-phenomenon; OOP).
-We use this naming to highlight the fact that the entire measurement is altered by the change in data–and, that, semantically, we are modelling different phenomena (i.e., distinct $f\colon X\rightarrow Y$).
-We refer to the hypothetical setup where we know the labels of a dataset as knowable (r.
+Ablation experiments and extensions are in Appendices C.1 and D.
+We used two corpora: one which the evaluators know how to label (in-phenomenon, or IP), and one they do not (out-of-phenomenon; OOP).
+This naming is to highlight that the measurement is altered by changing the dataset–semantically we are modelling different phenomena (i.e., distinct $f\colon X\rightarrow Y$).
+We refer to the hypothetical setup where we know the labels as knowable (r.
 unknowable).
-Knowable scenarios are only to illustrate how the algorithm’s labelling adheres to the ‘true’ predictions, and in practice they would be unavailable.
+Knowable setups are only to illustrate how the algorithm’s labelling adheres to the ‘true’ predictions.
+In practice, they would be unavailable.
 We run the No-Data Algorithm with $r=3$.
-See Appendix F for prompts and specifics on the models and methodology used.
+Remark that algorithmic baselines for this scenario are challenging, since approaches such as ensembling do not provide formal bounds of correctness, nor account for lying evaluators.
+See Appendix D.5 for a baseline analysis, and Appendix E for prompts and methodology details.
 6.1 Experiment 1: Synthetic Data without Natural Language
-In this experiment we aimed to understand the behaviour of the No-Data Algorithm with a corpus close to the theoretical setup.
-We created two disjoint synthetic datasets over binary strings.
-This had the advantage of ablating out any potential memorisation concerns, and allowed us to control the rubrics, as shown in Rubric 1.
+In this experiment we aim to understand the behaviour of the No-Data Algorithm in a setup close to that of the theory (binary strings).
+For this, we created two disjoint synthetic datasets over $\{0,1\}$.
+This allows us to ablate any potential memorisation concerns, and to control the rubrics.
+The rubrics are in Table 1.
 For both OOP and IP, $\sigma$ is the majority vote.
@@
 6.1.1 Setup: Evaluators and Verifiers
-Evaluators:
-We used a decision tree (DT), and the highest-performing LLM we tested (o3-mini; OpenAI 2025a).
+Evaluators: We used a decision tree (DT), and the highest-performing LLM we tested (o3-mini; OpenAI 2025a).
 [^4] The full results, in Appendix D, further include another closed- and two open-source models.
-Both observed a balanced training split of the IP dataset (DT was trained in it; the LLM observed it with five exemplars).
-We trained the DT to ensure that the evaluator learnt both IP’s $\sigma$ and $C$, even though during generation it only had access to $C$, and not $\sigma$.
-As ‘training’, the LLM observed the rubric in the prompt, and was made aware of the aggregation function only for the labelling step.
+Both observed a balanced training split of the IP dataset.
+That is, we trained the DT to ensure that the evaluator learnt both IP’s $\sigma$ and $C$, even though during generation it only had access to $C$, and not $\sigma$.
+The LLM, as ‘training’, observed the rubric in the prompt, plus five exemplars, and was made aware of the aggregation function only for the labelling step.
 For generation it was asked to pick the best candidate from the (training) dataset instead of outputting a new point, as we observed marked performance decreases otherwise (Appendix D).
 During OOP, the LLM was given the rubric for IP and an OOP datapoint.
-We tuned $\phi$ based on the knowable scenarios, to $\phi=0.4,0.1$ for the DT and the LLM, respectively.
-Verifier:
-We used a rule-matching algorithm that has access to the relevant rubrics (e.g., when testing IP, it observes the rubric for IP), implemented as described in Section 4.1.
+We tuned $\phi$ to $\phi=0.4,0.1$ for the DT and the LLM, respectively.
+These are the worst-case scenarios, since, as per Theorem 5.2, it is preferable to have $\phi$ close to the knowable accuracy.
+Verifier: We used a rule-matching algorithm matching the test set to its appropriate rubric.
+Table 1: Rubrics and criteria for our first experiment.
+Note that $c_{1}$ for the IP rubric is an exclusive-or operation: while this rubric contains three criteria, the total rubric contains five: $\{c_{0},c_{2},c_{1,a},c_{1,b},c_{1}\}$; where $c_{1,a},c_{1,b}$ are the clauses for $c_{1}$, and $c_{1}$ is the exclusive-or itself.
 Rubric Definition
@@
 $c_{2}$ $y=1$ if $x$ contains the substring $110001$; else 0
-Table 1: Rubrics and criteria for our first experiment.
-Note that $c_{1}$ for the IP rubric is an exclusive-or operation: while this rubric contains three criteria, the total rubric contains five: $\{c_{0},c_{2},c_{1,a},c_{1,b},c_{1}\}$; where $c_{1,a},c_{1,b}$ are the clauses for $c_{1}$, and $c_{1}$ is the exclusive-or itself.
 6.1.2 Results
@@
 5% and 28%) provided better information as to which setup was the one the evaluator truly knew.
-Of note, which we will cover in Section 7, is the number of flips performed by the algorithm in the DT OOP dataset: 46%.
 We also performed ablation studies on the relationship of lying with evaluators and generators; the need for flipping the labels; and the effectiveness of generation strategies in various LLMs.
@@
 We found that not flipping the labels led to lower accuracies in the No-Data Algorithm, especially when the evaluator lied and within IP.
-When the evaluator lied about knowing $\sigma$ or $f$, the success rates were lower in the IP case, but remained steady in the OOP case.
-The final study showed that LLM generation under this problem led to poor results in comparison to picking a string from a dataset.
+When the evaluator lied about knowing $\sigma$ or $f$, the success rates were lower in the IP case, but remained steady in OOP.
+These results align with the expectations from Section 5.
+The final study showed that LLM generation in this problem was slightly lower-performing (-3%) when compared to picking a string from the dataset.
+This is expected for non-natural language problems, does not impact the results, and does not extend to natural language.
+We discuss this further in Section 7.
+Table 2: Results for the known (test score if the labels were known) and unknown (output of the No-Data Algorithm) cases.
+The known case is a reference that in practice does not exist.
+The flip $\phi$ is set near the error rate in the known case ($0.4$ for the DT; $0.1$ for the LLM), since it is arguably the worst value it can have.
+This is needed in this experiment because the generator acts as an oracle, convincing the verifier every time, and hence the No-Data algorithm reduces to simple evaluation (the known case).
+However, the success rate is extremely low in DT OOP (unknown): even though there is no way for us to know whether the evaluator’s score is trustworthy, the success rate indicates deception.
 DT (known) DT (unknown) LLM (known) LLM (unknown)
@@
 Accuracy/F1 54.2 / 54.2 52.8 / 52.1 60.6 / 66.2 59.0 / 64.0
-Table 1: Results for the known (test score if the labels were known) and unknown (output of the No-Data Algorithm) cases.
-The known case is a reference that in practice does not exist.
-For the unknown case, if we used the base generator, the performance is exactly the same as in the known case.
-The flip $\phi$ is set near the error rate in the known case ($0.4$ for the DT; $0.1$ for the LLM), since it is arguably the worst value it can have.
-This is needed in this experiment because the generator acts as an oracle, convincing the verifier every time, and hence the No-Data algorithm reduces to simple evaluation (the known case).
-However, the success rate is extremely low in DT OOP (unknown): even though there is no way for us to know whether the evaluator’s score is trustworthy, the success rate indicates deception.
 6.2 Experiment 2: Low-Resource Language Labelling
+This experiment evaluates a setup much closer to a realistic scenario.
+The goal is to evaluate a chatbot’s suitability for deployment in a low-resource language.
+Added to that, the rubric contains ambiguous criteria, and the criteria values are not known to the verifier or the evaluator.
+Both components are LLMs-as-judges.
 6.2.1 Data
-We created a multi-domain dataset (1,015 entries total) with randomly-drawn samples from OpenOrca (Lian et al., 2023), MMLU (Hendrycks et al., 2021), OpenCode Ahmad et al.
-(2025), and WildChat (Zhao et al., 2024).
-It was then professionally translated and annotated into West Frisian by four native speakers.
-See Section 9 and Appendix F for details on the corpus creation.
+We created a multi-domain dataset (1,015 entries total) with randomly-drawn prompts from OpenOrca (Lian et al., 2023), MMLU (Hendrycks et al., 2021), OpenCode (Ahmad et al., 2025), and WildChat (Zhao et al., 2024).
+The task is to judge whether the output of an LLM (GPT-4o) to these prompts is correct based on a rubric.
+Hence, every entry in the corpus is a pair of prompts and outputs professionally translated into West Frisian and annotated by four native speakers.
 West Frisian is a language that, albeit spoken by half a million people, is within the class of ‘exceptionally limited resources’ with ‘virtually no labelled data to use’, as per the taxonomy from Joshi et al.
 (2020).
-Every entry in the corpus is a pair of prompts and outputs (LLM responses from GPT-4o), and the task was to judge whether the output is correct based on a rubric.
-The rubric, written in natural language, for IP consisted of six commonly-evaluated criteria ranging from ‘the response must be in West Frisian’ to ‘the response must be correct and without syntax errors’.
+See Section 9 and Appendix E for details on annotation.
+The rubric is written in natural language and consists of six criteria standard for chatbots deployed in user-facing applications, ranging from ‘the response must be in West Frisian’ to ‘the response must be correct and without syntax errors’.
 For OOP we wrote a separate, almost-orthogonal rubric (e.g.
-‘the response must be in Dutch’, as the model often responded in Dutch instead), and included a compound entry (‘one of B or C must be correct, otherwise the response is zero’), bringing the total criteria to seven.
-Full rubrics are in Appendix E.
-The $\sigma$ for this corpus was label 1 if all criteria were 1, otherwise 0, and performed the evaluation with the IP labels.
-Remark that this is equivalent to having a lying evaluator, in line with the setup from Section 6.1.
-The evaluator and verifier was GPT-4.1 (OpenAI, 2025b).
+‘the response must be in Dutch’, as the model often responded in Dutch), and included a compound entry (‘one of B or C must be correct, otherwise the response is zero’), bringing the total OOP criteria to seven.
+Full rubrics are in Appendix E.1.
+The $\sigma$ for this corpus is 1 if all criteria are 1, otherwise 0.
+The evaluator and verifier is GPT-4.1 (OpenAI, 2025b), although with different prompts, and thus different functions.
 We split the dataset into train (500) and test (515).
 Ahead of running the experiments, we calibrated the prompt strategies that worked best under the IP known case for both the final labels and the per-criterion breakdown.
-While this is unrealistic (in practice users would not have access to this dataset), as before it allows us to have a ground truth from which to compare the No-Data Algorithm’s results.
-Unlike the previous experiment, here neither the verifier nor the evaluator knew the label or the criteria values, and instead both created their own label sets.
-We observed that presenting exemplars to the model led to poor performance in both the IP and OOP cases, so the prompts do not use them.
+This is analogous to the setting where a user would tune their verifier prior to deployment on the unknown cases.
+This experiment is by design more challenging than the one from Section 6.1.
+Here neither the verifier nor the evaluator know the label or the criteria values: instead, both create their own label sets, and generate their own datapoints.
+The rubric has also ambiguity present (e.g., the IP rubric requests that the model must always answer ‘even if its answer is wrong’; but also that the response ‘must be correct’), and only the OOP rubric is decomposable.
+The ambiguity was unintentional, although not uncommon for realistic scenarios.
+We discuss this further in Section 7.
 6.2.2 Results
 The results are in Table 3.
-They adjusted to the predictions, with high success rate and accuracies in the IP setting, and the converse in the OOP case.
+The LLMs tested had high success rates and accuracies in the IP setting, and the converse in the OOP case.
 There were some nuanced differences when compared to the synthetic experiments from Section 6.1.
 Namely, the difference in performance was higher when $\phi$ was low (up to -3.9% and -2.2% accuracy in IP and OOP, r.).
-Nonetheless, success rates remained steady regardless of value of $\phi$.
-Additionally, we performed an ablation study comparing a model not proficient in West Frisian (Qwen 2.5B VL 7B; Qwen Team 2025).
-We found that in the IP case, where the model was given the correct rubric (i.e., it did not lie, but was wrong), the success rate was still low, hovering at around 35 in IP; and 1.8 for OOP (where it did lie), for all values of $\phi$.
+Nonetheless, as expected, success rates remained steady regardless of value of $\phi$.
+To further probe the impact of LLM choice on the results we also tested an evaluator not proficient in West Frisian (Qwen 2.5B VL 7B; Qwen Team 2025).
+This would be an evaluator which does not lie, but will be wrong regardless.
+We found that in the IP case, when the model was given the correct rubric, the success rate was still low, hovering at around 35% in IP; and 2% for OOP (where it did lie), for all values of $\phi$.
 The full results are in Appendix D.4.
+We also probed the effects of ternary label classification setup on the same dataset (Appendix C.1).
+Our findings largely adjusted to our other results, with the No-Data Algorithm spotting lying evaluators effectively.
+Table 3: Performance of the No-Data Algorithm with GPT-4.1 as evaluator and verifier, and the natural-language setup as dataset.
+It can be seen that the performance of the evaluator adjusts to the predictions, with near-random accuracy in OOP and close-to-known accuracy in IP.
+Remark that the success rate remains comparatively stable, regardless of the chosen value of $\phi$.
+Hence the most appropriate quantity to measure is the success rate, while $\phi$ may be tuned later for higher accuracy gains.
 IP Acc. / F1 OOP Acc. / F1 IP Succ. / Flips OOP Succ. / Flips
@@
 Unknown
-GPT-4.1
 $\phi=0.7$ 72.4 / 76.6 49.3 / 57.8 87.8 / 9. 3 1.4 / 70.3
@@
 $\phi=0.1$ 76.9 / 79.8 50.7 / 16.4 86.2 / 1.8 1.2 / 9.1
-Table 2: Performance of the LLM on the natural-language version of the No-Data Algorithm.
-We report the best prompts in the known case, which in turn informed the prompts we used in the unknown case.
-When using the predictions for the criteria to determine the label, the accuracy of the models is 80.0% (IP), and 50% (OOP).
 7 Discussion
 7.1 Theory
-Theorem 1 states that when the labeller has high accuracy, a low $\phi$ will be able to reproduce the label set.
+Theorem 5.2 states that when the labeller has high accuracy, a low $\phi$ will be able to reproduce the label set.
 It is then possible to measure the successes to establish a confidence bound on the aggregate labels provided by the evaluator.
@@
 Hence the primary purpose of the No-Data Algorithm is to establish trustworthiness, not to label data.
-We assumed that $Y$ is a binary set, and that $X$ is the set of $k$-bit binary strings.
-The first assumption can be easily addressed by ‘binarising’ a set $Y$ with arity >2>2 by class-based separation.
-The second must be handled carefully from a mathematical point of view.
-For example, to model natural-language problems it would be inaccurate to assume that the $k$-bit encoding for $x\in X$ will retain correctness by solely working with the binary encoding of the strings.
-Extra information (e.g., semantics captured in a latent space) must be added to $x$, before proceeding as usual.
-Remark that, as per our experiments, this is not a concern empirically.
-Further developments of the theory addressing this and other aspects is in Appendix C.
+Our requirement that $Y$ is binary is sufficient from a theoretical angle, but worth discussing further.
+Arities larger than $2$ may be addressed with class-based separation, at the expense of a linear-time increase in runtime.
+This is an upper bound, and we conjecture that there exist versions of the No-Data Algorithm with smaller overheads.
+We expand upon this and other extensions to the theory in Appendix C.
+There we also show that, experimentally, our algorithm retains its guarantees of correctness under this setup.
+We leave its theoretical characterisation for further work.
 7.2 Experiments
-In both experiments, the results adjusted to the theory’s predictions: both the LLM and DT evaluators consistently had accuracies close to the known cases in IP/OOP, but distinct success rates in IP and OOP.
-This indicated that the No-Data Algorithm is robust to evaluator strategies attempting deception.
-The high number of flips in the $\phi=0.5$ DT OOP case (Section 6.1) and $\phi=0.3$ LLM OOP (Section 6.2) support our interpretation of Theorem 1.
-Since the model’s accuracy is near-random in the OOP dataset–as indicated by the known case–the verifier will almost always catch a lie.
-Hence, most labels got flipped based on $\phi$–indeed, the exact proportion marked by $\phi$–and no information was gained from the labelling.
+In both experiments both the LLM and DT evaluators consistently had accuracies close to the known cases in IP/OOP, but distinct success rates in IP and OOP.
+This indicates that the No-Data Algorithm is robust to evaluator strategies attempting deception.
+The large number of flips in the $\phi=0.5$ DT OOP case (Section 6.1) and $\phi=0.3$ LLM OOP (Section 6.2) support our interpretation of Theorem 5.2.
+Since the model’s accuracy is near-random in the OOP dataset, the verifier will almost always catch a lie.
+Hence, most labels were flipped and no information was gained from the labelling.
 It is worth noting that the success rate remained steady regardless of the value of $\phi$, even when the flips and accuracies didn’t.
-This means that it is an interpretable quantity, while $\phi$ is tuneable to attain ‘good’ labelled values.
-Finally, it is worth noting that the evaluation of LLMs with LLMs, as per Section 6.2, depends strongly on the LLMs themselves.
-This is not a circular argument per se.
-The judge (of judges) is tuned on the IP case, and asked to evaluate an unkown model in the OOP case, as done in Section 6.2 and Appendix D.3, even when it had near-random performance in this task.
-This is because they are solving different tasks, and thus the accuracies did adjust to the known cases.
-From the perspective of future research directions, it does raise the question on what are the results measuring.
+Thus, success rate is the main metric to measure: $\phi$ may be tuneable post-hoc to attain better performance.
+In our synthetic data scenario (Section 6.1), we noted that LLMs struggled with generating data for that problem, and picking it from a set was more effective.
+However, this was not the case in the natural-language scenario (Section 6.2).
+We argue that this is because LLMs are primarily trained with natural language, and identification (classification) and generation are two distinct problems (Kleinberg & Mullainathan, 2024); and certain automata-based problems like ours are harder for LLMs (De Wynter, 2026).
+From the perspective of our work, however, how an evaluator generates a datapoint makes no difference, so long as it can prove that it knows what it is doing.
+Finally, remark that the evaluation of LLMs with LLMs, as in Section 6.2, depends strongly on the LLMs themselves.
+This is not a circular argument: the verifier (i.e., the prompt) was tuned on the IP case, and asked to evaluate an unknown LLM (r.
+evaluation prompt) in the OOP case, as done in Section 6.2 and Appendix D.4, even when it had near-random performance in this task.
+This is because the evaluator and verifier are solving different tasks, and hence why the accuracies adjusted to the known cases.
+We provide experiments in Appendix D–and specifically, Appendix D.4–to confirm this with different LLMs.
+However, from the perspective of future research directions, it raises the question on what are the (literature’s) results measuring.
 We come back to this point in the next section.
+7.3 The Need for a Rubric
+A key limitation of the No-Data Algorithm is the requirement of a well-designed rubric, and its performance hinges on it.
+If the rubric or aggregator is poorly defined, confidence in the algorithm’s results may be misplaced, and its reliability bounded by rubric quality.
+In other words, the No-Data Algorithm measures what it is told to measure.
+Having a rubric is a fundamental aspect of the scientific method: verifying a hypothesis $H_{0}$ requires measuring a set of experimental outcomes, which are then used to accept or reject $H_{0}$.
+For this, outcomes must be interpreted (i.e., evaluated as criteria) and then aggregated.
+Thus, the responsibility for designing an effective rubric depends on the scientist (user), not the algorithm.
+This requirement is not unique to our work, or even annotation in machine learning.
+For example, there is work on rubric-based learning in reinforcement learning (He et al., 2025; Shao et al., 2025; Huang et al., 2025; Masters et al., 2025); and in evaluation with (not of) LLMs-as-judges (Lei et al., 2025).
+However, needing a rubric and characterising it are two distinct considerations.
+For the latter, there are two cases that merit special attention: decomposability, and ambiguity.
+For decomposability, note that the algorithm requires this property for the proofs, but not necessarily at runtime.
+For example, the experiments from Section 6.2 used a decomposable rubric only in the OOP case.
+Since the No-Data Algorithm is agnostic to the data labels ($\sigma$ is unknown), it worked as intended.
+Regarding ambiguity, recall that the rubrics from Appendix E.1 were ambiguous, and the experiments nevertheless succeeded.
+This further illustrates our argument on $H_{0}$: the performance of the LLMs on the (human-annotated) data was measured with respect to the human labels, which themselves were based on the rubric.
+Both annotators and LLMs used the same rubric.
+Hence, the measurement performed in Section 6.2 is whether the evaluator could act as a labeller the same way a human would.
 8 Conclusion
-We introduced an algorithm to determine trustworthiness of an evaluator in the case where labelled data does not exist.
-It works by performing randomly-selected checks evaluating specific parts of the datapoint to be labelled.
-We showed that these checks were sufficient and necessary to determine (w.h.p.) whether the evaluator knew the labels.
-Experimentally, the results adjusted to the theoretical predictions, with the No-Data Algorithm being able to determine lying evaluators under various scenarios, both synthetic and realistic.
-Of note was when the evaluator was randomly guessing the labels: while its accuracy was deceptively close to that of its known case, the success rate of the No-Data Algorithm was low, thus flagging that said labelling could not be trusted.
-Since this quantity remained steady across all values of $\phi$, it makes it a reliable metric to determine trust without needing tuning.
-One implication of the No-Data Algorithm’s correctness proof is that, when the evaluator cannot be fully trusted, the data cannot be annotated.
-This scenario is unsolvable: there is no prior information on the data, and hence it would be effectively creating information out of thin air otherwise.
-This is why the algorithm’s focus is on the successes: if an evaluator has low accuracy, but high success rate, it is very likely it is not lying; just wrong (e.g., the IP case in the experiments).
-Alternatively, high accuracy but low success rate indicates deception (r.
-the OOP case).
-Hence, the primary purpose of this algorithm is to establish trustworthiness and is up to the user to decide (to its own margin of trust) whether to accept or reject the labelling.
-It is well-known that LLMs are sensitive to prompting.
-Establishing trust in an LLM evaluator in our setup was equivalent to establishing trust in the prompt, not the model.
-While effective in the two experiments we ran, this nuance could require extensions to the No-Data Algorithm to determine (w.h.p.) whether it is the model or the prompt what is failing to convince the verifier.
-As a starting point, we note that our realistic experiment slightly departed from the theoretical work by not assuming that the criteria was known, and instead relied on the evaluator and verifier to generate their own judgement of both $C$ and $\tilde{y}$.
-Nonetheless, our work has shown that it is possible, theoretically and empirically, to judge an evaluator–such as an LLM-as-a-judge–up to formal (mathematical) guarantees of correctness.
-The other implication of our work is that the algorithm’s setup allows the verifier to ‘self-calibrate’, by generating its own datapoints, since it knows which challenges will appear.
-This has both theoretical and practical implications that we leave for future exploration.
-9 Ethics
+Trust in a measurement, although foundational to the scientific method, is difficult to determine when labelled data is scarce.
+In this work we introduced the No-Data Algorithm to formally establish trust in an evaluator when labelled data does not exist.
+This algorithm works by posing carefully-designed challenges, inspired by zero-knowledge proofs, to an evaluator.
+We proved mathematically and empirically that this method is sufficient to establish an evaluator’s trustworthiness, or, alternatively, to detect deception by either sycophantic or uninformed models.
+Our experiments in synthetic and natural-language scenarios showed close agreement to the theory, as well as some robustness to realistic settings (e.g., ambiguous rubrics and undefined criteria).
+Ablation studies with other LLMs and label arities showed extensibility and reinforced the findings.
+In our theoretical work we noted that an implication of the No-Data Algorithm is that, when the evaluator cannot be fully trusted, the data cannot be annotated.
+This is because there is no prior information on the data, and hence it would be effectively creating information out of thin air.
+This is why the algorithm’s focus is on the success rate.
+If an evaluator has low accuracy, but high success rate, it is very likely uninformed (e.g., the IP case in the experiments).
+High accuracy but low success rate indicates deception (r.
+OOP).
+Hence, it is up to the user to decide (to their own margin of trust) whether to accept or reject the labelling with respect to the rubric they designed.
+As noted in Section 9, we urge caution when interpreting the algorithm’s outputs, as they depend on the rubric itself.
+We argue that our work reframes how evaluator reliability can be measured in situations with scarcely-labelled data.
+By establishing formal measures of trust, it is now possible to perform rigorous evaluations of ‘hard’ domains such as low-resource languages and medicine.
+This is particularly important due to the debates on the reliability and deployment-readiness of LLMs-as-judges, especially in sensitive applications.
+In terms of further work, there are two open questions remaining, not addressed here due to scope: (1) how can one fully evaluate a model’s (not a prompt’s) trustworthiness?
+This is particularly important for open-ended scenarios, such as dialogue.
+Here, the principles behind the EV protocol still apply, albeit the probabilistic guarantees and challenges would differ.
+Hence, (2) what practical adaptations are needed for other data types and output signatures?
+Mathematically the full problem is solvable by assuming a binary string, but empirically this is often not feasible.
+Our work shows that it is possible to judge an evaluator’s–such as an LLM-as-a-judge’s– trustworthiness with formal guarantees of correctness.
+It also enables practitioners to transition from assumed evaluator competence, to provable trust, thus enabling the building of more robust and trustworthy AI systems.
+9 Impact Statement
 This work’s contributions involve the ability to establish trust in an LLM.
-We are unaware of any potential misuse of this algorithm, albeit we admit we are unable to cover all potential applications.
-We, however, argue that releasing the code publicly has more benefits than potential harms.
-For the data annotation, all annotators were contracted through an annotator services company, and compensated at a rate of 28.7 USD/hr.
+A possible risk is over-reliance on this algorithm’s measures of trust, especially if the underlying rubric is ill-posed.
+Users should understand that the No-Data Algorithm is only as valid as the rubrics provided.
+We thus urge users to not correlate high success rates with LLM infallibility.
+For the data annotation, all annotators were contracted through an annotator services company, and compensated based on seniority, at a rate starting at 28.7 USD/hr.
 Annotators were encouraged to take breaks and emphasised quality over speed through the annotation instructions.
-10 Reproducibility Statement
-All code has been included as part of the supplementary materials.
-It will be released under the MIT licence.
-The data will be released on the three licences from the original works.
-Detailed methodology, included model versioning, is in Appendix F.
-Prompts are in Appendix F.1, and also in the supplementary materials.
-Whenever available, we set the temperature to zero to ensure further reproducibility.
-Given the tendency of closed-source models to update and phase out existing models, we also include experiments with open models in Appendix D.
-The theoretical work includes the assumptions in Sections 3 and 5, and detailed proofs in Appendix A.
 Appendix A Detailed Proofs
-In this section we provide more careful treatments of the proofs of Lemma 1 (Appendix A.1) and Theorem 1 (Appendix A.4).
+In this section we provide more careful treatments of the proofs of Lemma 5.1 (Appendix A.1) and Theorem 5.2 (Appendix A.4).
 A.1 Correctness of the EV protocol
 The key in this proof is to note that accepting a label does not mean it will be the correct label, but that the challenges in the EV protocol are sufficient to ascertain said correctness with some probability.
-We separate our proof in two: the first part concerns itself with correctness, or whether the challenges will provide sufficient information to determine if the evaluator can be lying.
+We separate our proof in two: the first part concerns itself with completeness, or whether the challenges will provide sufficient information to determine if the evaluator can be lying.
 The second part deals with robustness under lying, cheating, or lucky scenarios by the evaluator.
-A.1.1 Part 1 (Correctness)
+A.1.1 Part 1 (Completeness)
 We prove the following lemma:
-Lemma 2.
+Lemma A.1.
 Let $x,x^{\prime}\in X$ be two binary strings such that $x\neq x^{\prime}$.
-In order to test whether an evaluator knows $f(x)=f(x^{\prime})$ without calling $f$, it is necessary to challenge the evaluator in both $x\equiv x^{\prime}$ and $x\cong x^{\prime}$.
-Moreover, these conditions are sufficient.
+To determine if the evaluator is lying, the checks from Challenge 1 and Challenge 2 are necessary and sufficient.
+In other words, it is necessary and sufficient to check if $x$ is similar to $x^{\prime}$.
 Proof.
 (Necessary) It follows from the definition of $C$.
-Since $f(x)=f(x^{\prime})$, testing $x\equiv x^{\prime}$ is a necessary condition, but not sufficient.
+Since $f(x)=f(x^{\prime})$, the check for $x\equiv x^{\prime}$ is a necessary condition, but not sufficient.
 This challenge yields information around the internal structure of $x,x^{\prime}$ up to permutation.
 It is necessary because it is possible to have two $x,x^{\prime}$ such that $x^{\prime}\not\in\{\pi_{i}(x)\}$ (i.e., two strings that aren’t permutations of one another), and yet $C(x^{\prime})=C(x)$.
-However, this does not prove that x is similar to x’, since there could be a subset $t\in S_{x^{\prime}}.t\not\in S_{x}$ and $c(t)=c(s)$ for some $c\in\mathcal{C}$ (for example, if $c$ contains a disjunctive clause).
+However, this does not prove that $x$ is similar to $x^{\prime}$, since there could be a subset $t\in S_{x^{\prime}}.t\not\in S_{x}$ and $c(t)=c(s)$ for some $c\in\mathcal{C}$ (for example, if $c$ contains a disjunctive clause).
 Hence running this challenge provides information as to whether the encoding of $C(x^{\prime})$ contains all information fed into $\sigma$ for a specific $x,x^{\prime}$.
@@
 Since neither of them provide information about the other, it follows that both are needed to fully determine $f(x)=f(x^{\prime})$.
-(Sufficient) It follows directly from the earlier statement.
-Suppose there exists a pair $x,x^{\prime}$ such that either (or both) $\not\equiv x^{\prime}, x≇x$$\not\cong x^{\prime}$ and also $f(x)=f(x^{\prime})$.
-This would imply that (1) either the string , a contradiction.
+(Sufficient) It follows directly from the above.
+Knowing $C(x)=C(x^{\prime})$ solely implies that the final encodings for $x$ and $x^{\prime}$ are equivalent, not that they are structurally equivalent.
+Conversely, knowing $\bar{C}(x),\bar{C}(x)$ does not necessarily yield information as to how these are converted into $C(x)$ and $C(x^{\prime})$.
+Hence it is sufficient to know both quantities to determine if $x$ is similar to $x^{\prime}$.
+Consequentially, testing if $x$ is similar to $x^{\prime}$ naturally implies that these challenges are sufficient to test $f(x)=f(x^{\prime})$ (via $x\equiv x^{\prime}$), and necessary (via $x\cong x^{\prime}$) to determine if the evaluator is lying.
 This concludes the proof. ∎
-While a natural consequence of Lemma 2 is that only checking for isomorphism is required for testing equality, we note that the goal of the challenges is to determine if the evaluator is lying, not the label itself.
 A.1.2 Part 2 (Robustness)
 We prove the following lemma:
-Lemma 3.
+Lemma A.2.
 The probability of error (i.e., the verifier accepts a $y^{\prime}\neq y$) at the $r^{\text{th}}$ round is $\left(1/4\right)^{r}$.
@@
 There are then two (disjoint) possibilities for a generated $\langle x^{\prime},y^{\prime}\rangle$: either that the evaluator is lying, but gets lucky; or that it is cheating and can generate the answer for one of the challenges.
-It is worth noting that generating an answer to pass both challenges is equivalent to knowing $f$.
+Remark that generating an answer to pass both challenges is equivalent to knowing $f$.
+We thus focus on the two possible lies from the evaluator, which would allow it to pass either challenge.
 A.2 Lie 1: Generate a ‘good’ $x^{\prime}$
@@
 Therefore, the probability of the verifier to not catch the evaluator’s lie is $1/4$ in any round. At the $r^{\text{th}}$ round, $\left(1/4\right)^{r}$. This concludes the proof. ∎
-It follows from Lemma 2 that we know that both challenges together are sufficient to ascertain the truthfulness of a label; and from Lemma 3 we know that the ability of the verifier to do this is bounded.
-Putting both lemmas together we obtain the proof for Lemma 1.
+It follows from Lemma A.1 that both challenges are necessary and sufficient to decide if an evaluator is lying; and from Lemma A.2 that the ability of the verifier to do this is bounded by a probability of error of $(1/4)^{r}$.
+Putting both lemmas together we obtain the proof for Lemma 5.1.
 A.4 Correctness Bounds of the No-Data Algorithm
 By a simple application of the union bound on the probability that the evaluator was wrong, but not detected by the verifier, and the probability that the evaluator was wrong, detected by the verifier, but without having its label flipped.
-For this, it is easier to work with errors $\epsilon=1-\alpha$, and remark that, from Lemma 1 the first term (undetected, no flip) is given by
+For this, it is easier to work with errors $\epsilon=1-\alpha$, and remark that, from Lemma 5.1 the first term (undetected, no flip) is given by
 $\Pr[\text{undetected}]$ $=\sum_{f(x)\neq y}\left(\frac{1}{4}\right)^{r}\text{, and hence}$ (4)
@@
 Appendix B Characterisation Bounds
-Lemma 4.
+Lemma B.1.
 […]
 Proof.
-Straightforward. ∎
+Straightforward by noting the number of calls in the algorithm. ∎
+Remark B.2.
+The runtime of the No-Data Algorithm is linear (up to a factor of $r$), but this algorithm is designed to calibrate an evaluator, not label a full dataset.
+Thus, $D$ and $r$ may be small, as in Section 6.2.
 Appendix C Extensions
-C.1 Distance-based Constraints
+C.1 $k$-ary Label Sets
+To convert a dataset of arity $k>2$, it is simply a matter of following a ‘one-vs-all’ breakdown, rerunning the No-Data algorithm k1k-1 times.
+On each run, the label set $\{l_{1},\dots,l_{k}\}$ is mapped as $\{l_{1}\},\{l_{2},\dots,l_{k}\}\mapsto\{0\},\{1\}$.
+In these cases, the runtime of the algorithm is increased by a factor of k1k-1, but the theoretical guarantees hold.
+It is then worthwhile considering whether it is possible to circumvent the runtime by simply running the No-Data algorithm without any modifications, except the label flip, and still obtain similar results.
+To test this, we ran the same setup from Section 6.2, but splitting the aggregator into three labels:
+  1. 1.
+    Label 0: All of the $c_{i}\in\bar{C}$ evaluate to zero.
+  2. 2.
+    Label 1: All, except one, of the $c_{i}\in\bar{C}$ evaluate to zero.
+  3. 3.
+    Label 2: None of the $c_{i}\in\bar{C}$ are zero.
+The results are in Table 4.
+Overall, it is possible to see that the results largely adhere to the previously-observed binary label settings.
+Namely, lying evaluators have near-zero successes and their original performance is close to the known case, with a best-of difference of 6% accuracy.
+This difference is comparable to the binary case in Table 8 for an untuned $\phi$.
+Indeed can observe a slight improvement with respect to the known case when the label flip is disabled.
+As shown in Section 5 (and experimentally in Appendix D.3), the label flip is required to achieve full restoration of the original score, which will require further algorithmic developments for a $k$-ary setting, and are beyond the scope of this work.
+In sum, while empirically the No-Data Algorithm may be used for $k>2$, further theoretical development is needed to extend this algorithm to non-‘one-vs-all’ strategies with the same guarantees.
+We then may pose the following:
+Conjecture 1:
+Let $Y$ be a label set with arity $k>2$, and $D$ a dataset.
+Then, for any number of rounds $r$, there exists an $n$ such that a version of the No-Data Algorithm runs in $\text{O}(nr|D|)$ steps, where 1<m<n<k11<m<n<k-1 and $m$ depends on $k$.
+For a ‘version’ of the No-Data Algorithm, we mean that it maintains the same correctness bounds from Theorem 5.2.
+Table 4: Ablation study reframing the setup from Section 6.2 into a ternary-labelled experiment (0,1,2{0,1,2}).
+We report the results with accuracy and $F_{\mu}$ (macro) score.
+For comparison, a random guesser obtains 33 accuracy and 30 $F_{\mu}$.
+In this experiment, the known case was still reconstructable from the algorithm’s run, although with more noise.
+Disabling label flips improved restoration of the original scores.
+Full reconstruction would require the development of further algorithmic strategies to retain all guarantees from the No-Data Algorithm.
+IP Acc. / Fμ OOP Acc. / Fμ IP Succ. / Flips OOP Succ. / Flips
+Known
+Predicting $y$ 71.8 / 61.1 10.5 / 7.0 –/– –/–
+Unknown
+$\phi=0.1$
+   Flip 66.0 / 55.7 12.2 / 10.8 39.4 / 4.3 0.4 / 11.3
+   No flip 67.0 / 57.1 10.1 / 6.4 39.4 / – 0.4 / –
+$\phi=0.3$
+   Flip 59.8 / 52.3 20.0 / 17.8 38.1 / 18.1 0.8 / 34.2
+   No flip 61.6 / 55.1 10.5 / 7.0 38.1 / – 0.8 / –
+$\phi=0.7$
+   Flip 39.5 / 40.0 30.9 / 21.1 35.3 / 47.4 0.8 / 71.6
+   No flip 46.8 / 43.4 0.10 / 6.7 35.3 / – 0.8 / –
+C.2 Distance-based Constraints
 The constraint that $C(x)=C(x^{\prime})$ is too strict for some applications.
@@
 Suppose now that the failure is given whenever $\delta(C(x),C(x^{\prime}))\geq\epsilon$, for some $\epsilon\in[0,1/2]$.
-The bounds of correctness are roughly equivalent to these described in Lemma 1 and Theorem 1.
-However, the term $(1/4)^{r}$ now takes a dependency on the probability of generating a string that is far from the desired constraint.
-C.2 $k$-ary Label Sets
-To convert a dataset of arity $k>2$, it is necessary to rerun the No-Data algorithm k1k-1 times.
-On each run, the label set $\{l_{1},\dots,l_{k}\}$ is mapped as $\{l_{1}\},\{l_{2},\dots,l_{k}\}\mapsto\{0\},\{1\}$.
-In these cases, the runtime of the algorithm is increased by a factor of k1k-1.
+The bounds of correctness are roughly equivalent to these described in Lemma 5.1 and Theorem 5.2.
+However, the term $(1/4)^{r}$ now takes a dependency on the probability of generating a string that is far (based on $\delta$) from the desired constraint, and thus the bounds become weaker when said probability is larger than $1/2$.
 Appendix D Full Results
-D.1 Ablation: Lying Evaluators and Lying Generators
-The results are in Table 5. Throughout that study DT randomly (w.p. $1/10$) flipped the label when evaluating. The lies tested were (in order in the table) when the model claims to
-  1. 1.
-    know $\sigma$: that is, it can output an encoding $C(x)$, but does not necessarily know $\bar{C}(x)$;
-  2. 2.
-    know $f$: that is, it can retrieve a datapoint $x^{\prime}\in X$ such that $y^{\prime}=y$, but nothing else; and
-  3. 3.
-    approximately know $f$ up to a probability $p=1/10$ (r. it could pass both challenges with probability 1p1-p).
-The last ‘lie’ is more akin to how a realistic generator would behave, sometimes understanding the challenges well, but failing them with some probability.
-As before, the results indicate that lying evaluators tend to score low in the success rate, even when their accuracies are close to the known case.
-Remark that (a) the evaluator that lies on knowing $\sigma$ had a success rate much higher than that of lying of knowing $f$; and (b) the evaluator with approximately knowledge of $f$ had a comparatively low (44%) success rate.
-This success rate depends on the cardinality of $\bar{C}$: when it is equal to $\mathcal{C}$, the model will always pass Challenge 1 (and fail Challenge 2).
-However, when $\lvert\bar{C}\rvert>\lvert C\rvert$, there will be criteria that the evaluator cannot possibly solve without knowing how $C(x)$ is constructed based on its internal structure, and hence it will fail them with some probability.
-D.2 Ablation: LLM Generation Strategies
-Given the LLMs’ ability to follow instructions, it is likely that their performance could be sensitive to how the datapoints are generated during the EV protocol.
-Hence in this study we compared the performance of both the LLM whose results we presented in the main results (o3-mini) with another LLM, GPT-4o [OpenAI, 2024].
-We tested two different generation strategies: picking a matching datapoint from the data, as in Section 6.1.2, and actually generating a datapoint from the rubric.
-The results are in Table 3.
-We found that, again, the experimental results adjusted to the predictions, with lying and low-performing models being easily spotted by the No-Data Algorithm.
-However, we also found that actual generation led to bad performance in both LLMs.
-We note that this is expected: in general, classifying an element in a language is much harder than generating it [Kleinberg and Mullainathan, 2024, Kalavasis et al., 2025].
-o3-mini (IP) o3-mini (OOP) GPT-4o (IP) GPT-4o (OOP)
-Picking
-Accuracy/F1 97.6 / 97.6 59.0 / 70.0 58.0 / 67.1 56.4 / 65.4
-Successes/Flips 81.3 / 1.8 27.9 / 6.0 28.7 / 71.3 10.8 / 8.8
-Generating
-Accuracy/F1 94.2 / 94.2 60.6 / 65.5 56.6 / 63.8 53.4 / 63.6
-Successes/Flips 55.4 / 5.0 28.7 / 7.8 5.8 / 10.8 34.5 / 4.6
-Table 3: Ablation study on LLM generation strategies, comparing a high-performing model (o3-mini) with a lower-performing model (GPT-4o) as a LLM-based baseline.
-We report the highest-performing run at $\phi=0.1$.
-Both models were able to broadly adjust to their original accuracy, but success rates are markedly different when the models are generating a datapoint versus picking one.
-We also performed a study on the overall performance of various models under the generation strategy, comparing them to the known case.
-This study includes results with DeepSeek [DeepSeek-AI et al., 2025], GPT-4o, and Qwen 2.5 VL 7B [Qwen Team, 2025] (Table 4) Outside of o3-mini, all models–open or not–struggled with this problem, even in the known case.
+D.1 LLM-Based Results
+In a full version of our study from Section 6.1, we evaluated multiple open and closed LLMs.
+The LLMs studied were o3-mini (as reported in the main section), DeepSeek (DeepSeek-AI et al., 2025), GPT-4o, and Qwen 2.5 VL 7B (Qwen Team, 2025), all under the same experimental setup from Section 6.1.
+The results can be found in Table 5.
+Outside of o3-mini, all models struggled with this problem, even in the known case.
 Consequentially, their performance in the unknown case also adjusted to the predictions.
-When the accuracy was high in OOP settings, the success rate was low–as expected.
+It is worth noting that, even when their accuracy was (relatively) high in OOP, their success rate was low–as expected.
+This is particularly noticeable in o3-mini, GPT-4o, and DeepSeek.
+Table 5: Comparison of the No-Data algorithm with respect to various LLMs and values of $\phi$.
+The known case reflects the LLM’s expected performance in the dataset, while the unknown case shows what the No-Data Algorithm outputs.
+All LLMs broadly adjusted to their known case.
+The LLMs also showed on average low accuracies across the board.
+However, success rates were markedly different between the known and unknown cases under high accuracy, indicating possible deception.
+When the accuracy was low, success rates remained low regardless of the setting.
 o3-mini DeepSeek GPT-4o Qwen
@@
 Successes/Flips 28.7 / 7.8 18.1 / 7.6 34.5 / 4.6 16.3 / 8.0
-Table 4: Comparison of the No-Data algorithm with respect to various LLMs and values of $\phi$.
-The known case reflects the LLM’s expected performance in the dataset, while the unknown case shows what the No-Data Algorithm outputs.
-In the unknown case, all models broadly adjusted to their known case.
-The models also showed on average low accuracies across the board.
-However, success rates were markedly different between the known and unknown cases under high accuracy, indicating possible deception.
-When the accuracy was low, success rates remained low regardless of the setting.
-For all of the above, we used the generating strategy.
+D.2 Ablation: Lying Evaluators and Lying Generators
+In this study we evaluated the impact of various aspects of lying in evaluation, both during generation and evaluation (of $C$).
+The results are in Table 6.
+The lies tested were when the model claims to:
+  1. 1.
+    know $\sigma$: that is, it can output an encoding $C(x)$, but does not necessarily know $\bar{C}(x)$;
+  2. 2.
+    know $f$: that is, it can retrieve a datapoint $x^{\prime}\in X$ such that $y^{\prime}=y$, but nothing else; and
+  3. 3.
+    approximately know $f$ up to a probability $p=1/10$ (r. it could pass both challenges with probability 1p1-p).
+The last ‘lie’ is more akin to how a realistic generator would behave (e.g., an LLM), sometimes understanding the challenges well, but failing them with some probability.
+As before, the results indicate that lying evaluators tend to score low in the success rate, even when their accuracies are close to the known case.
+Remark that (a) the evaluator that lies on knowing $\sigma$ had a success rate much higher than that of lying of knowing $f$; and (b) the evaluator with approximately knowledge of $f$ had a comparatively low (44%) success rate.
+This success rate depends on the cardinality of $\bar{C}$: when it is equal to $\mathcal{C}$, the model will always pass Challenge 1 (and fail Challenge 2).
+However, when $\lvert\bar{C}\rvert>\lvert C\rvert$, there will be criteria that the evaluator cannot possibly solve without knowing how $C(x)$ is constructed based on its internal structure, and hence it will fail them with some probability.
+Table 6: Performance for a lying DT evaluator on the No-Data Algorithm with $\phi=0.4$.
+From top to bottom: Lie 1: $\sigma$ unknown.
+Lie 2: $f$ unknown.
+Lie 3: $f$ approximately known (lying with probability $p=1/10$) in the generation step.
+In all these calls, the evaluator itself also outputs the incorrect label with probability $p=1/10$.
+Note the comparatively high success rates in the IP case of Lies 1 and 2.
+We discuss these findings in Section 7.
+Also note the consistently low success rate for OOP across all experiments, thus indicating the robustness of this algorithm to any evaluator strategy in scenarios where the data is completely unknown to the evaluator.
 IP (Flips) IP (No Flips) OOP (Flips) OOP (No Flips)
@@
 Accuracy/ F1 55.8 / 54.2 49.0 / 49.4 47.6 / 49.5 46.6 / 48.6
-Table 5: Performance for a lying DT evaluator on the No-Data Algorithm with $\phi=0.4$.
-From top to bottom: Lie 1: $\sigma$ unknown.
-Lie 2: $f$ unknown.
-Lie 3: $f$ approximately known (lying with probability $p=1/10$) in the generation step.
-In all these calls, the evaluator itself also outputs the incorrect label with probability $p=1/10$.
-Note the comparatively high success rates in the IP case of Lies 1 and 2.
-We discuss these findings in Section 7.
-Also note the consistently low success rate for OOP across all experiments, thus indicating the robustness of this algorithm to any evaluator strategy in scenarios where the data is completely unknown to the evaluator.
-D.3 Ablation: Linguistic Competency
+D.3 Ablation: LLM Generation Strategies
+In this study we accounted for the sensitivity to the prompt.
+This is because it is likely that an LLM’s performance in the EV protocol could be marked by how it generates the datapoints.
+We tested two different generation strategies: picking a matching datapoint from the data, as in Section 6.1.2; and actually generating a datapoint from the rubric.
+The dataset is the synthetic dataset, and we compared the performance of the LLM from Section 6.1 (o3-mini); and GPT-4o (OpenAI, 2024).
+The results are in Table 7.
+We found that, again, the experimental results adjusted to the predictions, with lying and low-performing models being easily spotted by the No-Data Algorithm.
+However, we also found that actual generation (versus picking) led to lower performance in both LLMs; with o3-mini having a slightly lower drop (-3%) than GPT-4o (-5%).
+Table 7: Ablation study on LLM generation strategies, comparing a high-performing model (o3-mini) with a lower-performing model (GPT-4o).
+We report the results at $\phi=0.1$.
+Both LLMs were able to broadly adjust to their original accuracy, but success rates are different when they are generating a datapoint versus picking it.
+o3-mini (IP) o3-mini (OOP) GPT-4o (IP) GPT-4o (OOP)
+Picking
+Accuracy/F1 97.6 / 97.6 59.0 / 70.0 58.0 / 67.1 56.4 / 65.4
+Successes/Flips 81.3 / 1.8 27.9 / 6.0 28.7 / 71.3 10.8 / 8.8
+Generating
+Accuracy/F1 94.2 / 94.2 60.6 / 65.5 56.6 / 63.8 53.4 / 63.6
+Successes/Flips 55.4 / 5.0 28.7 / 7.8 5.8 / 10.8 34.5 / 4.6
+D.4 Ablation: Linguistic Competency
 In this section we present an expanded version of the results from Section 6.2, including the performance of Qwen 2.5B VL 7B on the same corpus, with GPT-4.1 as a verifier.
-This model has no indications of being proficient in West Frisian, although the calibration of GPT-4.1 in IP indicates that the verifier is capable of acting as a reliable component.
-Indeed, in the IP case, where the model was given the correct rubric (i.e., it did not lie), the success rate was still low.
-In the OOP case (where it was set up to lie), the success rate was lowest.
+This model has no indications of being proficient in West Frisian.
+On the other hand, the calibration of GPT-4.1 in IP through the known case indicates that this verifier is capable of acting as a comparatively reliable component.
+Indeed, in the IP case, where the model was given the correct rubric (i.e., it did not lie), the success rate for the Qwen-based evaluator was low–thus indicating inability to solve the problem, as opposed to deception.
+Consequentially, in the OOP case (where it was set up to lie), the success rate was lowest.
 As before, this number remained steady throughout the experimentation.
 The full results are in Table 8.
+Table 8: Performance of Qwen 2.5B and GPT-4.1 on the natural-language version of the No-Data Algorithm with our ablation study.
+In this setup, Qwen 2.5B and GPT-4.1 act as evaluators, and GPT-4.1 as a verifier.
+The success rate for Qwen 2.5B was low in both the IP case (implying that the the model is not lying; just wrong) and the OOP case (where Qwen 2.5B was lying and wrong).
+Since the success rate is much lower in the latter, it follows that the No-Data Algorithm is particularly proficient at detecting confident-but-wrong evaluators.
 IP Acc. / F1 OOP Acc. / F1 IP Succ. / Flips OOP Succ. / Flips
@@
 $\phi=0.1$ 55.5 / 67.7 49.7 / 19.8 35.0 / 6.2 2.1 / 11.5
-Table 6: Performance of the LLM on the natural-language version of the No-Data Algorithm with our ablation study including a model not proficient in the tasks.
-The success rate for this model was low in both the IP case (the model not lying; just wrong) and the OOP case (the model lying and wrong), although much lower in the latter.
-Appendix E Rubrics
+D.5 Comparison With Baselines
+It was noted in Section 6 that baselines for this problem remain challenging, as they are unable to spot lying evaluators.
+Ensemble methods, for example, provide ‘a’ label, but not a measure of trust.
+Typically, this would be given by agreement metrics such as pairwise Cohen’s $\kappa$ or percentage agreement (PA).
+However, these metrics must be interpreted and calibrated, which in turn depends on the problem at hand (Braylan et al., 2022).
+Likewise, they have other drawbacks, such as sensitivity to class imbalance (Cicchetti & Feinstein, 1990; Gwet, 2010).
+These issues are hard to circumvent when there do not exist labels or references for calibration.
+On the other hand, in this scenario the No-Data Algorithm matches the known-case performance, provides a trust metric (the success rate), and spots deception.
+In this section we illustrate this by comparing two consensus-based strategies (ensembles): aggregating the predictions of three evaluators (GPT-4.1, o3-mini, and Qwen-2.5) by averaging them, and by taking a majority vote.
+The comparisons are with respect to the known case from Section 6.2, for GPT-4.1 and including IP and OOP.
+The agreement metrics are Cohen’s $\kappa$ and PA.
+The full results are in Table 9.
+Perhaps unsurprisingly, in terms of baselines, ensembling three evaluators matched or outperformed the accuracy and F1 for the (single model) IP known case.
+Nonetheless, they were marked as presenting poor agreement ($\kappa=0.38$; PA = $0.72$), as per Fleiss & Cohen (1973).
+Without calibration, poor agreement would indicate caution.
+Indeed, this value should be on average higher since the classes are balanced and the ensemble performances are good (Sim & Wright, 2005).
+Moreover, when in the presence of lying evaluators (OOP), ensembling methods had a difficult time.
+Averaging their predictions correctly predicted random performance (50.1% accuracy); but majority vote aggregation reported accuracies well above what would be expected (61.6%), thus indicating vulnerability to deception.
+Even though Cohen’s $\kappa$ was reliable in this case (noting only chance agreement, $\kappa=0.0$), PA reported a moderate agreement ($0.37$).
+In practice, these contradictory findings would require calibrating the metrics, especially if the IP case’s $\kappa$ is used as justification for the OOP case.
+This is not possible without labels.
+[^5] Realistically, in an imbalanced-labels or multi-class scenario, metrics such as PA and Cohen’s $\kappa$ are even more sensitive, thus further complicating calibration (Cicchetti & Feinstein, 1990).
+In contrast, the No-Data Algorithm provided a way to accept/reject the evaluator’s predictions (the success rate; 86.2 for IP and 1.2 for OOP) as well as a slightly higher accuracy (76.9%) when compared to the single-model IP case.
+Since no calibration is needed for any parameter, we argue that, in the unknown case (the focus of our work), it is much more helpful to use this method over other algorithms.
+Table 9: Baseline comparisons between two ensemble methods and the results of the No-Data Algorithm from Table 8.
+For measures of trust, we report Cohen’s $\kappa$ and PA for the ensembles and success rate for the No-Data Algorithm.
+In the IP case, Cohen’s $\kappa$ marks agreement as poor (as per Fleiss & Cohen 1973); even when the ensemble’s performance nears or surpasses the known case.
+In OOP, the ensembles obtained above-random performance, which is impossible given the setup.
+Thus, trusting the value of a calibrated $\kappa$ in IP for OOP would yield spurious results.
+Hence better calibration is needed, since the information provided by the agreement metrics is insufficient (i.e., it is either contradictory or does not indicate full agreement).
+This is, however, not possible in the unknown case.
+In contrast, the No-Data Algorithm does not rely on calibrations and is robust to lying evaluators, while also being able to restore the known case accuracy.
+IP trust OOP trust IP Acc. / F1 OOP Acc. / F1
+Known
+GPT-4.1 76.3 / 80.3 51.5 / 67.2
+No-Data Algorithm 86.2 1.2
+GPT-4.1 at $\phi=0.1$ 76.9 / 79.8 50.7 / 16.4
+Ensembles 0.39 / 0.72 0.0 / 0.37
+Average 80.6 / 81.3 50.1 / 0.0
+Majority vote 75.5 / 79.9 61.6 / 48.7
+Appendix E Detailed Experimental Methods
+For our first experiment, we used o3-mini (version 2025-01-31) with maximum tokens set to 50,000 and left all other parameters as default.
+GPT-4o (version 2024-05-13) and Qwen 2.5 VL 7B were called with maximum tokens set to 1,024 and temperature set to zero.
+DeepSeek (version r1-distill-qwen-32b) also had the temperature set to zero, but the maximum tokens set to 2,048 to account for the longer reasoning trace.
+For the decision tree we used the implementation from sklearn [^6] https://scikit-learn.org/stable/index.html with all parameters, except the random seed, left as default.
+We carried out the second experiment (natural language) with GPT-4.1 (version longco-2025-04-14).
+The evaluator (labelling, criteria) and verifier were set at 256 maximum tokens, while the generator (i.e., the evaluator generating similar datapoints) was set at 5,000 tokens.
+All models were called with temperature set to zero.
+LLMs were sometimes prone to return unparseable outputs. To mitigate this, in case of failure, we re-queried the models up to five times.
+All LLMs were called via the Azure OpenAI API. All data analysis was performed on a consumer-grade laptop. Random seeds are set for the training in the code and indicated in the repository.
+E.1 Rubrics
 We present the full rubrics used in our experiments.
-Rubrics 2 and 3 have the specifications for the natural-language experiments.
-Both rubrics were passed verbatim to the LLM and used during annotation.
+Tables 10 and 11 have the specifications for the natural-language experiments.
+Both rubrics were used by the human annotators, and passed verbatim to the LLMs.
+Table 10: Rubric for the natural-language evaluation (IP) experiment.
+This rubric is meant to evaluate the quality of the output of an LLM.
+The rubric above was given to both the human annotators and the LLM from the No-Data Algorithm.
 Criterion Definition
@@
 It cannot refuse to respond: if there aren’t any instructions, it should continue writing, NOT respond.
-Table 2: Rubric for the natural-language evaluation (IF) experiment.
-This rubric is meant to evaluate the quality of the output of an LLM.
-The rubric above was given to both the human annotators and the LLM from the No-Data Algorithm.
+Table 11: Rubric for the natural-language evaluation setup (OOP).
+This rubric was designed to be mostly orthogonal to the IP version.
+That meant measuring a different problem altogether, albeit with some overlaps (e.g., part of c2b overlaps with c2a in IP, c2a here is c4 in IP).
+Some criteria are completely different, such as c1 and c5.
 Criterion Definition
@@
 c5) The response must provide a summary or conclusion. It must be explicitly marked as ‘summary’ or ‘conclusion’.
-Table 3: Rubric for the natural-language evaluation (OOF).
-This rubric was designed to be somewhat orthogonal to the IF version.
-That meant measuring a different problem altogether, albeit with some overlaps (e.g., part of c2b overlaps with c2a in IF, c2a here is c4 in IF).
-Some criteria are completely different, such as c1 and c5.
-Appendix F Detailed Experimental Methods
-For our first experiment, we used o3-mini (version 2025-01-31) with maximum tokens set to 50,000 and left all other parameters as default.
-GPT-4o (version 2024-05-13) and Qwen 2.5 VL 7B were called with maximum tokens set to 1,024 and temperature set to zero.
-DeepSeek (version r1-distill-qwen-32b) also had the temperature set to zero, but the maximum tokens set to 2,048 to account for the longer reasoning trace.
-For the decision tree we used the implementation from sklearn [^3] https://scikit-learn.org/stable/index.html with the default parameters.
-We carried out the second experiment (natural language) with GPT-4.1 (version longco-2025-04-14).
-The evaluator (labelling, criteria) and verifier were set at 256 maximum tokens, while the generator (i.e., the evaluator generating similar datapoints) was set at 5,000 tokens.
-All models were called with temperature set to zero.
-LLMs were sometimes prone to return unparseable outputs. To mitigate this, in case of failure, we re-queried the models up to five times.
-All LLMs were called via the Azure OpenAI API. All data analysis was performed on a consumer-grade laptop. Random seeds are set for the training in the code.
-F.1 Prompts
-The prompts we used for our LLM evaluation on synthetic experiments are in Prompt 1 (evaluation) and Prompt 2 (generation).
+E.2 Prompts
+The prompts we used for our LLM evaluation on synthetic experiments are in Table 12 (evaluation) and Table 13 (generation).
 All exemplars are omitted for brevity.
@@
 If it failed on the fifth, we returned a random string.
-For the natural language experiments the prompts are in Prompt 4 (evaluator/verifier) and Prompt 5 (generator; part of the evaluator).
+For the natural language experiments the prompts are in Table 15 (evaluator/verifier) and Table 16 (generator; part of the evaluator).
 These prompts required no exemplars.
 When parsing errors occurred, we retried up to five times, and otherwise returned a zero-labelled value.
-You are labelling binary strings based on a rubric (given below).
-First return the parts of the criterion that match the string, and the values.
-Then return the label based on the aggregate function.
-# Rubric:
-<NATURAL-LANGUAGE RUBRIC GOES HERE>
-# Aggregation Function:
-Majority vote
-Return your answer in the form:
-|reasons|
-(list of reasons)
-|reasons|
-|label|
-(the label)
-|label|
 Table 12: System prompt for the evaluator (synthetic experiment).
-Anchors (‘|reasons|’, ‘|label|’) are used for parsing.
-The text contained within ‘|reasons|’ is, as specified by the exemplars (not pictured) a chain-of-thought approach determining criterion-by-criterion matching.
+Anchors (‘—reasons—’, ‘—label—’) are used for parsing.
+The text contained within ‘—reasons—’ is, as specified by the exemplars (not pictured) a chain-of-thought approach determining criterion-by-criterion matching.
 The rubric, explained in natural language, is inserted in the marker.
 Throughout the experiments we only used the IP rubric.
-You are given five datapoints and a set of criteria.
-Based on these, pick the datapoint that matches the criteria exactly.
-# Rubric:
-<NATURAL-LANGUAGE RUBRIC GOES HERE>
-# Datapoints:
-<DATAPOINTS GO HERE>
-Return your rationale, and then your chosen datapoint ONLY in the form:
-|datapoint|
-(the datapoint)
-|datapoint|
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
 Table 13: System prompt for the generator (synthetic experiment).
@@
 Given that the model did not follow the output format exactly, some trial-and-error was needed to ensure a comparatively high parse success rate.
-You are a datapoint generator over binary strings.
-Given a rubric (given below), a datapoint, and a label, return a similar datapoint that has the same
-label, and fulfils the same conditions as the rubric.
-For convenience, always start with the same datapoint: 0100. It will be easier to work with.
-# Rubric:
-<NL RUBRIC GOES HERE>
-Return your rationale, and then the final datapoint in the form:
-|datapoint|
-(the datapoint)
-|datapoint|
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
 Table 14: System prompt for generation used in the ablation study from Appendix D.3.
-Unlike Prompt 2, this prompt does take exemplars (not pictured): the first five entries in the training dataset, along with chain-of-thought-style rationale.
-You are an LLM evaluator. You will be given a prompt and an response in West Frisian, meant for West Frisian readers.
-Your job will be to verify if the response follows certain criteria and give a final binary score.
-Check the output against the criteria below. If it fulfils the criteria, it should be a 1. Otherwise, 0.
-If any of the criteria score a zero, the response must be zero.
-# Criteria:
-<CRITERIA GO HERE>
-# Output format:
-<OUTPUT FORMAT GOES HERE>
+Unlike Prompt LABEL:pro:generationprompt, this prompt does take exemplars (not pictured): the first five entries in the training dataset, along with chain-of-thought-style rationale.
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
 Table 15: System prompt for the natural-language evaluator and verifier.
@@
 For example, if it is meant to evaluate a single criterion, say, c2b, the definition will include only the definition of said criterion and the schema will specify solely ‘c2b’ as a key.
-Remark that the verifier will always expect the criteria from the IF rubric.
-You are a paraphraser evaluating a prompt and an output for an LLM.
-You will be given a datapoint (prompt/output), a label, and a list of reasons why that datapoint’s output has that label.
-Your job will be to return a SIMILAR prompt and output, such that the OUTPUT (1) it matches the list of reasons, and (2) matches the label.
-The output must match the values in the list of reasons.
-Here’s the rubric used for these reasons:
-<RUBRIC GOES HERE>
-Your response must be in JSON using the following schema:
- "Prompt": the new, paraphrased user prompt.
- "Output": the new, paraphrased output fulfiling the criteria.
-Only use the keys "Prompt" and "Output"
-Table 5: System prompt for the natural-language generator. In this scenario the prompt will always include the rubric (either IF or OOF), as labelled by the evaluator itself.
+Remark that the verifier will always expect the criteria from the IP rubric.
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+Table 16: System prompt for the natural-language generator. In this scenario the prompt will always include the rubric (either IP or OOP), as labelled by the evaluator itself.
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]